### v1

# Efficient Prior Selection in Gaussian Process Bandits with Thompson Sampling

Abstract
Gaussian process (GP) bandits provide a powerful framework for solving blackbox optimization of unknown functions. The characteristics of the unknown function depends heavily on the assumed GP prior. Most work in the literature assume that this prior is known but in practice this seldom holds. Instead, practitioners often rely on maximum likelihood estimation to select the hyperparameters of the prior - which lacks theoretical guarantees. In this work, we propose two algorithms for joint prior selection and regret minimization in GP bandits based on GP Thompson sampling (GP-TS): Prior-Elimination GP-TS (PE-GP-TS) and HyperPrior GP-TS (HP-GP-TS). We theoretically analyze the algorithms and establish upper bounds for their respective regret. In addition, we demonstrate the effectiveness of our algorithms compared to the alternatives through experiments with synthetic and real-world data.
1 Introduction
In Gaussian process bandits, we consider a variant of the multi-armed bandit problem where the arms are correlated and their expected reward is sampled from a Gaussian process (GP). The flexibility of GPs have made GP bandits applicable in a wide range of areas that need to optimize blackbox functions with noisy estimates, including machine learning hyperparameter tuning Turner et al. (2021), drug discovery Hernández-Lobato et al. (2017); Pyzer-Knapp (2018), chemical design Griffiths & Miguel Hernández-Lobato (2020), battery charging protocols Jiang et al. (2022), online advertising Nuara et al. (2018), and portfolio optimization Gonzalvez et al. (2019). However, most of the theoretical results in the literature assume that the GP prior is known but this is seldom the case in practical applications. Even with expert domain knowledge, selecting the exact prior to use can be a difficult task. Most practitioners tend to utilize maximum likelihood estimation (MLE) to identify suitable prior parameters. However, in a sequential decision making problem MLE is not guaranteed to recover the correct parameters.
In the literature, Wang & de Freitas (2014); Berkenkamp et al. (2019); Ziomek et al. (2024a) provided algorithms with theoretical guarantees when the kernel lengthscale is unknown. More recently, Ziomek et al. (2024b) introduced an elimination-based algorithm with theoretical guarantees for an arbitrary set of discrete priors. Their algorithm, Prior-Elimination GP-UCB (PE-GP-UCB), selects the arm and prior which provide the most optimistic upper confidence bound (UCB). If a prior generates too many incorrect predictions, then it may be eliminated. The previous work has focused on optimistic UCB methods which are known to over-explore.
In this work, we investigate the use of Thompson sampling for solving GP-bandit problems with unknown priors and we propose two algorithms. The first algorithm, Prior-Elimination GP-TS (PE-GP-TS), is an extension of PE-GP-UCB that replaces the doubly optimistic selection rule with posterior sampling and one less layer of optimism. We obtain a regret bound for PE-GP-TS of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma}_{T}})$ where $T$ is the horizon and $\hat{\gamma}_{T}$ is the worst-case maximum information gain, which matches that of PE-GP-UCB. The second algorithm, HyperPrior GP-TS (HP-GP-TS), uses bi-level posterior sampling to efficiently explore the priors and arms. We derive a regret bound of order ${\mathcal{O}}(\sqrt{T\beta_{T}\bar{\gamma}_{T}})$ where $\bar{\gamma}_{T}$ is the maximum information gain averaged across a hyperprior. The obtained bound matches that of standard GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ whenever the hyperprior is deterministic. Finally, we evaluate our methods on three sets of synthetic experiments and two experiments with real-world data. Across the experiments, our Thompson sampling based methods outperform PE-GP-UCB. Additionally, we analyze the priors selected by the algorithms and observe that HP-GP-TS selects the correct prior more often than the other algorithms.
The contributions of this work can be summarized as:
  • We propose two novel algorithms for GP-bandits with unknown prior: PE- and HP-GP-TS.
  • We provide theoretical regret bounds of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma_{T}}})$ and ${\mathcal{O}}(\sqrt{T\beta_{T}\bar{\gamma_{T}}})$ for PE- and HP-GP-TS, respectively.
  • We experimentally evaluate our algorithms on both synthetic and real-world data, demonstrating that they achieve superior performance by avoiding the optimistic exploration of PE-GP-UCB.
2 Background and problem statement
Figure 1: Agent deciding between priors based on observed data. The unknown function $f$ is sampled from a periodic prior and the observed data is in the top left. The agent discards the priors on the right since they do not match the periodic observations.
Problem statement
We consider a sequential decision making problem where an agent repeatedly selects among a set of arms and receives a random reward whose mean depends on the selected arm and is unknown to the agent. The goal of the agent is to maximize the cumulative sum of rewards over a finite time horizon. We assume that the distribution of the means, the prior, is sampled from a set of priors, the hyperprior. An effective agent must distinguish which prior the means are sampled from to ensure it explores efficiently. We illustrate this in Fig. 1.
Now, let us formally state the problem. Let ${\mathcal{X}}\subseteq[0,r]^{d}\subset{\mathbb{R}}^{d}$ denote the finite set of arms and $P$ a finite set of priors with associated prior mean and kernel functions $\mu_{1,p}:{\mathcal{X}}\mapsto{\mathbb{R}}$ and $k_{1,p}:{\mathcal{X}}\times{\mathcal{X}}\mapsto{\mathbb{R}}$, $\forall p\in P$. Let $p^{*}\in P$ denote the true prior and assume the expected reward function $f:{\mathcal{X}}\mapsto{\mathbb{R}}\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ is a sample from a Gaussian process with prior $p^{*}$. Both the function $f$ and the true prior $p^{*}$ are considered unknown. We will consider two settings: In the frequentist selection setting, the prior $p^{*}\in P$ is picked arbitrarily. In the Bayesian selection setting, the prior is sampled from a hyperprior $p^{*}\sim\mathcal{P}(P)$. To simplify notation, let $P_{0}$ denote the hyperprior.
Let $T$ denote the horizon. For time step $t=1,2,\ldots,T$, the agent selects an arm $x_{t}\in{\mathcal{X}}$ and observes the reward $r_{t}=f(x_{t})+\epsilon_{t}$ where $\{\epsilon_{t}\}_{t=1}^{T}$ are i.i.d. zero-mean Gaussian noise with variance $\sigma^{2}$. The goal of the agent is to select a sequence of arms $\{x_{t}\}_{t=1}^{T}$ that minimizes the regret $R(T)=\sum_{t\in[T]}f(x^{*})-f(x_{t})$ where $[T]=\{1,\ldots,T\}$ and $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}f(x)$. In the Bayesian selection setting, we evaluate the agent based on the Bayesian regret ${\text{BR}}(T)={\mathbb{E}}\left[R(T)\right]$ where the expectation is taken over the prior $p^{*}$, the expected reward function $f$, the noise $\{\epsilon_{t}\}_{t=1}^{T}$ and the potentially stochastic selection of arms.
Gaussian processes
A Gaussian process $f(x)\sim{\mathcal{GP}}(\mu,k)$ is a collection of random variables such that for any subset $\{x_{1},\ldots,x_{n}\}\subset{\mathcal{X}}$, the vector $[f(x_{1}),\ldots f(x_{n})]\in{\mathbb{R}}^{n}$ has a multivariate Gaussian distribution. The probabilistic nature of GPs makes them very useful for defining and solving bandit problems where the arms are correlated. Given $t$ observations $\{(x_{i},y_{i})\}_{i=1}^{t}$, the posterior mean and kernel functions of a Gaussian process ${\mathcal{GP}}(\mu,k)$ are given by
$\mu_{t}(x)$ $=\mu(x)+{\mathbf{k}}^{\top}\left({\mathbf{K}}+\sigma^{2}I\right)^{-1}({\mathbf{y}}-{\bm{\mu}}),$ (1)
$k_{t}(x,\tilde{x})$ $=k(x,\tilde{x})-{\mathbf{k}}^{\top}\left({\mathbf{K}}+\sigma^{2}I\right)^{-1}\tilde{{\mathbf{k}}}.$ (2)
Above, ${\mathbf{k}},\tilde{\mathbf{k}}\in{\mathbb{R}}^{t}$ are vectors such that $({\mathbf{k}})_{i}=k(x_{i},x)$ and $(\tilde{{\mathbf{k}}})_{i}=k(x_{i},\tilde{x})$. Additionally, ${\mathbf{y}},{\bm{\mu}}\in{\mathbb{R}}^{t}$ are also vectors such that $({\mathbf{y}})_{i}=y_{i}$ and $({\bm{\mu}})_{i}=\mu(x_{i})$. The gram matrix is denoted by ${\mathbf{K}}\in{\mathbb{R}}^{t\times t}$ where $({\mathbf{K}})_{i,j}=k(x_{i},x_{j})$. Let $\mu_{t,p}$ and $k_{t,p}$ denote the posterior mean and kernel for a Gaussian process with prior $p\in P$ at time $t$ and let $\sigma^{2}_{t,p}(x)=k_{t,p}(x,x)$ denote the posterior variance at time $t$.
The kernel $k$ determines important characteristics of the functions $f$ and in the following we provide some examples. The RBF kernel, $k(x,\tilde{x})=\exp(-||x-\tilde{x}||^{2}/\ell^{2})$ guarantees that $f$ is smooth. The lengthscale parameter $\ell>0$ determines how quickly $f$ changes, smaller values lead to more fluctuations. The rational quadratic (RQ) kernel $k(x,\tilde{x})=\left(1+\frac{||x-\tilde{x}||^{2}}{2\alpha\ell^{2}}\right)^{-\alpha}$ where $\alpha>0$ is a mixture of RBF kernels with varying lengthscales. The Matérn kernel Matérn (1986) $k(x,\tilde{x})=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}||x-\tilde{x}||}{\ell}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}||x-\tilde{x}||}{\ell}\right)$ where $\nu>0$ is the smoothness parameter that imposes that $f$ is k-times differentiable if $\nu>k$ for integer $k$. The functions $\Gamma(\nu)$ and $K_{v}$ correspond to the gamma function and a modified Bessel function Williams & Rasmussen (2006). The periodic kernel $k(x,\tilde{x})=\exp\left(-\frac{1}{2}\sum_{i=1}^{d}\sin^{2}(\frac{\pi}{\rho}(x-\tilde{x}))/\ell\right)$ generates smooth and periodic functions with period $\rho>0$ Mackay (1998). The linear kernel $k(x,\tilde{x})=vx^{\top}\tilde{x}/\nu$ generates linear functions where $v$ is the variance parameter.
Information gain
The maximal information gain (MIG) is a measure of reduction in uncertainty of $f$ after observing the most informative data points up to a specified size. The MIG commonly occurs in regret bounds for GP bandit algorithms Srinivas et al. (2012); Vakili et al. (2021) and its growth rate is strongly determined by the prior kernel of the GP. Hence, we will define the MIG for any fixed GP prior $p\in P$. Let ${\mathbf{y}}_{A}$ denote noisy observations of $f$ at the locations $A\subset{\mathcal{X}}$. Then, the MIG given prior $p\in P$, $\gamma_{T,p}$, is defined as
$\gamma_{T,p}:=\sup_{A\subset{\mathcal{X}},|A|\leq T}I_{p}({\mathbf{y}}_{A};f),$ (3)
where $I_{p}({\mathbf{y}}_{A};f)=H({\mathbf{y}}_{A}|p)-H({\mathbf{y}}_{A}|f,p)$ is the mutual information between ${\mathbf{y}}_{A}$ and $f$ given $p$, and $H(\cdot)$ denotes the entropy. To aid our analysis later, we also define the worst-case MIG as $\hat{\gamma}_{T}:=\max_{p\in P}\gamma_{T,p}$ and the average MIG as $\bar{\gamma}_{T}:={\mathbb{E}}_{p\sim P_{0}}[\gamma_{T,p}]$. For the RBF and Matérn kernels, $\gamma_{T,p}={\mathcal{O}}(\log^{d+1}(T))$ and $\gamma_{T,p}={\mathcal{O}}(T^{\frac{d}{2\nu+d}}\log^{\frac{2\nu}{2\nu+d}}(T))$ Srinivas et al. (2012); Vakili et al. (2021).
Previous work
Wang & de Freitas (2014) was the first work that presented regret bounds for Gaussian process bandits with unknown lengthscale. Wang & de Freitas studied the popular Expected Improvement algorithm Močkus (1975). But their algorithm requires a lower bound on the lengthscale and their regret bound depends on the worst-case MIG. Later work by Berkenkamp et al. (2019) introduced Adaptive GP-UCB (A-GP-UCB) that continually lowers lengthscale parameter. Once the lengthscale parameter is sufficently small, then the function $f$ will lie within the reproducing kernel Hilbert space (RKHS) and then the regular GP-UCB theory can be applied. However, A-GP-UCB lacks a stopping mechanism and will overexplore as the lengthscale continues to shrink. Recent work by Ziomek et al. (2024b) introduced Prior-Elimination GP-UCB (PE-GP-UCB) for time-varying GP-bandits with unknown prior. Unlike the work before, the regret bound of PE-GP-UCB holds for arbitrary types of hyperparametersin the GP prior. PE-GP-UCB is doubly optimistic and selects the prior and arm with the highest upper confidence bound. PE-GP-UCB tracks the cumulative prediction error made by the selected priors and eliminates priors that exceed a threshold level.
Other works have introduced regret balancing algorithms that maintain a set of base learning algorithms and balance their selection frequency to achieve close to optimal regret Abbasi-Yadkori et al. (2020); Pacchiano et al. (2020). Ziomek et al. (2024a) built on this idea and introduced length-scale balancing GP-UCB which can adaptively explore smaller lengthscales but can return to longer ones, unlike A-GP-UCB.
Wang et al. (2018) provided regret bounds for probability of improvement Kushner (1964) and GP-UCB when the GP prior is learnt from offline data, which they called Meta BO. The aforementioned works are based on UCB, EI, PI or regret balancing. However, another line of work has studied Thompson sampling in standard and linear bandits with unknown prior distribution Kveton et al. (2021); Basu et al. (2021); Hong et al. (2022); Li et al. (2024). In their setting (meta or hierarchical bandits), the agent plays multiple bandit instances, either simultaneously or sequentially. The unknown means are sampled from the same (unknown) prior and by gathering knowledge across instances, the agent can solve later instances more efficiently once it has identified the prior. We emphasize that these methods have been studied only for standard stochastic and linear bandits, not for GP bandits.
3 Algorithms
As discussed by Russo & Roy (2014), TS can offer advantages over UCB algorithms for problems where constructing tight confidence bounds is difficult. In addition, Thompson sampling is often observed to perform better than UCB in practice Chapelle & Li (2011); Wen et al. (2015); Kandasamy et al. (2018); Åkerblom et al. (2023b, a). Motivated by this, we present two algorithms for efficient prior selection based on TS.
3.1 Prior-Elimination with Thompson sampling
Figure 2: Elimination procedure of PE-GP-TS. The solid lines correspond to posterior means and the shaded regions are confidence intervals. The dashed lines are samples from the posteriors. The figure has been adapted from Ziomek et al. (2024b) to illustrate the elimination step of PE-GP-TS.
Algorithm 1 Prior Elimination GP-TS (PE-GP-TS)
0:  Horizon $T$, prior functions $\{\mu_{1,p},k_{1,p}\}_{p\in P}$, confidence parameters $\{\beta_{t}\}_{t=1}^{T}$ and $\{\xi_{t}\}_{t=1}^{T}$.
1:  $P_{0}=P$, $S_{0,p}=\varnothing$ $\forall p\in P$
2:  for $t=1,2\ldots,T$ do
3:     Sample $\tilde{f}_{t,p}\sim{\mathcal{GP}}(\mu_{t,p},k_{t,p})$ $\forall p\in P_{t-1}$
4:     Set $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t-1}}\tilde{f}_{t,p}(x)$
5:     $S_{t,p_{t}}=S_{t-1,p_{t}}\cup\{t\}$ and $S_{t,p}=S_{t-1,p}$ for $p\in P\setminus\{p_{t}\}$
6:     Observe $y_{t}=f(x_{t})+\epsilon_{t}$
7:     Set $\eta_{t}=y_{t}-\mu_{t,p_{t}}(x_{t})$
8:     Set $V_{t}=\sqrt{\xi_{t}|S_{t,p_{t}}|}+\sum_{i\in S_{t,p_{t}}}\sqrt{\beta_{i}}\sigma_{i,p_{t}}(x_{i})$
9:     if $\left|\sum_{i\in S_{t,p_{t}}}\eta_{i}\right|>V_{t}$ then
10:        $P_{t}=P_{t-1}\setminus\{p_{t}\}$
11:     else
12:        $P_{t}=P_{t-1}$
Our first algorithm is an extension of the PE-GP-UCB algorithm Ziomek et al. (2024b) to be employed with Thompson sampling - instead of UCB. The key difference is that instead of maximizing the upper confidence bound $U_{t}(x,p)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$ over ${\mathcal{X}}\times P$, we instead sample $\tilde{f}_{t,p}$ from the posterior ${\mathcal{GP}}(\mu_{t,p},k_{t,p})$ for all priors $p\in P_{t-1}$ where $P_{t-1}$ is the set of active priors. Then, we select the arm and prior $x_{t},p_{t}$ such that $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t-1}}\tilde{f}_{t,p}(x)$. Whilst PE-GP-UCB has two layers of optimism, the upper confidence bound and joint maximization of $x$ and $p$. PE-GP-TS has only a single layer of optimism - which should alleviate potential overexploration issues.
The elimination procedure of PE-GP-TS is illustrated in Fig. 2. Samples $\tilde{f}_{t,p}$ are drawn from the active prior $p\in P_{t-1}$. Then, the unknown function $f$ is queried at the selected arm $x_{t}$. If the observed value differs too much from the prediction made by the selected prior, then the selected prior is eliminated. Otherwise, it remains active.
The PE-GP-TS algorithm is presented in Algorithm 1. Similar to PE-GP-UCB, the set $S_{t,p}$ is used to store the time steps where prior $p$ was selected up to and including time $t$. When prior $p_{t}$ is selected, the prediction error $\eta_{t}=y_{t}-\mu_{t,p_{t}}(x_{t})$ between the observed and predicted value made by the prior $p_{t}$ is computed. If the sum of prediction errors made by the prior $p_{t}$ exceeds the threshold value $V_{t}$, then $p_{t}$ is eliminated from the active priors $P_{t}$, see line 9. Note that at time step $t$, only the selected prior $p_{t}$ can be eliminated. As such, if a prior is very pessimistic it may never be selected and therefore will never be eliminated. Thus, the final set of active priors $P_{T}$ should be viewed as non-eliminated priors rather than necessarily being reasonable priors.
3.2 HyperPrior Thompson sampling
Algorithm 2 HyperPrior GP-TS (HP-GP-TS)
0:  Horizon $T$, prior functions $\{\mu_{1,p},k_{1,p}\}_{p\in P}$, hyperprior $P_{0}$.
1:  for $t=1,2\ldots,T$ do
2:     Sample $p_{t}\sim P_{t-1}$
3:     Sample $\tilde{f}_{t}\sim{\mathcal{GP}}(\mu_{t,p_{t}},k_{t,p_{t}})$
4:     Set $x_{t}=\operatorname*{arg\,max}_{x}\tilde{f}_{t}$
5:     Observe $y_{t}=f(x_{t})+\epsilon_{t}$
6:     Set ${P}_{t}(p)\propto{\mathbb{P}}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)P_{t-1}(p)$triangleright Update hyperposterior
In our first algorithm, we removed one layer of optimism. In our second algorithm, we adopt a fully Bayesian algorithm by using a hyperposterior sampling scheme where both the prior and the mean function are sampled from their respective posteriors. By shedding the optimism over the selected prior $p_{t}$, HP-GP-TS should be able avoid costly exploration by selecting likely priors instead of optimistic ones.
The algorithm is presented in Algorithm 2. In the first step, the current prior $p_{t}$ is sampled from the hyperposterior $P_{t-1}$. Then, a single sample $\tilde{f}_{t}$ is taken from the selected posterior ${\mathcal{GP}}(\mu_{t,p_{t}},k_{t,p_{t}})$ and is used to select the current arm: $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\tilde{f}_{t}(x)$. After observing $y_{t}$, the hyperposterior is updated by computing the likelihood of $y_{t}$ under the different priors. Note that since the set of priors $P$ is finite, computing the posterior is tractable albeit computationally costly with a complexity of ${\mathcal{O}}(t^{3}|P|)$. The likelihood ${\mathbb{P}}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)={\mathcal{N}}(y_{t};\mu_{t,p}(x_{t}),\sigma_{t,p}^{2}(x_{t})+\sigma^{2})$ is simply the Gaussian likelihood of the posterior at $x_{t}$ plus Gaussian noise with variance $\sigma^{2}$.
4 Regret bounds
In this section, we present regret bounds for the proposed algorithms. Recall from the problem statement that we consider two slightly different settings for the two algorithms. Specifically, for PE-GP-TS we assume the unknown prior $p^{*}$ is sampled arbitrarily from $P$ whilst for HP-GP-TS we assume that the unknown prior $p^{*}$ is selected from a known hyperprior distribution $P_{0}$.
Ziomek et al. (2024b) structured the proof of the regret bound of PE-GP-UCB into 4 larger steps; First, showing that $p^{*}$ is never eliminated with high probability. Second, establishing a bound on the simple regret. Third, bounding the cumulative regret. Finally, the cumulative bound is re-expressed in terms of the worst-case MIG. For PE-GP-TS, we note that it is sufficient to establish a new bound on the simple regret and then adapt the steps of Ziomek et al. to accommodate the new simple regret bound.
To bound the simple regret, we require two concentration inequalities to hold for both the posteriors and the posterior samples which we present in the following lemma.
{lemma}
[] For any $p^{*}\in P$, if $f(x)\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ and $\beta_{t}=2\log\left(\frac{|{\mathcal{X}}||P|\pi^{2}t^{2}}{3\delta}\right)$. Then, with probability at least $1-\delta$, the following holds for all $t,x\in[T]\times{\mathcal{X}}$:
$|f(x)-\mu_{t,p^{*}}(x)|\leq\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x),$ (4)
and for all $t,x,p\in[T]\times{\mathcal{X}}\times P:$
$|\tilde{f}_{t,p}(x)-\mu_{t,p}(x)|\leq\sqrt{\beta_{t}}\sigma_{t,p}(x).$ (5)
All proofs can be found in Appendix A. Section 4 is based on Lemma 5.1 of Srinivas et al. (2012) but adapted to TS by specifying that it holds for any sequence of $x_{1},\ldots,x_{T}$, as discussed by Russo & Roy (2014). Additionally, we add Eq. 5 which can be shown through the same steps and an additional union bound over $P$. Next, we state our bound for the simple regret of PE-GP-TS.
{lemma}
[] If the event of Section 4 holds, then the following holds for the simple regret of PE-GP-TS for all $t\in[T]$:
$f(x^{*})-f(x_{t})\leq 2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-\eta_{t}+\epsilon_{t}.$ (6)
Compared to the simple regret bound for PE-GP-UCB, we obtain the additional term $2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})$. Since $p^{*}$ is fixed, the sum over $t$ of the new term is ${\mathcal{O}}(\sqrt{T\beta_{T}\gamma_{T,p^{*}}})$ and we obtain the following regret bound:
Theorem 4.1.
If $p^{*}\in P$ and $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, then PE-GP-TS with confidence parameters $\beta_{t}=2\log(2|{\mathcal{X}}||P|\pi^{2}t^{2}/3\delta)$ and $\xi_{t}=2\sigma^{2}\log(|P|\pi^{2}t^{2}/3\delta)$, satisfies the following regret bound with probability at least $1-\delta$:
$R(T)\leq$ $2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}$ (7)
$+2\sqrt{CT\beta_{T}\gamma_{T,p^{*}}}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}$
where $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$ and $C=2/\log(1+\sigma^{-2})$.
Since $\gamma_{T,p^{*}}\leq\hat{\gamma}_{T}$, the bound is of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma}_{T}})$ w.r.t. $T$ which matches that of PE-GP-UCB. To our knowledge, the best lower bound for standard GP bandits in the Bayesian setting, where $f$ is sampled from a GP, is $\Omega(\sqrt{T})$ for $d=1$ Scarlett (2018). This would suggest that our bound is tight up to a factor ${\mathcal{O}}(\sqrt{\beta_{T}\hat{\gamma}_{T}})$.
[…]
{lemma}
[] Let $C=2/\log(1+\sigma^{-2})$ and assume $\sigma_{t,p}(x)\leq 1$, then for any sequence of arms $x_{1},\ldots,x_{T}\in{\mathcal{X}}$ and any $p\in P$
$I_{p}({\mathbf{y}}_{T};f)=\frac{1}{2}\sum_{t\in[T]}\log\left(1+\sigma^{-2}\sigma_{t,p}^{2}(x_{t})\right),$ (8)
${\sum_{t\in[T]}}\sigma^{2}_{t,p}(x_{t})\leq CI_{p}({\mathbf{y}}_{t};f).$ (9)
{corollary}
[] For any stochastic or deterministic sequence of arms $x_{1},\ldots,x_{T}\in{\mathcal{X}}$ and $p^{*}\sim P_{0}$,
${\mathbb{E}}\Big{[}{\sum_{t\in[T]}}\sigma_{t,p^{*}}^{2}(x_{t})\Big{]}\leq C\bar{\gamma_{T}}.$ (10)
The proof of Theorem 4.1 is based on Lemma 5.3 and 5.4 in Srinivas et al. (2012) but is adapted to randomized selection by specifying that it holds for any sequence of arms. Theorem 4.1 follows by Theorem 4.1 and the definition of maximum information gain. Next, we state our regret bound for HP-GP-TS.
Theorem 4.2.
If $p^{*}\sim P_{0}$ and $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, then the Bayesian regret of HP-GP-TS is bounded by
${\text{BR}}(T)\leq\frac{\pi^{2}}{6}+\sqrt{C_{1}T\beta_{T}\bar{\gamma}_{T}}$ (11)
where $C_{1}=2\log(1+\sigma^{-2})$ and $\beta_{t}=2\log(|{\mathcal{X}}|t^{2}/\sqrt{2\pi)}$.
Unlike PE-GP-TS and PE-GP-UCB, the regret bound for FB-GP-TS depends on the average $\bar{\gamma}_{T}$ rather than the worst case $\hat{\gamma}_{T}$ which can impact the theoretical regret significantly if the complexity of learning the priors differ. Although, this is reasonable since the elimination methods assume arbitrary selection of $p^{*}$ as opposed to sampling from a hyperprior. If the hyperprior is deterministic then the regret bound for FB-GP-TS matches that of GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ Takeno et al. (2024) and the average $\bar{\gamma}_{T}$ would be equal to the worst case $\hat{\gamma}_{T}$. Again, using the lower bound of Scarlett (2018), our upper bound would tight up to a factor of ${\mathcal{O}}(\sqrt{\beta_{T}\bar{\gamma}_{T}})$.
Figure 3: Prior mean functions for the hill experiment. Mean function 0 consists of 10 smaller hills whereas mean functions 1 to 4 have 9 smaller hills and 1 taller hills.
Figure 4: Cumulative regret for synthetic experiments with varying kernel (left), lengthscale (center) and mean function (right). The average final regret for PE-GP-UCB is 116.5 in the lengthscale experiment. Errorbars correspond to ±$\pm 1$ standard error.
Figure 5: Average amount of priors remaining in $P_{t}$ over time for PE-GP-UCB and -TS.
5 Experiments
In this section, we describe our experiments based on synthetic and real-world data.
Synthetic experiments
We consider three synthetic setups with different choices of priors in $P$. For the first setup, the priors have varying kernels selected according to one of the following kernels: i) squared-exponential kernel, Matérn kernel with $\nu=5/2$, ii) Matérn kernel with $\nu=3/2$, iii) periodic kernel with period $\rho=5$, iv) linear kernel with $v=0.05^{2}$, and, v) the rational quadratic kernel with $\alpha=0.5$. All kernels use a lengthscale of $1.0$ and are scaled s.t. $k(x,\tilde{x})\leq 1$. In addition, the mean function for all priors are zero everywhere. For the second setup, the priors use the squared-exponential kernel with lengthscales $4$, $2$, $1$ or $1/2$. The third setup is based on the toy problem in an earlier version of Ziomek et al. (2024b) where the mean functions vary. The first mean function has 10 smaller ‘hills’ across the first half of the interval. The four other mean functions are the same except the height of one of the hills is doubled in height, see Fig. 3. All priors in the third setup use a squared-exponential kernel with lengthscale 1. For all of the setups, the true prior $p^{*}$ is sampled uniformly from $P$, 500 arms are equidistantly spaced in $[0,20]$, the noise variance $\sigma^{2}=0.25^{2}$, and the horizon $T=500$. The prior elimination methods use $\delta=0.05$. All models are evaluated on 500 seeds on each setup. As baselines, we use PE-GP-UCB and Maximum A Posteriori (MAP) GP-TS where MAP GP-TS is identical to HP-GP-TS except for greedily selecting $p_{t}$ from the posterior: $p_{t}=\operatorname*{arg\,max}_{p}P_{t-1}(p)$. Regardless of the selected prior, HP-GP-TS observes an observation $y_{t}$ to update the hyperposterior $P_{t}$. Hence, greedily selecting the prior could reduce unnecessary exploration. In addition, we investigate the oracle variants of PE-GP-TS and PE-GP-UCB with $\delta=0.05$ that are only given the true prior: $P_{0}=\{p^{*}\}$.
Figure 6: Average entropy in the hyperposterior $P_{t}$ over time for HP- and MAP GP-TS. The dashed reference values correspond to entropies of discrete distributions with prob. $q$ on one choice and prob. $\frac{1-q}{|P|-1}$ on the other $|P|-1$ choices.
The cumulative regret for the three synthetic experiments is shown in Fig. 4. Across all three experiments, we observe that HP-GP-TS has lower regret than the other methods and performs close to the oracle GP-TS. For the kernel and hill experiments, PE-GP-TS has lower regret than the oracle GP-UCB. Hence, even if PE-GP-UCB was optimized to perform as well as the oracle, it would still not achieve the regret of our proposed methods. MAP GP-TS has slightly higher regret than HP-GP-TS for the lengthscale and hill experiments but has significantly higher regret for the kernel experiment. The greedy selection causes under-exploration for MAP GP-TS in certain instances.
In the three left-most columns of Fig. 5, the average amount of priors remaining is shown over time for the synthetic experiments (for the PE-methods methods). For the hill experiment, no priors are eliminated for either PE-GP-TS nor PE-GP-UCB. For the kernel and lengthscale experiment, PE-GP-TS eliminates more priors than its UCB-counterpart. However, note that, on average, less than 0.15 priors are eliminated in the lengthscale experiment and slightly more than 1.2 priors in the kernel experiment.
Similarly, in Fig. 6, the average entropy in the hyperposterior $P_{t}$ is shown over time. Across the three experiments, the hyperposterior of HP-GP-TS concentrates more than that of MAP GP-TS - which could be due to the greedy prior selection. As a comparison, HP-GP-TS places more than 95% of the probability mass on one prior (based on the reference values) in the lengthscale experiment where as the PE-methods barely removed priors.
Figure 7: Confusion matrices for the true prior $p^{*}$ and the selected priors $p_{t}$ across all time steps of the kernel experiment.
In Fig. 7, we visualize how often the methods select the true prior $p^{*}$ (or kernel) in the synthetic experiments as confusion matrices. PE-GP-UCB almost always selects the prior using the Matérn-3/2 kernel. This prior induces a distribution over functions that are less smooth compared to the other kernels. Hence, the Matérn-3/2 kernel produces much higher confidence intervals outside the observed data and is thus more optimistic, leading to excessive exploration. PE-GP-TS also shows a bias towards the Matérn-3/2 kernel but does not select it as frequently as PE-GP-UCB - demonstrating that one layer of optimism has been removed. The overall “accuracy” of the selected priors, i.e. ${\sum_{t\in[T]}}\mathds{1}\{p_{t}=p^{*}\}/T$ where $\mathds{1}$ is the indicator function, of the elimination-based methods is around $17\%$ in the kernel experiment compared to $62.5\%$ and $63.2\%$ for MAP and HP-GP-TS respectively. For HP-GP-TS, we observe that it can easily identify the periodic and linear kernels. However, the RBF, Matérn and RQ kernels are often confused with each other. These kernels do not have as easily distinguishable characteristics and are likely to produce similar posteriors given enough data.
Figure 8: Cumulative regret on Intel temperature data (top) and PeMS data (bottom). Errorbars correspond to ±$\pm 1$ standard error.
Real-world data
We consider two setups with real-world data. Using the Intel Berkeley dataset Madden et al. (2004), the first setup considers 46 temperature sensors with measurements over 19 days. The goal is to select the sensor with the highest temperature. The data is split into 12 training days and 7 test days. For each day in the training set, the empirical mean and covariance is computed and set as a prior. For each run, the unknown function $f$ is selected as a set of measurements from the test data and the noise variance is set to $\sigma^{2}=0.7^{2}$.
The second setup uses sensor data from 211 sensors along the I-880 highway in March 2023 from the California Performance Measurement System (PEMS) Chen et al. (2001); California Department of Transportation (2024). The goal is to select the sensor with the lowest speed in order to identify congestions. The sensor data prior to March 21st is used as training data and the rest is used as test data. Only sensor readings between 6AM and 13AM are used. For each hour in the training set (across all training days), the empirical mean and covariance is again used to construct a prior - yielding 7 priors. Again, the test data is used to sample the unknown function $f$ and the noise variance is set to $\sigma^{2}=2.25^{2}$. PE-GP-UCB and MAP-GP-TS are used as baselines to compare against PE-GP-TS and HP-GP-TS. The elimination methods use $\delta=0.05$, the horizon is set to $T=500$ and all methods are evaluated on $500$ seeds [^1] One outlier was removed from the Intel experiment, see Appendix B for more details..
The cumulative regret for the real-world data experiments is presented in Fig. 8. For the Intel temperature data, HP-GP-TS and MAP-GP-TS obtain the lowest and second lowest cumulative regret respectively. For the PeMS data, PE-GP-TS obtains the lowest cumulative regret of all methods. HP-GP-TS and MAP-GP-TS have lower regret than PE-GP-UCB but their slopes are almost linear for the final time steps. To understand this better, we visualize quantiles of the total regret in Fig. 9 and the median cumulative regret in Fig. 11 in Appendix B. The figures show that MAP- and HP-GP-TS have the lowest median regret for both experiments and hence perform best in a majority of instances. However, the 90th and 95th quantiles are considerably larger for the PeMS data which impacts the average regret significantly. Hence, for the PeMS data, the prior elimination methods seem to yield more stable results.
The two right-most columns of Figs. 5 and 6 shows the average amount of active priors and the average entropy over time in the Intel and PeMS experiment. For the PeMS data, barely any priors are eliminated whilst for the Intel data around 0.75 to 0.80 priors are eliminated on average for the PE-methods. For the Intel data, the HP- and MAP GP-TS place around than 80% to 90% of the posterior probability mass in one prior based on the reference values. On the PeMS data, the posteriors concentrate more heavily - which could be premature given the average regret we observed.
(a) Intel
(b) PeMS
Figure 9: Quantiles of the final cumulative regret on the real-world data experiments. The median is highlighted in green. The whiskers correspond to the 5th and 95th percentile and the lower and upper edge of the box show the first and third quantile.
6 Discussion
Limitations
The main limitation of all the methods we have considered is that they require the set of priors $P$ to be discrete. In addition, their computational cost scales linearly with the number of priors considered. In theory, the elimination-based methods could have lower computational cost compared to HP-GP-TS. However, in practice, priors are rarely eliminated as shown by Fig. 5. At best, one sixth of the priors is eliminated in the kernel experiment for PE-GP-TS. Since only priors that are selected can be eliminated and the confidence parameters $\beta_{t}$, $\xi_{t}$ increasing over time, including more priors could lead to less priors being eliminated overall.
Future work
To lower the computational complexity of the elimination methods, one would either need tighter confidence bounds for the elimination criteria or need to allow more than the selected priors to be eliminated. The bi-level sampling approach of HP-GP-TS provides efficient exploration but suffers from higher computational cost since it cannot discard priors. An interesting line of future work could study how the probabilities in the hyperposterior could be used to discard priors that are unlikely to be selected in the remaining rounds. It would also be interesting to study selection criteria for the prior $p_{t}$ that maximize the disagreement among the priors such that evidence for the true prior is generated quicker.
A key difference between PE-GP-TS and the meta bandits methods of Kveton et al. (2021); Basu et al. (2021); Hong et al. (2022) is that they consider multiple bandit instances sampled from the true prior $p^{*}$. Specifically, Hong et al. (2022) study a setting with arbitrary interleaving of the instances. Generalizing our results for HP-GP-TS to a similar setting would provide an interesting complement to these methods beyond the standard and linear setting.
7 Conclusion
In this paper, we have proposed two algorithms for joint prior selection and regret minimization in GP bandits based on GP-TS. We have analyzed the algorithms theoretically and provided regret bounds whose order closely match previous work. We have experimentally evaluated both algorithms on both synthetic and real-world data. We find that they both select the true prior more often and obtain lower regret than previous work due to lowering the amount of optimistic exploration.
Acknowledgements
The work of Jack Sandberg and Morteza Haghir Chehreghani was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Appendix A Proofs
In the following section, we state and prove the results shown in the main text.
A.1 PE-GP-TS
First, we state and prove concentration inequalities for $f(x)$ and $\tilde{f}_{t,p}(x)$.
See 4
Proof.
Follows by the same steps as Lemma 5.1 of Srinivas except we condition on the complete history $H_{t}$ instead of only ${\mathbf{y}}_{1:t-1}$. Additionally, for Eq. 5 we must take an additional union bound over $p\in P$.
Fix $t,x,p\in[T]\times{\mathcal{X}}\times P$. Given the history $H_{t}$, $\tilde{f}_{t,p}(x)\sim{\mathcal{N}}(\mu_{t,p}(x),\sigma^{2}_{t,p}(x))$. Using that ${\mathbb{P}}(Z>c)\leq 1/2e^{-c^{2}/2}$ for $Z\sim{\mathcal{N}}(0,1)$, we get that
${\mathbb{P}}\left(\left|\frac{\tilde{f}_{t,p}(x)-\mu_{t,p^{*}}(x)}{\sigma_{t,p^{*}}(x)}\right|>\sqrt{\beta_{t}}\right)$ $\leq\exp(-\beta_{t}/2)$ (12)
$=\frac{3\delta}{|{\mathcal{X}}||P|\pi^{2}t^{2}}$ (13)
Note that $\sum_{t\geq 1}\frac{1}{t^{2}}=\frac{\pi^{2}}{6}$y taking the union bound over ${\mathcal{X}}$, $P$ and $t\geq 1$, Eq. 5 holds w.p. at least $21-\delta/$ 2. By the same reasoning and skipping the union bound over $P$, Eq. 4 holds w.p. at least $21-\delta/$ 2. Thus, both events hold w.p. at least $1-\delta$. ∎
Next, we state three lemmas from Ziomek et al. (2024b) that are used in the proof of our regret bound.
{lemma}
(Lemma 5.1 of Ziomek et al. (2024b)) If $\xi_{t}=2\sigma^{2}\log\left(\frac{|P|\pi^{2}t^{2}}{6\delta}\right)$, then the following holds with probability at least $1-\delta$:
$\left|\sum_{i\in S_{t,p}}\epsilon_{i}\right|\leq\sqrt{\xi_{t}|S_{t,p}|}\quad\forall t,p\in[T]\times P.$ (14)
{lemma}
(Lemma 5.2 of Ziomek et al. (2024b)) Let $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$, then if $\mu_{1,p^{*}}$ and $k_{1,p^{*}}$ satisfy $|\mu_{1,p^{*}}(\cdot)|<\infty$ and $k_{1,p^{*}}(\cdot,\cdot)\leq 1$ and Section 4 holds, then
$\sup_{x\in{\mathcal{X}}}|f(x)|\leq B_{p^{*}}.$ (15)
{lemma}
(Lemma 5.3 of Ziomek et al. (2024b)) For $C=2/\log(1+\sigma^{-2})$, $\sum_{t\notin{\mathcal{C}}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\leq\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}$ where $\beta_{T}=\max_{p\in P}\beta_{T}$ and $\hat{\gamma}_{T}=\max_{p\in P}{\gamma}_{T,p}$.
Then, we state and prove the new simple regret bound for PE-GP-TS.
See 4
Proof.
First, we upper bound $f(x^{*})$ as follows
[…]
Then, we lower bound $f(x_{t})$
[…]
Combining, Eqs. 18 and 20 we obtain
$f(x^{*})-f(x_{t})\leq 2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-\eta_{t}+\epsilon_{t}.$ (21)
Finally, we state and prove the cumulative regret bound for PE-GP-TS.
See 4.1
Proof.
First, we show that the true prior $p^{*}$ is never rejected if Sections 4 and A.1 hold.
$\left|\sum_{i\in S_{t,p^{*}}}\eta_{i}\right|$ $=\left|\sum_{i\in S_{t,p^{*}}}(y_{i}-f(x_{i})+f(x_{i})-\mu_{i,p^{*}}(x_{i})\right|$ (22)
[…]
Next, we bound the cumulative regret. To establish a bound on the cumulative regret, we must separate out the rounds where priors are eliminated. Hence, define the set of critical iterations as
${\mathcal{C}}=\left\{t\in[T]:\left|\sum_{i\in S_{t,p_{t}}}\eta_{i}\right|>\sqrt{\xi_{t}S_{t,p_{t}}}+\sum_{i\in S_{t,p_{t}}}\sqrt{\beta_{i}}\sigma_{i,p_{t}}(x_{i})\right\}.$ (25)
Using Eqs. 14 and 21, we can bound the cumulative regret as follows:
${\text{BR}}(T)$ $=\sum_{t\in{\mathcal{C}}}{\text{BR}}_{t}+\sum_{t\notin{\mathcal{C}}}{\text{BR}}_{t}$ (26)
$\leq 2|P|B_{p^{*}}+\sum_{t\notin{\mathcal{C}}}2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+\sum_{t\notin{\mathcal{C}}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})+\sum_{p\in P}\sum_{t\in S_{T,p}\setminus{\mathcal{C}}}(\epsilon_{t}-\eta_{t}).$ (27)
where $B_{p^{*}}:=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$. If $t\notin{\mathcal{C}}$, line 9 in Algorithm 1 evaluates to false and hence
$\sum_{p\in P}\sum_{t\in S_{T,p}\setminus{\mathcal{C}}}-\eta_{t}\leq\sum_{p\in P}\sqrt{\xi_{T}|S_{T,p}|}+\sum_{p\in P}\sum_{t\in S_{T,p}\setminus{\mathcal{C}}}\sqrt{\beta_{t}}\sigma_{t,p}(x_{t}).$ (28)
Additionally, using Section A.1, we can bound the Gaussian noise:
$\sum_{p\in P}\sum_{t\in S_{T,p}\setminus{\mathcal{C}}}\epsilon_{t}$ $\leq\sum_{p\in P}\left|\sum_{t\in S_{T,p}\setminus{\mathcal{C}}}\epsilon_{t}\right|\leq\sum_{p\in P}\left|\sum_{t\in S_{T,p}}\epsilon_{t}\right|$ (29)
[…]
Combining the above, the cumulative regret is bounded by
${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sum_{t\notin{\mathcal{C}}}\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+2\sum_{t\notin{\mathcal{C}}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t}).$ (32)
Finally, applying Eqs. 15 and 4.1, we obtain the result
${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}\gamma_{T,p^{*}}}+2\sqrt{CT\beta_{T}\gamma_{T}|P|}.$ (33)
A.2 HP-GP-TS
First, we state and prove Theorem 4.1. See 4.1
Proof.
First, note that $I_{p}({\mathbf{y}}_{T};f)=H({\mathbf{y}}_{T}|p)-H({\mathbf{y}}_{T}|f,p)$. As in Srinivas et al. (2012), $H({\mathbf{y}}_{T}|f,p)$ is simply the entropy of $T$ zero-mean Gaussians with variance $\sigma^{2}$: $\frac{T}{2}\log(2\pi e\sigma^{2})$. Using the chain rule of entropy, $H({\mathbf{y}}_{T}|p)=H({\mathbf{y}}_{T-1}|p)+H(y_{T}|{\mathbf{y}}_{T-1},p)=H({\mathbf{y}}_{T-1}|p)+\frac{1}{2}\log(2\pi e(\sigma^{2}+\sigma^{2}_{t}(x_{T}))$ where the second equality is due to $y_{T}|{\mathbf{y}}_{T-1},p\sim{\mathcal{N}}(\mu_{T,p}(x_{T}),\sigma^{2}_{T,p}(x_{T})+\sigma^{2})$. The rest follows by induction.
The second part follows from the proof of Lemma 5.4 in Srinivas et al. (2012):
$\sigma_{t,p}^{2}(x_{t})$ $=\sigma^{2}(\sigma^{-2}\sigma^{2}_{t,p}(x_{t}))$ (34)
$\leq\sigma^{2}\frac{\sigma^{-2}}{\log(1+\sigma^{-2})}\log(1+\sigma^{-2}\sigma^{2}_{t,p}(x_{t}))$ (35)
$\leq\frac{1}{\log(1+\sigma^{-2})}\log(1+\sigma^{-2}\sigma^{2}_{t,p}(x_{t})).$ (36)
Using Eq. 8, we get the desired result. ∎
Based on Theorem 4.1, Theorem 4.1 follows almost immediately.
See 4.1
Proof.
From Theorem 4.1, we have for any fixed sequence of arms $x_{1},\ldots,x_{T}\in{\mathcal{X}}$ and prior $p^{*}\in P$, that ${\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}\leq CI_{p^{*}}({\mathbf{y}}_{T};f)$. The mutual information $I_{p^{*}}({\mathbf{y}}_{T};f)$ given the prior $p^{*}$ is in turn bounded by the maximal information gain $\gamma_{T,p^{*}}$ for the same prior. Hence,
${\mathbb{E}}\left[{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})\right]\leq{\mathbb{E}}\left[C\gamma_{T,p^{*}}\right]={\mathbb{E}}_{p^{*}\sim P_{0}}\left[C\gamma_{T,p^{*}}\right].$ (37)
Then, we are ready to state and prove our regret bound for HP-GP-TS.
See 4.2
Proof.
[…]
${\text{BR}}(T)$ $={\sum_{t\in[T]}}{\mathbb{E}}\left[f(x^{*})-f(x_{t})\right]$ (38)
[…]
$=\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}\left[f(x^{*})-U_{t,p^{*}}(x^{*})\right]}_{(1)}+\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p_{t}}(x_{t})-f(x_{t})\right]}_{(2)}.$ (40)
We begin by bounding term $(1)$,
$(1)$ $={\sum_{t\in[T]}}{\mathbb{E}}\left[f(x^{*})-\mu_{t,p^{*}}(x^{*})-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})\right]$ (41)
$\leq{\sum_{t\in[T]}}{\mathbb{E}}\left[\left[f(x^{*})-\mu_{t,p^{*}}(x^{*})-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})\right]_{+}\right]$ $\left([\cdot]_{+}:=\max(\cdot,0)\right)$ (42)
$\leq{\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}{\mathbb{E}}\left[\left[f(x)-\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)\right]_{+}\right]$ $(x^{*}\in{\mathcal{X}},[\cdot]_{+}\geq 0)$ (43)
[…]
Recall that for $Z\sim{\mathcal{N}}(\mu,\sigma)$ with $\mu\leq 0$, ${\mathbb{E}}[[Z]_{+}]=\frac{\sigma}{\sqrt{2\pi}}\exp\left(\frac{-\mu^{2}}{2\sigma^{2}}\right)$. In our case, note that $f|p^{*},H_{t}\sim{\mathcal{N}}(\mu_{t,p^{*}},k_{t,p^{*}})$ and $\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)$ is deterministic given $p^{*},H_{t}$. Hence,
$(1)$ $\leq{\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}{\mathbb{E}}_{p^{*},H_{t}}\left[\frac{\sigma_{t,p^{*}(x)}}{\sqrt{2\pi}}\exp\left(\frac{-\beta_{t}}{2}\right)\right]$ (45)
$\leq{\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}{\mathbb{E}}_{p^{*},H_{t}}\left[\frac{1}{\sqrt{2\pi}}\exp\left(-\frac{\beta_{t}}{2}\right)\right]$ $(\sigma_{t,p^{*}}(x)\leq\sigma_{0,p^{*}}(x)\leq 1)$ (46)
$={\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}\frac{1}{\sqrt{2\pi}}\exp(-\beta_{t}/2)$ (47)
$\leq{\sum_{t\in[T]}}\frac{1}{t^{2}}\leq\frac{\pi^{2}}{6}.$ $(\beta_{t}=2\log(|{\mathcal{X}}|t^{2}/\sqrt{2\pi}))$ (48)
Next, we bound $(2)$ as follows:
$(2)$ $={\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p_{t}}(x_{t})-f(x_{t})\right]$ (49)
[…]
$={\sum_{t\in[T]}}{\mathbb{E}}_{H_{t}}\left[{\mathbb{E}}_{t}\left[U_{t,p_{t}}(x_{t})-\mu_{t,p_{t}}(x_{t})|H_{t}\right]\right]$ $({\mathbb{E}}[f(x_{t})|H_{t}]={\mathbb{E}}[\mu_{t,p_{t}}(x_{t})|H_{t}])$ (51)
$={\sum_{t\in[T]}}{\mathbb{E}}_{H_{t}}[{\mathbb{E}}_{t}[\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})|H_{t}]]$ $(U_{t,p_{t}}(\cdot)=\mu_{t,p_{t}}(\cdot)+\sqrt{\beta_{t}}\sigma_{t,p_{t}}(\cdot))$ (52)
[…]
$={\sum_{t\in[T]}}{\mathbb{E}}[\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x_{t})]$ (54)
Continuing,
$(2)$ $={\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x_{t})\right]$ (55)
[…]
$\leq\sqrt{{\sum_{t\in[T]}}\beta_{t}}{\mathbb{E}}\left[\sqrt{{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})}\right]$ $(\beta_{t}\text{ deterministic})$ (57)
[…]
Combining the bounds for (1) and (2), we obtain
${\text{BR}}(T)\leq\frac{\pi^{2}}{6}+\sqrt{CT\beta_{t}{\mathbb{E}}_{p^{*}}[\gamma_{T,p^{*}}]}.$ (60)
Appendix B Additional experimental details
In this section, we provide some additional experimental details and results. First, we include the full set of confusion matrices for the synthetic lengthscale and hill experiments in Fig. 10. In the lengthscale experiments, we observe that PE-GP-UCB and -TS oversample the shortest lengthscale. This is similar to the kernel experiment where the Matérn 3/2 kernel was also oversampled. However, we see that HP-GP-TS and MAP GP-TS do not suffer from this optimistic bias. In the hill experiment, we see that PE-GP-UCB almost never selects Hill 0, i.e. the mean function with only smaller hills. PE-GP-TS selects the prior almost uniformly whilst HP-GP-TS and MAP-GP-TS select the correct mean function in a majority of the time steps.
In Fig. 11, we visualize the median cumulative regret on the real-world data experiments.
In the Intel experiment, we removed one outlier seed. All methods had a final cumulative regret around 6000°C, note that the average for the worst performing model across the other seeds was $\approx 150$°C. The outlier is shown Fig. 12. We can see that one of the sensors display very high temperatures compared to all other sensors, which is why all methods performed poorly on this seed. It should be noted that many of the sensors in the Intel data logged degrees above 100°C after a certain time - likely due to sensor failure rather boiling temperatures in an office environment. Also note that these days were excluded from both our training and test data. The outlier could be indication that this particular sensor was starting to fail earlier than others.
(a) Lengthscale experiment
(b) Hill experiment
Figure 10: Confusion matrices for the true prior $p^{*}$ and $p_{t}$ across all time steps of the synthetic lengthscale and hill experiments.
Figure 11: Median cumulative regret on Intel temperature data (top) and PeMS data (bottom). Errorbars correspond to first and last decile.
Figure 12: Removed sample from the test data in the Intel experiment. One of the sensors display very high temperatures but

### v2

@@
 Abstract
-Gaussian process (GP) bandits provide a powerful framework for solving blackbox optimization of unknown functions.
-The characteristics of the unknown function depends heavily on the assumed GP prior.
+Gaussian process (GP) bandits provide a powerful framework for performing blackbox optimization of unknown functions.
+The characteristics of the unknown function depend heavily on the assumed GP prior.
 Most work in the literature assume that this prior is known but in practice this seldom holds.
@@
 In this work, we propose two algorithms for joint prior selection and regret minimization in GP bandits based on GP Thompson sampling (GP-TS): Prior-Elimination GP-TS (PE-GP-TS) and HyperPrior GP-TS (HP-GP-TS).
-We theoretically analyze the algorithms and establish upper bounds for their respective regret.
+We theoretically analyze the algorithms and establish upper bounds for the regret of HP-GP-TS.
 In addition, we demonstrate the effectiveness of our algorithms compared to the alternatives through experiments with synthetic and real-world data.
@@
 In Gaussian process bandits, we consider a variant of the multi-armed bandit problem where the arms are correlated and their expected reward is sampled from a Gaussian process (GP).
-The flexibility of GPs have made GP bandits applicable in a wide range of areas that need to optimize blackbox functions with noisy estimates, including machine learning hyperparameter tuning Turner et al.
-(2021), drug discovery Hernández-Lobato et al.
-(2017); Pyzer-Knapp (2018), chemical design Griffiths & Miguel Hernández-Lobato (2020), battery charging protocols Jiang et al.
-(2022), online advertising Nuara et al.
-(2018), and portfolio optimization Gonzalvez et al.
-(2019).
-However, most of the theoretical results in the literature assume that the GP prior is known but this is seldom the case in practical applications.
+The flexibility of GPs have made GP bandits applicable in a wide range of areas that need to optimize blackbox functions with noisy estimates, including machine learning hyperparameter tuning (Turner et al., 2021), drug discovery (Hernández-Lobato et al., 2017; Pyzer-Knapp, 2018), online advertising (Nuara et al., 2018), portfolio optimization (Gonzalvez et al., 2019) and energy-efficient navigation (Sandberg et al., 2025).
+Most of the theoretical results in the literature assume that the GP prior is known but this is seldom the case in practical applications.
 Even with expert domain knowledge, selecting the exact prior to use can be a difficult task.
@@
 (2019); Ziomek et al.
-(2024a) provided algorithms with theoretical guarantees when the kernel lengthscale is unknown.
+(2024) provided algorithms with theoretical guarantees when the kernel lengthscale is unknown.
 More recently, Ziomek et al.
-(2024b) introduced an elimination-based algorithm with theoretical guarantees for an arbitrary set of discrete priors.
+(2025) introduced an elimination-based algorithm with theoretical guarantees for an arbitrary set of discrete priors.
 Their algorithm, Prior-Elimination GP-UCB (PE-GP-UCB), selects the arm and prior which provide the most optimistic upper confidence bound (UCB).
@@
 The first algorithm, Prior-Elimination GP-TS (PE-GP-TS), is an extension of PE-GP-UCB that replaces the doubly optimistic selection rule with posterior sampling and one less layer of optimism.
-We obtain a regret bound for PE-GP-TS of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma}_{T}})$ where $T$ is the horizon and $\hat{\gamma}_{T}$ is the worst-case maximum information gain, which matches that of PE-GP-UCB.
+We analyze the regret of PE-GP-TS.
+For the terms we can bound, we obtain a regret bound for PE-GP-TS of order ${\mathcal{O}}(\sqrt{T\log T|P|\hat{\gamma}_{T}})$ where $T$ is the horizon, $|P|$ is the number of priors and $\hat{\gamma}_{T}$ is the worst-case maximum information gain, which matches that of PE-GP-UCB.
 The second algorithm, HyperPrior GP-TS (HP-GP-TS), uses bi-level posterior sampling to efficiently explore the priors and arms.
-We derive a regret bound of order ${\mathcal{O}}(\sqrt{T\beta_{T}\bar{\gamma}_{T}})$ where $\bar{\gamma}_{T}$ is the maximum information gain averaged across a hyperprior.
-The obtained bound matches that of standard GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ whenever the hyperprior is deterministic.
-Finally, we evaluate our methods on three sets of synthetic experiments and two experiments with real-world data.
+An UCB-based analysis of HP-GP-TS yields a regret bound of order ${\mathcal{O}}(\sqrt{T\log T\bar{\gamma}_{T}})$ (where $\bar{\gamma}_{T}$ is the average maximum information gain) plus a term that corresponds to the cost of learning the optimal prior.
+An information-theoretic analysis yields a regret bound of order ${\mathcal{O}}(\sqrt{T|{\mathcal{X}}|\log|{\mathcal{X}}|})$ where $|{\mathcal{X}}|$ is the number of arms.
+We evaluate our methods on three sets of synthetic experiments and three experiments with real-world data.
 Across the experiments, our Thompson sampling based methods outperform PE-GP-UCB.
-Additionally, we analyze the priors selected by the algorithms and observe that HP-GP-TS selects the correct prior more often than the other algorithms.
+Additionally, we find that the regret of HP-GP-TS does not increase with $|P|$ in our experiments.
+Finally, we analyze the priors selected by the algorithms and observe that HP-GP-TS selects the correct prior more often than the other algorithms.
 The contributions of this work can be summarized as:
   • We propose two novel algorithms for GP-bandits with unknown prior: PE- and HP-GP-TS.
-  • We provide theoretical regret bounds of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma_{T}}})$ and ${\mathcal{O}}(\sqrt{T\beta_{T}\bar{\gamma_{T}}})$ for PE- and HP-GP-TS, respectively.
-  • We experimentally evaluate our algorithms on both synthetic and real-world data, demonstrating that they achieve superior performance by avoiding the optimistic exploration of PE-GP-UCB.
+  • We theoretically analyze the regret of HP-GP-TS using a UCB framework and an information-theoretic framework which provides a regret bound of order ${\mathcal{O}}(\sqrt{T|{\mathcal{X}}|\log|{\mathcal{X}}|})$.
+Additionally, we analyze the regret of PE-GP-TS.
+  • We experimentally evaluate our algorithms on both synthetic and real-world data, demonstrating that they achieve superior performance and that the regret of HP-GP-TS does not increase with $|P|$.
 2 Background and problem statement
@@
 The unknown function $f$ is sampled from a periodic prior and the observed data is in the top left.
-The agent discards the priors on the right since they do not match the periodic observations.
+The agent discards the priors on the right since they do not match the new observation (star).
 Problem statement
@@
 An effective agent must distinguish which prior the means are sampled from to ensure it explores efficiently.
-We illustrate this in Fig.
-1.
+We illustrate this in Fig.˜1.
 Now, let us formally state the problem.
@@
 In the Bayesian selection setting, the prior is sampled from a hyperprior $p^{*}\sim\mathcal{P}(P)$.
-To simplify notation, let $P_{0}$ denote the hyperprior.
+To simplify notation, let $P_{1}$ denote the hyperprior.
 Let $T$ denote the horizon.
-For time step $t=1,2,\ldots,T$, the agent selects an arm $x_{t}\in{\mathcal{X}}$ and observes the reward $r_{t}=f(x_{t})+\epsilon_{t}$ where $\{\epsilon_{t}\}_{t=1}^{T}$ are i.i.d.
+For time step $t=1,2,\ldots,T$, the agent selects an arm $x_{t}\in{\mathcal{X}}$ and observes the reward $y_{t}=f(x_{t})+\epsilon_{t}$ where $\{\epsilon_{t}\}_{t=1}^{T}$ are i.i.d.
 zero-mean Gaussian noise with variance $\sigma^{2}$.
-The goal of the agent is to select a sequence of arms $\{x_{t}\}_{t=1}^{T}$ that minimizes the regret $R(T)=\sum_{t\in[T]}f(x^{*})-f(x_{t})$ where $[T]=\{1,\ldots,T\}$ and $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}f(x)$.
-In the Bayesian selection setting, we evaluate the agent based on the Bayesian regret ${\text{BR}}(T)={\mathbb{E}}\left[R(T)\right]$ where the expectation is taken over the prior $p^{*}$, the expected reward function $f$, the noise $\{\epsilon_{t}\}_{t=1}^{T}$ and the potentially stochastic selection of arms.
+The goal of the agent is to select a sequence of arms $\{x_{t}\}_{t=1}^{T}$ that minimizes the regret $R(T)=\sum_{t\in[T]}f(x^{*})-f(x_{t})$ where $T]=\{1,\ldots,T\}$ and $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}f(x)$.
+In the Bayesian selection setting, we evaluate the agent based on the Bayesian regret ${\text{BR}}(T)={\mathbb{E}}\left[R(T)\right]$ where the expectation is taken over the prior $p^{*}$, the expected reward function $f$, the noise $\{\epsilon_{t}\}_{t=1}^{T}$ and the (potentially) stochastic selection of arms.
 Gaussian processes
-A Gaussian process $f(x)\sim{\mathcal{GP}}(\mu,k)$ is a collection of random variables such that for any subset $\{x_{1},\ldots,x_{n}\}\subset{\mathcal{X}}$, the vector $[f(x_{1}),\ldots f(x_{n})]\in{\mathbb{R}}^{n}$ has a multivariate Gaussian distribution.
-The probabilistic nature of GPs makes them very useful for defining and solving bandit problems where the arms are correlated.
-Given $t$ observations $\{(x_{i},y_{i})\}_{i=1}^{t}$, the posterior mean and kernel functions of a Gaussian process ${\mathcal{GP}}(\mu,k)$ are given by
+A Gaussian process $f(x)\sim{\mathcal{GP}}(\mu,k)$ is a collection of random variables such that for any subset $\{x_{1},\ldots,x_{n}\}\subset{\mathcal{X}}$, the vector $f(x_{1}),\ldots,f(x_{n})]\in{\mathbb{R}}^{n}$ has a multivariate Gaussian distribution.
+The probabilistic nature of GPs make them very useful for defining and solving bandit problems where the arms are correlated.
+Given the history $H_{t}=\{(x_{i},y_{i})\}_{i=1}^{t-1}$, the posterior mean and kernel functions of a Gaussian process ${\mathcal{GP}}(\mu,k)$ are given by
 $\mu_{t}(x)$ $=\mu(x)+{\mathbf{k}}^{\top}\left({\mathbf{K}}+\sigma^{2}I\right)^{-1}({\mathbf{y}}-{\bm{\mu}}),$ (1)
-$k_{t}(x,\tilde{x})$ $=k(x,\tilde{x})-{\mathbf{k}}^{\top}\left({\mathbf{K}}+\sigma^{2}I\right)^{-1}\tilde{{\mathbf{k}}}.$ (2)
-Above, ${\mathbf{k}},\tilde{\mathbf{k}}\in{\mathbb{R}}^{t}$ are vectors such that $({\mathbf{k}})_{i}=k(x_{i},x)$ and $(\tilde{{\mathbf{k}}})_{i}=k(x_{i},\tilde{x})$.
-Additionally, ${\mathbf{y}},{\bm{\mu}}\in{\mathbb{R}}^{t}$ are also vectors such that $({\mathbf{y}})_{i}=y_{i}$ and $({\bm{\mu}})_{i}=\mu(x_{i})$.
-The gram matrix is denoted by ${\mathbf{K}}\in{\mathbb{R}}^{t\times t}$ where $({\mathbf{K}})_{i,j}=k(x_{i},x_{j})$.
-Let $\mu_{t,p}$ and $k_{t,p}$ denote the posterior mean and kernel for a Gaussian process with prior $p\in P$ at time $t$ and let $\sigma^{2}_{t,p}(x)=k_{t,p}(x,x)$ denote the posterior variance at time $t$.
-The kernel $k$ determines important characteristics of the functions $f$ and in the following we provide some examples.
-The RBF kernel, $k(x,\tilde{x})=\exp(-||x-\tilde{x}||^{2}/\ell^{2})$ guarantees that $f$ is smooth.
-The lengthscale parameter $\ell>0$ determines how quickly $f$ changes, smaller values lead to more fluctuations.
-The rational quadratic (RQ) kernel $k(x,\tilde{x})=\left(1+\frac{||x-\tilde{x}||^{2}}{2\alpha\ell^{2}}\right)^{-\alpha}$ where $\alpha>0$ is a mixture of RBF kernels with varying lengthscales.
-The Matérn kernel Matérn (1986) $k(x,\tilde{x})=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}||x-\tilde{x}||}{\ell}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}||x-\tilde{x}||}{\ell}\right)$ where $\nu>0$ is the smoothness parameter that imposes that $f$ is k-times differentiable if $\nu>k$ for integer $k$.
-The functions $\Gamma(\nu)$ and $K_{v}$ correspond to the gamma function and a modified Bessel function Williams & Rasmussen (2006).
-The periodic kernel $k(x,\tilde{x})=\exp\left(-\frac{1}{2}\sum_{i=1}^{d}\sin^{2}(\frac{\pi}{\rho}(x-\tilde{x}))/\ell\right)$ generates smooth and periodic functions with period $\rho>0$ Mackay (1998).
-The linear kernel $k(x,\tilde{x})=vx^{\top}\tilde{x}/\nu$ generates linear functions where $v$ is the variance parameter.
+[…]
+[…]
 Information gain
 The maximal information gain (MIG) is a measure of reduction in uncertainty of $f$ after observing the most informative data points up to a specified size.
-The MIG commonly occurs in regret bounds for GP bandit algorithms Srinivas et al.
-(2012); Vakili et al.
-(2021) and its growth rate is strongly determined by the prior kernel of the GP.
+The MIG commonly occurs in regret bounds for GP bandit algorithms (Srinivas et al., 2012; Vakili et al., 2021) and its growth rate is strongly determined by the prior kernel of the GP.
 Hence, we will define the MIG for any fixed GP prior $p\in P$.
@@
 where $I_{p}({\mathbf{y}}_{A};f)=H({\mathbf{y}}_{A}|p)-H({\mathbf{y}}_{A}|f,p)$ is the mutual information between ${\mathbf{y}}_{A}$ and $f$ given $p$, and $H(\cdot)$ denotes the entropy.
-To aid our analysis later, we also define the worst-case MIG as $\hat{\gamma}_{T}:=\max_{p\in P}\gamma_{T,p}$ and the average MIG as $\bar{\gamma}_{T}:={\mathbb{E}}_{p\sim P_{0}}[\gamma_{T,p}]$.
-For the RBF and Matérn kernels, $\gamma_{T,p}={\mathcal{O}}(\log^{d+1}(T))$ and $\gamma_{T,p}={\mathcal{O}}(T^{\frac{d}{2\nu+d}}\log^{\frac{2\nu}{2\nu+d}}(T))$ Srinivas et al.
-(2012); Vakili et al.
-(2021).
+To aid our analysis later, we also define the worst-case MIG as $\hat{\gamma}_{T}:=\max_{p\in P}\gamma_{T,p}$ and the average MIG as $\bar{\gamma}_{T}:={\mathbb{E}}_{p\sim P_{1}}[\gamma_{T,p}]$.
+For the RBF and Matérn kernels, $\gamma_{T,p}={\mathcal{O}}(\log^{d+1}(T))$ and $\gamma_{T,p}={\mathcal{O}}(T^{\frac{d}{2\nu+d}}\log^{\frac{2\nu}{2\nu+d}}(T))$ (Srinivas et al., 2012; Vakili et al., 2021).
 Previous work
-Wang & de Freitas (2014) was the first work that presented regret bounds for Gaussian process bandits with unknown lengthscale.
-Wang & de Freitas studied the popular Expected Improvement algorithm Močkus (1975).
-But their algorithm requires a lower bound on the lengthscale and their regret bound depends on the worst-case MIG.
+Plenty of previous work have proposed fully Bayesian approaches that integrate the acquisition function over the hyperposterior (Osborne et al., 2009; Benassi et al., 2011; Snoek et al., 2012; Hernández-Lobato et al., 2014; Wang & Jegelka, 2017; De Ath et al., 2021; Hvarfner et al., 2023).
+In contrast, HP-GP-TS optimizes a single hyperposterior sample instead of expected values over the hyperposterior.
+Wang & de Freitas (2014) first derived regret bounds for GP bandits with unknown lengthscale for the Expected Improvement algorithm (Močkus, 1975).
+However, the proposed algorithm requires a lower bound on the lengthscale and the regret bound depends on the worst-case MIG.
 Later work by Berkenkamp et al.
-(2019) introduced Adaptive GP-UCB (A-GP-UCB) that continually lowers lengthscale parameter.
-Once the lengthscale parameter is sufficently small, then the function $f$ will lie within the reproducing kernel Hilbert space (RKHS) and then the regular GP-UCB theory can be applied.
+(2019) introduced Adaptive GP-UCB (A-GP-UCB) that continually lowers the lengthscale parameter.
+Given a sufficiently small lengthscale, the function $f$ lies within the reproducing kernel Hilbert space (RKHS) and the regular GP-UCB theory can be applied.
 However, A-GP-UCB lacks a stopping mechanism and will overexplore as the lengthscale continues to shrink.
 Recent work by Ziomek et al.
-(2024b) introduced Prior-Elimination GP-UCB (PE-GP-UCB) for time-varying GP-bandits with unknown prior.
+(2025) introduced Prior-Elimination GP-UCB (PE-GP-UCB) for time-varying GP-bandits with unknown prior.
 Unlike the work before, the regret bound of PE-GP-UCB holds for arbitrary types of hyperparameters in the GP prior.
@@
 PE-GP-UCB tracks the cumulative prediction error made by the selected priors and eliminates priors that exceed a threshold level.
-Other works have introduced regret balancing algorithms that maintain a set of base learning algorithms and balance their selection frequency to achieve close to optimal regret Abbasi-Yadkori et al.
-(2020); Pacchiano et al.
-(2020).
+Other works have introduced regret balancing algorithms that maintain a set of base learning algorithms and balance their selection frequency to achieve close to optimal regret (Abbasi-Yadkori et al., 2020; Pacchiano et al., 2020).
 Ziomek et al.
-(2024a) built on this idea and introduced length-scale balancing GP-UCB which can adaptively explore smaller lengthscales but can return to longer ones, unlike A-GP-UCB.
-Wang et al.
-(2018) provided regret bounds for probability of improvement Kushner (1964) and GP-UCB when the GP prior is learnt from offline data, which they called Meta BO.
+(2024) built on this idea and introduced length-scale balancing GP-UCB which can adaptively explore smaller lengthscales but can return to longer ones, unlike A-GP-UCB.
 The aforementioned works are based on UCB, EI, PI or regret balancing.
-However, another line of work has studied Thompson sampling in standard and linear bandits with unknown prior distribution Kveton et al.
-(2021); Basu et al.
-(2021); Hong et al.
-(2022); Li et al.
-(2024).
+However, another line of work has studied Thompson sampling in standard and linear bandits with unknown prior distribution (Kveton et al., 2021; Basu et al., 2021; Hong et al., 2022; Li et al., 2024).
 In their setting (meta or hierarchical bandits), the agent plays multiple bandit instances, either simultaneously or sequentially.
@@
 We emphasize that these methods have been studied only for standard stochastic and linear bandits, not for GP bandits.
+Figure 2: a) Elimination procedure of PE-GP-TS.
+The solid lines correspond to posterior means and the shaded regions are confidence intervals.
+The figure has been adapted from Ziomek et al.
+(2025).
+The dashed lines are samples from the posteriors.
+b) Overview of HP-GP-TS.
+The orange star corresponds to $y_{t}$.
 3 Algorithms
-As discussed by Russo & Roy (2014), TS can offer advantages over UCB algorithms for problems where constructing tight confidence bounds is difficult.
-In addition, Thompson sampling is often observed to perform better than UCB in practice Chapelle & Li (2011); Wen et al.
-(2015); Kandasamy et al.
-(2018); Åkerblom et al.
-(2023b, a).
+As discussed by Russo & Van Roy (2014), TS can offer advantages over UCB algorithms for problems where constructing tight confidence bounds is difficult.
+In addition, Thompson sampling is often observed to perform better than UCB in practice (Chapelle & Li, 2011; Wen et al., 2015; Kandasamy et al., 2018; Åkerblom et al., 2023b, a).
 Motivated by this, we present two algorithms for efficient prior selection based on TS.
 3.1 Prior-Elimination with Thompson sampling
-Figure 2: Elimination procedure of PE-GP-TS.
-The solid lines correspond to posterior means and the shaded regions are confidence intervals.
-The dashed lines are samples from the posteriors.
-The figure has been adapted from Ziomek et al.
-(2024b) to illustrate the elimination step of PE-GP-TS.
 Algorithm 1 Prior Elimination GP-TS (PE-GP-TS)
 0: Horizon $T$, prior functions $\{\mu_{1,p},k_{1,p}\}_{p\in P}$, confidence parameters $\{\beta_{t}\}_{t=1}^{T}$ and $\{\xi_{t}\}_{t=1}^{T}$.
-1:  $P_{0}=P$, $S_{0,p}=\varnothing$ $\forall p\in P$
+[…]
 2:for $t=1,2\ldots,T$ do
-3:     Sample $\tilde{f}_{t,p}\sim{\mathcal{GP}}(\mu_{t,p},k_{t,p})$ $\forall p\in P_{t-1}$
+3:  Sample $\tilde{f}_{t,p}\sim{\mathcal{GP}}(\mu_{t,p},k_{t,p})$ $\forall p\in P_{t}$
 4:  Set $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t-1}}\tilde{f}_{t,p}(x)$
@@
 9:  if $\left|\sum_{i\in S_{t,p_{t}}}\eta_{i}\right|>V_{t}$ then
-10:        $P_{t}=P_{t-1}\setminus\{p_{t}\}$
+10:   $P_{t+1}=P_{t}\setminus\{p_{t}\}$
 11:  else
-12:        $P_{t}=P_{t-1}$
-Our first algorithm is an extension of the PE-GP-UCB algorithm Ziomek et al.
-(2024b) to be employed with Thompson sampling - instead of UCB.
-The key difference is that instead of maximizing the upper confidence bound $U_{t}(x,p)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$ over ${\mathcal{X}}\times P$, we instead sample $\tilde{f}_{t,p}$ from the posterior ${\mathcal{GP}}(\mu_{t,p},k_{t,p})$ for all priors $p\in P_{t-1}$ where $P_{t-1}$ is the set of active priors.
+12:   $P_{t+1}=P_{t}$
+Our first algorithm is an extension of PE-GP-UCB (Ziomek et al., 2025) to be employed with Thompson sampling - instead of UCB.
+The key difference is that instead of maximizing the upper confidence bound $U_{t}(x,p)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$ over ${\mathcal{X}}\times P_{t-1}$, we instead sample $\tilde{f}_{t,p}$ from the posterior ${\mathcal{GP}}(\mu_{t,p},k_{t,p})$ for all priors $p\in P_{t-1}$ where $P_{t-1}$ is the set of active priors.
 Then, we select the arm and prior $x_{t},p_{t}$ such that $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t-1}}\tilde{f}_{t,p}(x)$.
-Whilst PE-GP-UCB has two layers of optimism, the upper confidence bound and joint maximization of $x$ and $p$.
-PE-GP-TS has only a single layer of optimism - which should alleviate potential overexploration issues.
-The elimination procedure of PE-GP-TS is illustrated in Fig.
-2.
+Whilst PE-GP-UCB has two layers of optimism, the upper confidence bound and joint maximization of $x$ and $p$, PE-GP-TS has only a single layer of optimism - which should alleviate potential overexploration issues.
+The elimination procedure of PE-GP-TS is illustrated in Fig.˜2.
 Samples $\tilde{f}_{t,p}$ are drawn from the active prior $p\in P_{t-1}$.
@@
 Otherwise, it remains active.
-The PE-GP-TS algorithm is presented in Algorithm 1.
+The PE-GP-TS algorithm is presented in Algorithm˜1.
 Similar to PE-GP-UCB, the set $S_{t,p}$ is used to store the time steps where prior $p$ was selected up to and including time $t$.
 When prior $p_{t}$ is selected, the prediction error $\eta_{t}=y_{t}-\mu_{t,p_{t}}(x_{t})$ between the observed and predicted value made by the prior $p_{t}$ is computed.
-If the sum of prediction errors made by the prior $p_{t}$ exceeds the threshold value $V_{t}$, then $p_{t}$ is eliminated from the active priors $P_{t}$, see line 9.
+If the sum of prediction errors made by the prior $p_{t}$ exceeds the threshold value $V_{t}$, then $p_{t}$ is eliminated from the active priors $P_{t}$, see line˜9.
 Note that at time step $t$, only the selected prior $p_{t}$ can be eliminated.
@@
 Algorithm 2 HyperPrior GP-TS (HP-GP-TS)
-0:  Horizon $T$, prior functions $\{\mu_{1,p},k_{1,p}\}_{p\in P}$, hyperprior $P_{0}$.
+0: Horizon $T$, prior functions $\{\mu_{1,p},k_{1,p}\}_{p\in P}$, hyperprior $P_{1}$.
 1:for $t=1,2\ldots,T$ do
-2:     Sample $p_{t}\sim P_{t-1}$
+2:  Sample $p_{t}\sim P_{t}$
 3:  Sample $\tilde{f}_{t}\sim{\mathcal{GP}}(\mu_{t,p_{t}},k_{t,p_{t}})$
@@
 5:  Observe $y_{t}=f(x_{t})+\epsilon_{t}$
-6:     Set ${P}_{t}(p)\propto{\mathbb{P}}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)P_{t-1}(p)$triangleright Update hyperposterior
+6:  Set ${P}_{t+1}(p)\propto\mathbb{P}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)\cdot P_{t}(p)$ $\triangleright$ Update hyperposterior
 In our first algorithm, we removed one layer of optimism.
@@
 By shedding the optimism over the selected prior $p_{t}$, HP-GP-TS should be able avoid costly exploration by selecting likely priors instead of optimistic ones.
-The algorithm is presented in Algorithm 2.
+The algorithm is visualized in Fig.˜2 and presented with more details in Algorithm˜2.
 In the first step, the current prior $p_{t}$ is sampled from the hyperposterior $P_{t-1}$.
@@
 Note that since the set of priors $P$ is finite, computing the posterior is tractable albeit computationally costly with a complexity of ${\mathcal{O}}(t^{3}|P|)$.
-The likelihood ${\mathbb{P}}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)={\mathcal{N}}(y_{t};\mu_{t,p}(x_{t}),\sigma_{t,p}^{2}(x_{t})+\sigma^{2})$ is simply the Gaussian likelihood of the posterior at $x_{t}$ plus Gaussian noise with variance $\sigma^{2}$.
-4 Regret bounds
-In this section, we present regret bounds for the proposed algorithms.
+The likelihood $\mathbb{P}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)={\mathcal{N}}(y_{t};\mu_{t,p}(x_{t}),\sigma_{t,p}^{2}(x_{t})+\sigma^{2})$ is simply the Gaussian likelihood of the posterior at $x_{t}$ with added Gaussian noise with variance $\sigma^{2}$.
+4 Regret analysis
+In this section, we analyze the regret for the proposed algorithms.
 Recall from the problem statement that we consider two slightly different settings for the two algorithms.
-Specifically, for PE-GP-TS we assume the unknown prior $p^{*}$ is sampled arbitrarily from $P$ whilst for HP-GP-TS we assume that the unknown prior $p^{*}$ is selected from a known hyperprior distribution $P_{0}$.
+Specifically, for PE-GP-TS we assume the unknown prior $p^{*}$ is selected arbitrarily from $P$ whilst for HP-GP-TS we assume that the unknown prior $p^{*}$ is selected from a known hyperprior distribution $P_{1}$.
+4.1 Analysis of PE-GP-TS
 Ziomek et al.
-(2024b) structured the proof of the regret bound of PE-GP-UCB into 4 larger steps; First, showing that $p^{*}$ is never eliminated with high probability.
+(2025) structured the proof of the regret bound of PE-GP-UCB into 4 larger steps; First, showing that $p^{*}$ is never eliminated with high probability.
 Second, establishing a bound on the simple regret.
@@
 Finally, the cumulative bound is re-expressed in terms of the worst-case MIG.
-For PE-GP-TS, we note that it is sufficient to establish a new bound on the simple regret and then adapt the steps of Ziomek et al.
+For PE-GP-TS, we establish a new bound on the simple regret and then adapt the steps of Ziomek et al.
 to accommodate the new simple regret bound.
@@
 {lemma}
-[] For any $p^{*}\in P$, if $f(x)\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ and $\beta_{t}=2\log\left(\frac{|{\mathcal{X}}||P|\pi^{2}t^{2}}{3\delta}\right)$.
-Then, with probability at least $1-\delta$, the following holds for all $t,x\in[T]\times{\mathcal{X}}$:
+[] If $f(x)\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ and $\beta_{t}=2\log\left(\frac{|{\mathcal{X}}||P|\pi^{2}t^{2}}{3\delta}\right)$.
+Then, with probability at least $1-\delta$, the following holds for all $t,x,p\in[T]\times{\mathcal{X}}\times P$:
 $|f(x)-\mu_{t,p^{*}}(x)|$ $\leq\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x),$ (4)
-and for all $t,x,p\in[T]\times{\mathcal{X}}\times P:$
 $|\tilde{f}_{t,p}(x)-\mu_{t,p}(x)|$ $\leq\sqrt{\beta_{t}}\sigma_{t,p}(x).$ (5)
-All proofs can be found in Appendix A.
-Section 4 is based on Lemma 5.1 of Srinivas et al.
-(2012) but adapted to TS by specifying that it holds for any sequence of $x_{1},\ldots,x_{T}$, as discussed by Russo & Roy (2014).
-Additionally, we add Eq.
-5 which can be shown through the same steps and an additional union bound over $P$.
+All proofs can be found in Appendix˜A.
+Section˜4.1 is based on Lemma 5.1 of Srinivas et al.
+(2012) but adapted to TS by specifying that it holds for any sequence of $x_{1},\ldots,x_{T}$, as discussed by Russo & Van Roy (2014).
+Additionally, we add Eq.˜5 which can be shown through the same steps and an additional union bound over $P$.
 Next, we state our bound for the simple regret of PE-GP-TS.
 {lemma}
-[] If the event of Section 4 holds, then the following holds for the simple regret of PE-GP-TS for all $t\in[T]$:
+[] If the event of Section˜4.1 holds, then the following holds for the simple regret of PE-GP-TS for all $t\in[T]$:
 $f(x^{*})-f(x_{t})\leq 2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-\eta_{t}+\epsilon_{t}.$ (6)
-Compared to the simple regret bound for PE-GP-UCB, we obtain the additional term $2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})$.
-Since $p^{*}$ is fixed, the sum over $t$ of the new term is ${\mathcal{O}}(\sqrt{T\beta_{T}\gamma_{T,p^{*}}})$ and we obtain the following regret bound:
+Compared to the simple regret bound for PE-GP-UCB, we obtain the additional term $2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})$ which leads to the following regret bound:
 Theorem 4.1.
+Let $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$ and $C=2/\log(1+\sigma^{-2})$.
 If $p^{*}\in P$ and $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, then PE-GP-TS with confidence parameters $\beta_{t}=2\log(2|{\mathcal{X}}||P|\pi^{2}t^{2}/3\delta)$ and $\xi_{t}=2\sigma^{2}\log(|P|\pi^{2}t^{2}/3\delta)$, satisfies the following regret bound with probability at least $1-\delta$:
-$R(T)\leq$ $2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}$ (7)
-$+2\sqrt{CT\beta_{T}\gamma_{T,p^{*}}}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}$
-where $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$ and $C=2/\log(1+\sigma^{-2})$.
-Since $\gamma_{T,p^{*}}\leq\hat{\gamma}_{T}$, the bound is of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma}_{T}})$ w.r.t.
+$R(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}+2\sqrt{CT\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}$ (7)
+The bound of the first three terms is of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma}_{T}})$ w.r.t.
 $T$ which matches that of PE-GP-UCB.
-To our knowledge, the best lower bound for standard GP bandits in the Bayesian setting, where $f$ is sampled from a GP, is $\Omega(\sqrt{T})$ for $d=1$ Scarlett (2018).
+To our knowledge, the best lower bound for standard GP bandits in the Bayesian setting, where $f$ is sampled from a GP, is $\Omega(\sqrt{T})$ for $d=1$ (Scarlett, 2018).
 This would suggest that our bound is tight up to a factor ${\mathcal{O}}(\sqrt{\beta_{T}\hat{\gamma}_{T}})$.
-[…]
-{lemma}
-[] Let $C=2/\log(1+\sigma^{-2})$ and assume $\sigma_{t,p}(x)\leq 1$, then for any sequence of arms $x_{1},\ldots,x_{T}\in{\mathcal{X}}$ and any $p\in P$
-$I_{p}({\mathbf{y}}_{T};f)=\frac{1}{2}\sum_{t\in[T]}\log\left(1+\sigma^{-2}\sigma_{t,p}^{2}(x_{t})\right),$ (8)
-${\sum_{t\in[T]}}\sigma^{2}_{t,p}(x_{t})\leq CI_{p}({\mathbf{y}}_{t};f).$ (9)
-{corollary}
-[] For any stochastic or deterministic sequence of arms $x_{1},\ldots,x_{T}\in{\mathcal{X}}$ and $p^{*}\sim P_{0}$,
-${\mathbb{E}}\Big{[}{\sum_{t\in[T]}}\sigma_{t,p^{*}}^{2}(x_{t})\Big{]}\leq C\bar{\gamma_{T}}.$ (10)
-The proof of Theorem 4.1 is based on Lemma 5.3 and 5.4 in Srinivas et al.
-(2012) but is adapted to randomized selection by specifying that it holds for any sequence of arms.
-Theorem 4.1 follows by Theorem 4.1 and the definition of maximum information gain.
-Next, we state our regret bound for HP-GP-TS.
+4.2 Analysis of HP-GP-TS
+We analyze the regret of HP-GP-TS in two ways: using the UCB-based framework of Russo & Van Roy (2014) and the information-theoretic framework of Russo & Van Roy (2016).
+First, note that HP-GP-TS inherits the probability matching property of GP-TS that $x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t}$ where $\overset{d}{=}$ denotes equal in distribution.
+In addition, $p_{t}|H_{t}\overset{d}{=}p^{*}|H_{t}$ since $p_{t}$ is sampled from the posterior distribution of $p^{*}$.
+In the UCB-based framework, we decompose the regret into three terms:
+${\mathbb{E}}[f(x^{*})-f(x_{t})]={\mathbb{E}}\Big[\underbrace{f(x^{*})-U_{t,p^{*}}(x^{*})}_{(1)}+\underbrace{U_{t,p^{*}}(x^{*})-U_{t,p^{*}}(x_{t})}_{(2)}+\underbrace{U_{t,p^{*}}(x_{t})-f(x_{t})}_{(3)}\Big]$ (8)
+where $U_{t,p}(x)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$.
+Summing over $t\in[T]$, the first term can be bounded by a constant whilst the third term is bounded by $\sqrt{CT\beta_{T}\bar{\gamma}_{T}}$.
+Together, these two terms match the Bayesian regret bound for GP-TS with known prior.
+For the second term, one can utilize that $p^{*},x_{t}|H_{t}\overset{d}{=}p_{t},x^{*}|H_{t}$ to re-express it as ${\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$.
+Hence, the second term can be seen as the cost of learning the true prior.
+Intuitively, if the priors are similar then the upper-confidence bounds of the true and selected prior will not differ significantly.
+Similarly, if the priors are sufficiently distinguishable, one would expect $P_{t}(p^{*})$ to increase quickly from a few samples.
+In the following lemma, we demonstrate the latter for $|P|=2$ with shared kernel functions.
+{lemma}[] If $|P|=2$ and the two priors share the same kernel function ($k_{p}=k$ $\forall p\in P$), then for any fixed sequence of arms $x_{1:t}=\{x_{i}\}_{i=1}^{t}$ the posterior probability of the true prior $p^{*}$ satisfies
+${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|p^{*}=p,x_{1:t}\right]\geq 1+P_{0}(p)e^{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}\Phi\Big(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\Big)-\frac{1}{P_{0}(p)}\Phi\Big(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\Big)$ (9)
+where ${\bm{\mu}}\in{\mathbb{R}}^{t}$ such that $({\bm{\mu}})_{i}=\mu_{p^{*}}(x_{i})-\mu_{\tilde{p}}(x_{i})$ for $\tilde{p}\neq p^{*}$ and $({\bm{\Sigma}})_{i,j}=k(x_{i},x_{j})$.
+From Eq.˜8, we note that $P_{t}(p^{*})$ increases the most when $\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$ is maximized, which corresponds to selecting arms with low correlation and large difference in prior mean (relative to the variance).
+Hence, $P_{t}(p^{*})$ increases quickly at first when HP-GP-TS explores but levels out when the algorithm starts exploiting.
+Next, we state the Bayesian regret bound for HP-GP-TS.
 Theorem 4.2.
-If $p^{*}\sim P_{0}$ and $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, then the Bayesian regret of HP-GP-TS is bounded by
-${\text{BR}}(T)\leq\frac{\pi^{2}}{6}+\sqrt{C_{1}T\beta_{T}\bar{\gamma}_{T}}$ (11)
-where $C_{1}=2\log(1+\sigma^{-2})$ and $\beta_{t}=2\log(|{\mathcal{X}}|t^{2}/\sqrt{2\pi)}$.
-Unlike PE-GP-TS and PE-GP-UCB, the regret bound for FB-GP-TS depends on the average $\bar{\gamma}_{T}$ rather than the worst case $\hat{\gamma}_{T}$ which can impact the theoretical regret significantly if the complexity of learning the priors differ.
-Although, this is reasonable since the elimination methods assume arbitrary selection of $p^{*}$ as opposed to sampling from a hyperprior.
-If the hyperprior is deterministic then the regret bound for FB-GP-TS matches that of GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ Takeno et al.
-(2024) and the average $\bar{\gamma}_{T}$ would be equal to the worst case $\hat{\gamma}_{T}$.
-Again, using the lower bound of Scarlett (2018), our upper bound would tight up to a factor of ${\mathcal{O}}(\sqrt{\beta_{T}\bar{\gamma}_{T}})$.
-Figure 3: Prior mean functions for the hill experiment. Mean function 0 consists of 10 smaller hills whereas mean functions 1 to 4 have 9 smaller hills and 1 taller hills.
+If $p^{*}\sim P_{1}$, $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ and $\beta_{t}=2\log(|{\mathcal{X}}|t^{2}/\sqrt{2\pi)}$, then the Bayesian regret of HP-GP-TS is bounded by
+${\text{BR}}(T)\leq\pi^{2}/6+{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})\right]+\sqrt{CT\beta_{T}\bar{\gamma}_{T}}.$ (10)
+Unlike PE-GP-TS and PE-GP-UCB, the regret bound of the third term for HP-GP-TS depends on the average MIG $\sqrt{\bar{\gamma}_{T}}$ rather than the worst case $\sqrt{|P|\hat{\gamma}_{T}}$ which can impact the theoretical regret significantly if the complexity of the priors differ and the prior is weighted towards simple priors.
+This is reasonable since the elimination methods assume arbitrary selection of $p^{*}$ as opposed to sampling from a hyperprior.
+If the hyperprior is deterministic then the regret bound for HP-GP-TS matches that of GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ (Takeno et al., 2024) and $\bar{\gamma}_{T}$ would be equal to the worst case $\hat{\gamma}_{T}$.
+Again, using the lower bound of Scarlett (2018), our upper bound would be tight up to a factor of ${\mathcal{O}}(\sqrt{\beta_{T}\bar{\gamma}_{T}})$.
+The information-theoretic framework of Russo & Van Roy (2016) can be applied generally if the probability matching property is satisfied and the rewards are subgaussian (Vashishtha & Maillard, 2025).
+Theorem 4.3.
+If $p^{*}\sim P_{1}$, $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, the Bayesian regret of HP-GP-TS is bounded by
+${\text{BR}}(T)\leq\sqrt{2|{\mathcal{X}}|\log(|{\mathcal{X}}|)(\sigma^{2}_{0}+\sigma^{2})T}.$ (11)
+The proof of Theorem˜4.3 follows the proof in section D.2 of Russo & Van Roy (2016) with subgaussian noise.
+For completeness, we provide a proof using the GP-bandit notation in Algorithm˜2.
+The information-theoretic regret bound is ${\mathcal{O}}(\sqrt{T})$ which improves upon the ${\mathcal{O}}(\sqrt{T\beta_{T}\bar{\gamma}_{T}})$ obtained previously and would match the lower bound of Scarlett (2018) in terms of $T$.
+However, we also obtain a ${\mathcal{O}}(\sqrt{|{\mathcal{X}}|\log|{\mathcal{X}}|})$ dependency.
 Figure 3: Cumulative regret for synthetic experiments with varying kernel (left), lengthscale (center) and mean function (right).
-The average final regret for PE-GP-UCB is 116.5 in the lengthscale experiment.
-Errorbars correspond to ±$\pm 1$ standard error.
-Figure 5: Average amount of priors remaining in $P_{t}$ over time for PE-GP-UCB and -TS.
+The average final regret for PE-GP-UCB is 116.5 and 389.0 in the lengthscale and subspace experiments.
+Errorbars correspond to $\pm 1$ standard error.
+Figure 4: Mean number of priors remaining in $P_{t}$ over time for PE-GP-UCB and -TS (left).
+Mean entropy in the hyperposterior $P_{t}$ over time for HP- and MAP GP-TS (right).
+The dashed reference values correspond to entropies of discrete distributions with prob.
+$q$ on one choice and prob.
+$\frac{1-q}{|P|-1}$ on the other |P|1|P|-1 choices.
 5 Experiments
-In this section, we describe our experiments based on synthetic and real-world data.
 Synthetic experiments
 We consider three synthetic setups with different choices of priors in $P$.
-For the first setup, the priors have varying kernels selected according to one of the following kernels: i) squared-exponential kernel, Matérn kernel with $\nu=5/2$, ii) Matérn kernel with $\nu=3/2$, iii) periodic kernel with period $\rho=5$, iv) linear kernel with $v=0.05^{2}$, and, v) the rational quadratic kernel with $\alpha=0.5$.
+For the first setup, the priors have one of the following kernels: i) RBF kernel, ii) the rational quadratic kernel with $\alpha=0.5$, iii) Matérn kernel with $\nu=5/2$, iv) Matérn kernel with $\nu=3/2$, v) periodic kernel with period $\rho=5$, vi) linear kernel with $v=0.05^{2}$.
 All kernels use a lengthscale of $1.0$ and are scaled s.t.
 $k(x,\tilde{x})\leq 1$.
-In addition, the mean function for all priors are zero everywhere.
-For the second setup, the priors use the squared-exponential kernel with lengthscales $4$, $2$, $1$ or $1/2$.
-The third setup is based on the toy problem in an earlier version of Ziomek et al.
-(2024b) where the mean functions vary.
-The first mean function has 10 smaller ‘hills’ across the first half of the interval.
-The four other mean functions are the same except the height of one of the hills is doubled in height, see Fig.
-3.
-All priors in the third setup use a squared-exponential kernel with lengthscale 1.
-For all of the setups, the true prior $p^{*}$ is sampled uniformly from $P$, 500 arms are equidistantly spaced in $[0,20]$, the noise variance $\sigma^{2}=0.25^{2}$, and the horizon $T=500$.
+In addition, the mean function for all priors is zero everywhere.
+For the second setup, the priors use the RBF kernel with lengthscales $4$, 22, 11 or $1/2$.
+For the third setup, the total dimensions $d=16$ but each prior $p_{i}$ assumes $f(x)$ depends on $d_{s}=4$ subdimensions: $[i,i+1,i+2,i+3]$ for $i\in[5]$.
+Dimensions larger than $5$ are wrapped around 1, i.e.
+$((j-1)\mod 5)+1$, such that the priors are equally difficult to distinguish and optimize.
+All priors use the RBF kernel with lengthscale $\ell=8$.
+For all three setups, the true prior $p^{*}$ is sampled uniformly from $P$, the noise variance $\sigma^{2}=0.25^{2}$, and the horizon $T=500$.
+For the first two setups, 500 arms are equidistantly spaced in $[0,20]$ and for the third 500 arms are sampled uniformly on $0,20]^{16}$.
 The prior elimination methods use $\delta=0.05$.
 All models are evaluated on 500 seeds on each setup.
-As baselines, we use PE-GP-UCB and Maximum A Posteriori (MAP) GP-TS where MAP GP-TS is identical to HP-GP-TS except for greedily selecting $p_{t}$ from the posterior: $p_{t}=\operatorname*{arg\,max}_{p}P_{t-1}(p)$.
-Regardless of the selected prior, HP-GP-TS observes an observation $y_{t}$ to update the hyperposterior $P_{t}$.
+As baselines, we use PE-GP-UCB and Maximum A Posteriori (MAP) GP-TS where MAP GP-TS is identical to HP-GP-TS except for greedily selecting $p_{t}$ from the posterior: $p_{t}=\operatorname*{arg\,max}_{p}P_{t-1}(p)$ [^1] Note that since the hyperprior is uniform, MAP is equivalent to discrete maximum likelihood estimation..
+Regardless of the selected prior, the weighting of the hyperposterior $P_{t}$ is updated for all priors.
 Hence, greedily selecting the prior could reduce unnecessary exploration.
-In addition, we investigate the oracle variants of PE-GP-TS and PE-GP-UCB with $\delta=0.05$ that are only given the true prior: $P_{0}=\{p^{*}\}$.
-Figure 6: Average entropy in the hyperposterior $P_{t}$ over time for HP- and MAP GP-TS.
-The dashed reference values correspond to entropies of discrete distributions with prob.
-$q$ on one choice and prob.
-$\frac{1-q}{|P|-1}$ on the other $|P|-1$ choices.
-The cumulative regret for the three synthetic experiments is shown in Fig.
-4.
+In addition, we investigate the oracle variants of PE-GP-TS and PE-GP-UCB with $\delta=0.05$ that are only given the true prior: $P_{1}=\{p^{*}\}$.
+The cumulative regret for the three synthetic experiments is shown in Fig.˜3.
 Across all three experiments, we observe that HP-GP-TS has lower regret than the other methods and performs close to the oracle GP-TS.
-For the kernel and hill experiments, PE-GP-TS has lower regret than the oracle GP-UCB.
+For the kernel and subspace experiments, PE-GP-TS has lower regret than the oracle GP-UCB.
 Hence, even if PE-GP-UCB was optimized to perform as well as the oracle, it would still not achieve the regret of our proposed methods.
-MAP GP-TS has slightly higher regret than HP-GP-TS for the lengthscale and hill experiments but has significantly higher regret for the kernel experiment.
-The greedy selection causes under-exploration for MAP GP-TS in certain instances.
-In the three left-most columns of Fig.
-5, the average amount of priors remaining is shown over time for the synthetic experiments (for the PE-methods methods).
-For the hill experiment, no priors are eliminated for either PE-GP-TS nor PE-GP-UCB.
-For the kernel and lengthscale experiment, PE-GP-TS eliminates more priors than its UCB-counterpart.
-However, note that, on average, less than 0.15 priors are eliminated in the lengthscale experiment and slightly more than 1.2 priors in the kernel experiment.
-Similarly, in Fig.
-6, the average entropy in the hyperposterior $P_{t}$ is shown over time.
-Across the three experiments, the hyperposterior of HP-GP-TS concentrates more than that of MAP GP-TS - which could be due to the greedy prior selection.
-As a comparison, HP-GP-TS places more than 95% of the probability mass on one prior (based on the reference values) in the lengthscale experiment where as the PE-methods barely removed priors.
-Figure 7: Confusion matrices for the true prior $p^{*}$ and the selected priors $p_{t}$ across all time steps of the kernel experiment.
-In Fig.
-7, we visualize how often the methods select the true prior $p^{*}$ (or kernel) in the synthetic experiments as confusion matrices.
-PE-GP-UCB almost always selects the prior using the Matérn-3/2 kernel.
-This prior induces a distribution over functions that are less smooth compared to the other kernels.
-Hence, the Matérn-3/2 kernel produces much higher confidence intervals outside the observed data and is thus more optimistic, leading to excessive exploration.
+MAP GP-TS has slightly higher regret than HP-GP-TS for the lengthscale and subspace experiments but has significantly higher regret and variance for the kernel experiment.
+The greedy selection of MAP (MLE) leads to under-exploration for MAP GP-TS in certain instances.
+The number of priors remaining $|P_{t}|$ and the hyperposterior entropy for the kernel experiment is shown in Fig.˜4.
+The PE-methods eliminate at most one prior on average.
+In contrast, the hyperposterior entropy of HP-GP-TS is equivalent to 80-90% of the probability mass being assigned to one prior.
+HP- and MAP-GP-TS thus effectively discards priors at a much faster rate.
+The same pattern holds for the lengthscale and subspace experiments, see Figs.˜9 and 10 in Appendix˜D.
+Figure 5: Confusion matrices for the true prior $p^{*}$ and the selected priors $p_{t}$ for the kernel experiment. Row-wise normalized to 100%.
+In Fig.˜5, we visualize how often the methods select the true prior $p^{*}$ (or kernel) in the kernel experiment as confusion matrices.
+PE-GP-UCB selects the Matérn-3/2 kernel more than 96% of the rounds.
+The Matérn-3/2 kernel induces a distribution over functions that are less smooth compared to the other kernels and produces much higher confidence intervals outside the observed data leading to excessive optimistic exploration.
 PE-GP-TS also shows a bias towards the Matérn-3/2 kernel but does not select it as frequently as PE-GP-UCB - demonstrating that one layer of optimism has been removed.
-The overall “accuracy” of the selected priors, i.e.
-${\sum_{t\in[T]}}\mathds{1}\{p_{t}=p^{*}\}/T$ where $\mathds{1}$ is the indicator function, of the elimination-based methods is around $17\%$ in the kernel experiment compared to $62.5\%$ and $63.2\%$ for MAP and HP-GP-TS respectively.
+The overall “accuracy" of the selected priors, i.e.
+${\sum_{t\in[T]}}\mathds{1}\{p_{t}=p^{*}\}/T$, for the elimination-based methods is around $17\%$ in the kernel experiment compared to $62.5\%$ and $63.2\%$ for MAP and HP-GP-TS respectively.
 For HP-GP-TS, we observe that it can easily identify the periodic and linear kernels.
 However, the RBF, Matérn and RQ kernels are often confused with each other.
-These kernels do not have as easily distinguishable characteristics and are likely to produce similar posteriors given enough data.
-Figure 8: Cumulative regret on Intel temperature data (top) and PeMS data (bottom). Errorbars correspond to ±$\pm 1$ standard error.
+These kernels do not have as easily distinguishable characteristics and are likely to produce similar posteriors even with a small amount of data.
+See Fig.˜11 in Appendix˜D for confusion matrices in the lengthscale and subspace experiments.
+Figure 6: Total regret for the lengthscale and subspace experiments as the number of priors $|P|$ increases.
+Scaling $|P|$
+We perform two experiments to understand how the regret of our algorithms scale with the number of priors.
+In both experiments, the average difficulty of the problem is kept constant such that the regret of the oracle models is constant.
+In the first experiment, we increase the discretization of the lengthscale values.
+The lengthscales are equidistantly spaced in $[0.5,4]$ with $|P|\in\{8,16,32,64,128\}$.
+As $|P|$ increases, the difference between similar priors is reduced.
+In the second experiment, we increase the number of priors in the subspace experiment from 5 up to 16.
+Each prior can share at most 3 out of 4 dimensions with other priors which ensures the priors remain meaningfully different.
+The total regret as the number of priors increases is shown in Fig.˜6.
+For the lengthscale experiment, increasing the number of priors above 8 does not affect the regret for any algorithm.
+This is likely due to the redundancy in the priors.
+However in the subspace experiment, the regret of the prior elimination algorithms scales approximately as $\sqrt{|P|}$ whilst MAP- and HP-GP-TS are consistently close to the constant regret of the oracle.
+Figure 7: Cumulative regret on Intel temperature data (top) and PeMS data (bottom). Errorbars correspond to $\pm 1$ standard error. The average final regret for PE-GP-UCB is 511.9 for PNW.
 Real-world data
-We consider two setups with real-world data.
-Using the Intel Berkeley dataset Madden et al.
-(2004), the first setup considers 46 temperature sensors with measurements over 19 days.
-The goal is to select the sensor with the highest temperature.
-The data is split into 12 training days and 7 test days.
-For each day in the training set, the empirical mean and covariance is computed and set as a prior.
-For each run, the unknown function $f$ is selected as a set of measurements from the test data and the noise variance is set to $\sigma^{2}=0.7^{2}$.
-The second setup uses sensor data from 211 sensors along the I-880 highway in March 2023 from the California Performance Measurement System (PEMS) Chen et al.
-(2001); California Department of Transportation (2024).
-The goal is to select the sensor with the lowest speed in order to identify congestions.
-The sensor data prior to March 21st is used as training data and the rest is used as test data.
-Only sensor readings between 6AM and 13AM are used.
-For each hour in the training set (across all training days), the empirical mean and covariance is again used to construct a prior - yielding 7 priors.
-Again, the test data is used to sample the unknown function $f$ and the noise variance is set to $\sigma^{2}=2.25^{2}$.
-PE-GP-UCB and MAP-GP-TS are used as baselines to compare against PE-GP-TS and HP-GP-TS.
-The elimination methods use $\delta=0.05$, the horizon is set to $T=500$ and all methods are evaluated on $500$ seeds [^1] One outlier was removed from the Intel experiment, see Appendix B for more details..
-The cumulative regret for the real-world data experiments is presented in Fig.
-8.
-For the Intel temperature data, HP-GP-TS and MAP-GP-TS obtain the lowest and second lowest cumulative regret respectively.
-For the PeMS data, PE-GP-TS obtains the lowest cumulative regret of all methods.
-HP-GP-TS and MAP-GP-TS have lower regret than PE-GP-UCB but their slopes are almost linear for the final time steps.
-To understand this better, we visualize quantiles of the total regret in Fig.
-9 and the median cumulative regret in Fig.
-11 in Appendix B.
-The figures show that MAP- and HP-GP-TS have the lowest median regret for both experiments and hence perform best in a majority of instances.
+We perform three experiments with real-world data from the Intel Berkeley dataset (Madden et al., 2004), California Performance Measurement System (PeMS) (Chen et al., 2001; California Department of Transportation, 2024) and Pacific Northwest (PNW) daily precipitation dataset (Widmann & Bretherton, 1999, 2000).
+Each dataset contains measurements from a set of sensors over time.
+We split each dataset into a training and test set where the test set contains the last third of the data.
+Hence, the distribution of the test data may have shifted slightly from the training data.
+The training sets are split further into separate buckets to define our priors.
+For each bucket $p$, we compute the empirical mean $\hat{\mu}_{p}$ and covariance $\hat{\Sigma}_{p}$ which defines the prior ${\mathcal{GP}}(\hat{\mu}_{p},\hat{\Sigma}_{p})$.
+The buckets in the Intel data corresponds to the 12 days in the training dataset.
+For the PeMS data, each hour between 06:00 and 13:00 defines one prior, giving 7 priors.
+For the daily precipication data, each month in the year constitutes a prior, yielding 12 priors.
+When running the experiments, we select a measurement of all sensors from the test data uniformly at random.
+The selected measurements correspond to the unknown function $f(x)$ where $x$ is the sensor index and the goal is then to identify sensors measuring large temperatures, small speeds or high precipication respectively for the three datasets.
+When the algorithms select an arm to evaluate, we add Gaussian noise to $y_{t}$ with variance $\sigma^{2}$ around 5% of the signal variance, similar to Srinivas et al.
+(2012); Bogunovic et al.
+(2016).
+See Appendix˜C for more details about the experimental setup.
+The cumulative regret for the experiments with real-world data is presented in Fig.˜7.
+For the Intel and PNW data, HP- and MAP GP-TS obtain the lowest and second lowest cumulative regret respectively.
+HP- and MAP GP-TS have lower regret than PE-GP-TS initially but PE-GP-TS catches up and has the lowest total regret.
+To understand this better, we visualize quantiles of the total regret in Fig.˜13.
+MAP- and HP-GP-TS have the lowest median regret for all three experiments and hence perform best in a majority of instances.
 However, the 90th and 95th quantiles are considerably larger for the PeMS data which impacts the average regret significantly.
 Hence, for the PeMS data, the prior elimination methods seem to yield more stable results.
-The two right-most columns of Figs.
-5 and 6 shows the average amount of active priors and the average entropy over time in the Intel and PeMS experiment.
-For the PeMS data, barely any priors are eliminated whilst for the Intel data around 0.75 to 0.80 priors are eliminated on average for the PE-methods.
-For the Intel data, the HP- and MAP GP-TS place around than 80% to 90% of the posterior probability mass in one prior based on the reference values.
-On the PeMS data, the posteriors concentrate more heavily - which could be premature given the average regret we observed.
-(a) Intel
-(b) PeMS
-Figure 9: Quantiles of the final cumulative regret on the real-world data experiments.
-The median is highlighted in green.
-The whiskers correspond to the 5th and 95th percentile and the lower and upper edge of the box show the first and third quantile.
-6 Discussion
-Limitations
-The main limitation of all the methods we have considered is that they require the set of priors $P$ to be discrete.
-In addition, their computational cost scales linearly with the number of priors considered.
-In theory, the elimination-based methods could have lower computational cost compared to HP-GP-TS.
-However, in practice, priors are rarely eliminated as shown by Fig.
-5.
-At best, one sixth of the priors is eliminated in the kernel experiment for PE-GP-TS.
-Since only priors that are selected can be eliminated and the confidence parameters $\beta_{t}$, $\xi_{t}$ increasing over time, including more priors could lead to less priors being eliminated overall.
-Future work
-To lower the computational complexity of the elimination methods, one would either need tighter confidence bounds for the elimination criteria or need to allow more than the selected priors to be eliminated.
-The bi-level sampling approach of HP-GP-TS provides efficient exploration but suffers from higher computational cost since it cannot discard priors.
-An interesting line of future work could study how the probabilities in the hyperposterior could be used to discard priors that are unlikely to be selected in the remaining rounds.
-It would also be interesting to study selection criteria for the prior $p_{t}$ that maximize the disagreement among the priors such that evidence for the true prior is generated quicker.
-A key difference between PE-GP-TS and the meta bandits methods of Kveton et al.
-(2021); Basu et al.
-(2021); Hong et al.
-(2022) is that they consider multiple bandit instances sampled from the true prior $p^{*}$.
-Specifically, Hong et al.
-(2022) study a setting with arbitrary interleaving of the instances.
-Generalizing our results for HP-GP-TS to a similar setting would provide an interesting complement to these methods beyond the standard and linear setting.
+In Fig.˜4, the number of priors remaining in $|P_{t}|$ and the hyperposterior entropy is shown for the Intel experiment.
+Similar to the synthetic experiment, on average, the prior elimination methods eliminate less than 1 prior whereas the hyperposteriors of HP- and MAP GP-TS concentrate the equivalent of 80-90% of the probability mass to one prior.
+The results for PeMS and PNW experiment are shown in Figs.˜9 and 10 in Appendix˜D.
+Here, effectively no priors are eliminated.
+For the PNW experiment, the hyperposteriors do not concentrate as much compared to the other experiments.
+This could indicate that knowing the exact prior is not as important for the PNW data.
 6 Conclusion
 In this paper, we have proposed two algorithms for joint prior selection and regret minimization in GP bandits based on GP-TS.
-We have analyzed the algorithms theoretically and provided regret bounds whose order closely match previous work.
-We have experimentally evaluated both algorithms on both synthetic and real-world data.
+We have analyzed the algorithms theoretically and have experimentally evaluated both algorithms on synthetic and real-world data.
 We find that they both select the true prior more often and obtain lower regret than previous work due to lowering the amount of optimistic exploration.
-Acknowledgements
+Acknowledgments
 The work of Jack Sandberg and Morteza Haghir Chehreghani was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
@@
 First, we state and prove concentration inequalities for $f(x)$ and $\tilde{f}_{t,p}(x)$.
-See 4
+See 4.1
 Proof.
 Follows by the same steps as Lemma 5.1 of Srinivas except we condition on the complete history $H_{t}$ instead of only ${\mathbf{y}}_{1:t-1}$.
-Additionally, for Eq.
-5 we must take an additional union bound over $p\in P$.
+Additionally, for Eq.˜5 we must take an additional union bound over $p\in P$.
 Fix $t,x,p\in[T]\times{\mathcal{X}}\times P$.
@@
 $=\frac{3\delta}{|{\mathcal{X}}||P|\pi^{2}t^{2}}$ (13)
-Note that $\sum_{t\geq 1}\frac{1}{t^{2}}=\frac{\pi^{2}}{6}$y taking the union bound over ${\mathcal{X}}$, $P$ and $t\geq 1$, Eq.
-5 holds w.p.
-at least $21-\delta/$ 2.
-By the same reasoning and skipping the union bound over $P$, Eq.
-4 holds w.p.
-at least $21-\delta/$ 2.
+Note that $\sum_{t\geq 1}\frac{1}{t^{2}}=\frac{\pi^{2}}{6}$By taking the union bound over $\mathcal{X}$}, $P$ and $t\geq 1$, Eq.˜5 holds w.p.
+at least $1-\delta/2$.
+By the same reasoning and skipping the union bound over $P$, Eq.˜4 holds w.p.
+at least $1-\delta/2$.
 Thus, both events hold w.p.
@@
 ∎
-Next, we state three lemmas from Ziomek et al. (2024b) that are used in the proof of our regret bound.
+Next, we state three lemmas from Ziomek et al. (2025) that are used in the proof of our regret bound.
 {lemma}
-(Lemma 5.1 of Ziomek et al. (2024b)) If $\xi_{t}=2\sigma^{2}\log\left(\frac{|P|\pi^{2}t^{2}}{6\delta}\right)$, then the following holds with probability at least $1-\delta$:
+(Lemma 5.1 of Ziomek et al. (2025)) If $\xi_{t}=2\sigma^{2}\log\left(\frac{|P|\pi^{2}t^{2}}{6\delta}\right)$, then the following holds with probability at least $1-\delta$:
 $\left|\sum_{i\in S_{t,p}}\epsilon_{i}\right|\leq\sqrt{\xi_{t}|S_{t,p}|}\quad\forall t,p\in[T]\times P.$ (14)
@@
 (Lemma 5.2 of Ziomek et al.
-(2024b)) Let $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$, then if $\mu_{1,p^{*}}$ and $k_{1,p^{*}}$ satisfy $|\mu_{1,p^{*}}(\cdot)|<\infty$ and $k_{1,p^{*}}(\cdot,\cdot)\leq 1$ and Section 4 holds, then
+(2025)) Let $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$, then if $\mu_{1,p^{*}}$ and $k_{1,p^{*}}$ satisfy $|\mu_{1,p^{*}}(\cdot)|<\infty$ and $k_{1,p^{*}}(\cdot,\cdot)\leq 1$ and Section˜4.1 holds, then
 $\sup_{x\in{\mathcal{X}}}|f(x)|\leq B_{p^{*}}.$ (15)
@@
 (Lemma 5.3 of Ziomek et al.
-(2024b)) For $C=2/\log(1+\sigma^{-2})$, $\sum_{t\notin{\mathcal{C}}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\leq\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}$ where $\beta_{T}=\max_{p\in P}\beta_{T}$ and $\hat{\gamma}_{T}=\max_{p\in P}{\gamma}_{T,p}$.
+(2025)) For $C=2/\log(1+\sigma^{-2})$, $\sum_{t\notin\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\leq\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}$ where $\beta_{T}=\max_{p\in P}\beta_{T}$ and $\hat{\gamma}_{T}=\max_{p\in P}{\gamma}_{T,p}$.
 Then, we state and prove the new simple regret bound for PE-GP-TS.
-See 4
+See 4.1
 Proof.
@@
 […]
+$\leq\tilde{f}_{t,p_{t}}(x_{t})+2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*}).$ $(\text{TS selection rule})$ (18)
+Then, we lower bound $f(x_{t})$
+$f(x_{t})$ $=\mu_{t,p_{t}}(x_{t})+\eta_{t}-\epsilon_{t}$ $(\text{Def. of }\eta_{t})$ (19)
 […]
-Then, we lower bound $f(x_{t})$
-[…]
-[…]
-Combining, Eqs. 18 and 20 we obtain
+Combining, Eqs.˜18 and 20 we obtain
 $f(x^{*})-f(x_{t})\leq 2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-\eta_{t}+\epsilon_{t}.$ (21)
@@
 Proof.
-First, we show that the true prior $p^{*}$ is never rejected if Sections 4 and A.1 hold.
+First, we show that the true prior $p^{*}$ is never rejected if Sections˜4.1 and A.1 hold.
 $\left|\sum_{i\in S_{t,p^{*}}}\eta_{i}\right|$ $=\left|\sum_{i\in S_{t,p^{*}}}(y_{i}-f(x_{i})+f(x_{i})-\mu_{i,p^{*}}(x_{i})\right|$ (22)
-[…]
+$\leq\left|\sum_{i\in S_{t,p^{*}}}\epsilon_{i}\right|+\sum_{i\in S_{t,p^{*}}}\left|f(x_{i})-\mu_{i,p^{*}}(x_{i})\right|$ $(\text{Triangle ineq.)}$ (23)
 […]
@@
 $\mathcal{C}=\left\{t\in[T]:\left|\sum_{i\in S_{t,p_{t}}}\eta_{i}\right|>\sqrt{\xi_{t}S_{t,p_{t}}}+\sum_{i\in S_{t,p_{t}}}\sqrt{\beta_{i}}\sigma_{i,p_{t}}(x_{i})\right\}.$ (25)
-Using Eqs. 14 and 21, we can bound the cumulative regret as follows:
+Using Eqs.˜14 and 21, we can bound the cumulative regret as follows:
 ${\text{BR}}(T)$ $=\sum_{t\in\mathcal{C}}{\text{BR}}_{t}+\sum_{t\notin\mathcal{C}}{\text{BR}}_{t}$ (26)
 $\leq 2|P|B_{p^{*}}+\sum_{t\notin\mathcal{C}}2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+\sum_{t\notin\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})+\sum_{p\in P}\sum_{t\in S_{T,p}\setminus\mathcal{C}}(\epsilon_{t}-\eta_{t}).$ (27)
-where $B_{p^{*}}:=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$. If $t\notin{\mathcal{C}}$, line 9 in Algorithm 1 evaluates to false and hence
+where $B_{p^{*}}:=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$. If $t\notin\mathcal{C}$, line˜9 in Algorithm˜1 evaluates to false and hence
 $\sum_{p\in P}\sum_{t\in S_{T,p}\setminus\mathcal{C}}-\eta_{t}\leq\sum_{p\in P}\sqrt{\xi_{T}|S_{T,p}|}+\sum_{p\in P}\sum_{t\in S_{T,p}\setminus\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p}(x_{t}).$ (28)
-Additionally, using Section A.1, we can bound the Gaussian noise:
+Additionally, using Section˜A.1, we can bound the Gaussian noise:
 $\sum_{p\in P}\sum_{t\in S_{T,p}\setminus\mathcal{C}}\epsilon_{t}$ $\leq\sum_{p\in P}\left|\sum_{t\in S_{T,p}\setminus\mathcal{C}}\epsilon_{t}\right|\leq\sum_{p\in P}\left|\sum_{t\in S_{T,p}}\epsilon_{t}\right|$ (29)
 […]
-[…]
+$\leq\sqrt{\xi_{T}|P|T}$ $(\text{Cauchy-Schwarz})$ (31)
 Combining the above, the cumulative regret is bounded by
 ${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sum_{t\notin\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+2\sum_{t\notin\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t}).$ (32)
-Finally, applying Eqs. 15 and 4.1, we obtain the result
-${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}\gamma_{T,p^{*}}}+2\sqrt{CT\beta_{T}\gamma_{T}|P|}.$ (33)
-A.2 HP-GP-TS
-First, we state and prove Theorem 4.1. See 4.1
-Proof.
-First, note that $I_{p}({\mathbf{y}}_{T};f)=H({\mathbf{y}}_{T}|p)-H({\mathbf{y}}_{T}|f,p)$.
-As in Srinivas et al.
-(2012), $H({\mathbf{y}}_{T}|f,p)$ is simply the entropy of $T$ zero-mean Gaussians with variance $\sigma^{2}$: $\frac{T}{2}\log(2\pi e\sigma^{2})$.
-Using the chain rule of entropy, $H({\mathbf{y}}_{T}|p)=H({\mathbf{y}}_{T-1}|p)+H(y_{T}|{\mathbf{y}}_{T-1},p)=H({\mathbf{y}}_{T-1}|p)+\frac{1}{2}\log(2\pi e(\sigma^{2}+\sigma^{2}_{t}(x_{T}))$ where the second equality is due to $y_{T}|{\mathbf{y}}_{T-1},p\sim{\mathcal{N}}(\mu_{T,p}(x_{T}),\sigma^{2}_{T,p}(x_{T})+\sigma^{2})$.
-The rest follows by induction.
-The second part follows from the proof of Lemma 5.4 in Srinivas et al. (2012):
-$\sigma_{t,p}^{2}(x_{t})$ $=\sigma^{2}(\sigma^{-2}\sigma^{2}_{t,p}(x_{t}))$ (34)
-$\leq\sigma^{2}\frac{\sigma^{-2}}{\log(1+\sigma^{-2})}\log(1+\sigma^{-2}\sigma^{2}_{t,p}(x_{t}))$ (35)
-$\leq\frac{1}{\log(1+\sigma^{-2})}\log(1+\sigma^{-2}\sigma^{2}_{t,p}(x_{t})).$ (36)
-Using Eq. 8, we get the desired result. ∎
-Based on Theorem 4.1, Theorem 4.1 follows almost immediately.
-See 4.1
-Proof.
-From Theorem 4.1, we have for any fixed sequence of arms $x_{1},\ldots,x_{T}\in{\mathcal{X}}$ and prior $p^{*}\in P$, that ${\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}\leq CI_{p^{*}}({\mathbf{y}}_{T};f)$.
-The mutual information $I_{p^{*}}({\mathbf{y}}_{T};f)$ given the prior $p^{*}$ is in turn bounded by the maximal information gain $\gamma_{T,p^{*}}$ for the same prior.
-Hence,
-${\mathbb{E}}\left[{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})\right]\leq{\mathbb{E}}\left[C\gamma_{T,p^{*}}\right]={\mathbb{E}}_{p^{*}\sim P_{0}}\left[C\gamma_{T,p^{*}}\right].$ (37)
-Then, we are ready to state and prove our regret bound for HP-GP-TS.
+Finally, applying Eq.˜15, we obtain the result
+${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}.$ (33)
+A.2 UCB-analysis of HP-GP-TS
+Next, we state and prove our UCB-based regret bound for HP-GP-TS.
 See 4.2
 Proof.
-[…]
+To begin, we note that $x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t}$ and $p_{t}|H_{t}\overset{d}{=}p^{*}|H_{t}$ since both $x_{t}$ and $p_{t}$ are sampled from their respective posteriors.
+Let $U_{t,p}(x):=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$.
+Then, we start decomposing the instant regret into two terms:
 ${\text{BR}}(T)$ $={\sum_{t\in[T]}}{\mathbb{E}}\left[f(x^{*})-f(x_{t})\right]$ (34)
-[…]
-$=\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}\left[f(x^{*})-U_{t,p^{*}}(x^{*})\right]}_{(1)}+\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p_{t}}(x_{t})-f(x_{t})\right]}_{(2)}.$ (40)
+$={\sum_{t\in[T]}}{\mathbb{E}}[f(x^{*})-U_{t,p^{*}}(x^{*})+U_{t,p^{*}}(x^{*})$ (35)
+$\qquad\qquad-U_{t,p_{t}}(x^{*})+U_{t,p^{*}}(x_{t})-f(x_{t})]\hskip 60.00009pt(x^{*},p_{t}|H_{t}\overset{d}{=}x_{t},p^{*}|H_{t})$ (36)
+$=\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}\left[f(x^{*})-U_{t,p^{*}}(x^{*})\right]}_{(1)}+\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p_{t}}(x_{t})-U_{t,p^{*}}(x_{t})\right]}_{(2)}+\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p^{*}}(x_{t})-f(x_{t})\right]}_{(3)}$ (37)
 We begin by bounding term $(1)$,
@@
 $\leq{\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}{\mathbb{E}}\left[\left[f(x)-\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)\right]_{+}\right]$ $(x^{*}\in{\mathcal{X}},[\cdot]_{+}\geq 0)$ (40)
-[…]
+$\leq{\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}{\mathbb{E}}_{p^{*},H_{t}}\left[{\mathbb{E}}_{t}\left[\left[f(x)-\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)\right]_{+}\big|\,p^{*},H_{t}\right]\right]$ $(\text{Tower rule})$ (41)
 Recall that for $Z\sim{\mathcal{N}}(\mu,\sigma)$ with $\mu\leq 0$, ${\mathbb{E}}[[Z]_{+}]=\frac{\sigma}{\sqrt{2\pi}}\exp\left(\frac{-\mu^{2}}{2\sigma^{2}}\right)$.
-In our case, note that $f|p^{*},H_{t}\sim{\mathcal{N}}(\mu_{t,p^{*}},k_{t,p^{*}})$ and $\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)$ is deterministic given $p^{*},H_{t}$.
+In our case, note that $f(x)|p^{*},H_{t}\sim{\mathcal{N}}(\mu_{t,p^{*}}(x),\sigma^{2}_{t,p^{*}}(x))$ and $\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)$ is deterministic given $p^{*},H_{t}$.
 Hence,
@@
 Next, we bound $(3)$ as follows:
-$(2)$ $={\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p_{t}}(x_{t})-f(x_{t})\right]$ (49)
-[…]
-$={\sum_{t\in[T]}}{\mathbb{E}}_{H_{t}}\left[{\mathbb{E}}_{t}\left[U_{t,p_{t}}(x_{t})-\mu_{t,p_{t}}(x_{t})|H_{t}\right]\right]$ $({\mathbb{E}}[f(x_{t})|H_{t}]={\mathbb{E}}[\mu_{t,p_{t}}(x_{t})|H_{t}])$ (51)
-$={\sum_{t\in[T]}}{\mathbb{E}}_{H_{t}}[{\mathbb{E}}_{t}[\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})|H_{t}]]$ $(U_{t,p_{t}}(\cdot)=\mu_{t,p_{t}}(\cdot)+\sqrt{\beta_{t}}\sigma_{t,p_{t}}(\cdot))$ (52)
-[…]
+$(3)$ $={\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p^{*}}(x_{t})-f(x_{t})\right]$ (46)
+$={\sum_{t\in[T]}}{\mathbb{E}}_{H_{t}}\left[{\mathbb{E}}_{t}\left[U_{t,p^{*}}(x_{t})-f(x_{t})|H_{t}\right]\right]$ $(\text{Tower rule})$ (47)
+$={\sum_{t\in[T]}}{\mathbb{E}}_{H_{t}}\left[{\mathbb{E}}_{t}\left[U_{t,p^{*}}(x_{t})-\mu_{t,p^{*}}(x_{t})|H_{t}\right]\right]$ $({\mathbb{E}}[f(x_{t})|H_{t}]={\mathbb{E}}[\mu_{t,p^{*}}(x_{t})|H_{t}])$ (48)
+$={\sum_{t\in[T]}}{\mathbb{E}}_{H_{t}}[{\mathbb{E}}_{t}[\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x_{t})|H_{t}]]$ $(U_{t,p^{*}}(\cdot)=\mu_{t,p^{*}}(\cdot)+\sqrt{\beta_{t}}\sigma_{t,p^{*}}(\cdot))$ (49)
 $={\sum_{t\in[T]}}{\mathbb{E}}[\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x_{t})]$ (50)
@@
 $(3)$ $={\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x_{t})\right]$ (51)
+$\leq{\mathbb{E}}\left[\sqrt{{\sum_{t\in[T]}}\beta_{t}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})}\right]$ $(\text{Cauchy-Schwarz})$ (52)
+$=\sqrt{{\sum_{t\in[T]}}\beta_{t}}{\mathbb{E}}\left[\sqrt{{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})}\right]$ $(\beta_{t}\text{ deterministic})$ (53)
+$\leq\sqrt{{\sum_{t\in[T]}}\beta_{t}}\sqrt{{\mathbb{E}}\left[{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})\right]}$ $(\text{Jensen's inequality})$ (54)
+$\leq\sqrt{\beta_{T}T}\sqrt{{\mathbb{E}}_{p^{*}}\left[{\mathbb{E}}\left[{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})\,\bigg|\,p^{*}\right]\right]}$ $(\beta_{t}\text{ increasing})$ (55)
 […]
-$\leq\sqrt{{\sum_{t\in[T]}}\beta_{t}}{\mathbb{E}}\left[\sqrt{{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x_{t})}\right]$ $(\beta_{t}\text{ deterministic})$ (57)
+$\leq\sqrt{\beta_{T}T}\sqrt{C\bar{\gamma}_{T}}$ (57)
+Combining the bounds for (1) and (3), we obtain the desired result
+${\text{BR}}(T)\leq\frac{\pi^{2}}{6}+{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p_{t}}(x_{t})-U_{t,p^{*}}(x_{t})\right]+\sqrt{CT\beta_{T}\bar{\gamma}_{T}}.$ (58)
+A.3 Analysis of expected posterior probabilities
+See 8
+Proof.
+Fix $p^{*}=p$ and let $\tilde{p}$ denote the incorrect prior.
+Without loss of generality, assume $\mu_{1,p}(x)=0$ $\forall x\in{\mathcal{X}}$ such that the two priors are: ${\mathcal{GP}}_{p}(0,k)$ and ${\mathcal{GP}}_{\tilde{p}}(\mu,k)$.
+For the observed rewards ${\mathbf{y}}\in{\mathbb{R}}^{t}$, we use $N_{p}({\mathbf{y}})={\mathcal{N}}({\mathbf{y}};\mathbf{0},{\mathbf{K}}+\sigma^{2}I)$ and $N_{\tilde{p}}({\mathbf{y}})={\mathcal{N}}({\mathbf{y}};{\bm{\mu}}(x_{1:t}),{\mathbf{K}}+\sigma^{2}I)$ to denote the likelihood (or multivariate Gaussian density) under the priors $p$ and $\tilde{p}$ where ${\bm{\mu}}(x_{1:t})\in{\mathbb{R}}^{t}$ and ${\mathbf{K}}\in{\mathbb{R}}^{t\times t}$ s.t.
+$({\bm{\mu}})_{i}=\mu(x_{i})$ and $({\mathbf{K}})_{i,j}=k(x_{i},x_{j})$.
+To simplify the notation, let ${\bm{\mu}}:={\bm{\mu}}(x_{1:t})$ and ${\bm{\Sigma}}=K+\sigma^{2}I$ s.t.
+$N_{p}({\mathbf{y}})={\mathcal{N}}({\mathbf{y}};\bf{0},{\bm{\Sigma}})$ and $N_{\tilde{p}}({\mathbf{y}})={\mathcal{N}}({\mathbf{y}};{\bm{\mu}},{\bm{\Sigma}})$.
+By Bayes’ theorem, the posterior probability of $p$ satisfies:
+$P_{t+1}(p)$ $=\frac{N_{p}({\mathbf{y}})P_{0}(p)}{N_{p}({\mathbf{y}})P_{0}(p)+N_{\tilde{p}}({\mathbf{y}})P_{0}(\tilde{p})}=\frac{\frac{N_{p}({\mathbf{y}})}{N_{\tilde{p}}({\mathbf{y}})}P_{0}(p)}{\frac{N_{p}({\mathbf{y}})}{N_{\tilde{p}}({\mathbf{y}})}P_{0}(p)+(1-P_{0}(p))}.$ (59)
+Note that $N_{p}({\mathbf{y}})/N_{\tilde{p}}({\mathbf{y}})=\exp\left(-\frac{1}{2}\left({\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}}-({\mathbf{y}}^{T}-{\bm{\mu}}){\bm{\Sigma}}^{-1}({\mathbf{y}}-{\bm{\mu}})\right)\right)=:\exp(Q({\mathbf{y}}))$ and that the quotient in the integrand can be written as $q(x)=\frac{e^{x}c}{e^{x}c+(1-c)}$ for $c:=P_{0}(p)\in(0,1)$.
+Since we wish to lower bound the expectation of $P_{t+1}(p)$, we will utilize the lower bounds $q(x)\geq ce^{x}$ $\forall x\leq 0$ and $q(x)\geq 1-\frac{1-c}{c}e^{-x}$ $\forall x\in{\mathbb{R}}$.
+The first lower bound will be applied to the region $H:=\{{\mathbf{y}}\big|Q({\mathbf{y}})\leq 0\}$ whilst the second is applied to the complement $H^{c}$:
+${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|x_{1:t}\right]$ $=\frac{1}{\sqrt{(2\pi)^{t}}}\int_{{\mathbb{R}}^{t}}\frac{1}{\sqrt{\det{\bm{\Sigma}}}}\frac{\exp(Q({\mathbf{y}}))c}{\exp(Q({\mathbf{y}})c+1-c}\exp\left(-\frac{1}{2}{\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}}\right)d{\mathbf{y}}$ (60)
+$\geq\frac{1}{\sqrt{(2\pi)^{t}}}\Bigg(c\underbrace{\int_{H}\frac{1}{\sqrt{\det{\bm{\Sigma}}}}\exp(Q({\mathbf{y}}))\exp\left(-\frac{1}{2}{\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}})\right)d{\mathbf{y}}}_{I_{1}:=}$ (61)
+$+\underbrace{\int_{H^{c}}\frac{1}{\sqrt{\det{\bm{\Sigma}}}}\left(1-\frac{1-c}{c}\exp(-Q({\mathbf{y}}))\right)\exp\left(-\frac{1}{2}{\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}})\right)d{\mathbf{y}}}_{I_{2}:=}\Bigg).$ (62)
+Let $\bm{\Sigma}^{\frac{1}{2}}$} be the positive definite and symmetric square root of $\bm{\Sigma}$} s.t.
+${\bm{\Sigma}}={\bm{\Sigma}^{\frac{1}{2}}}{\bm{\Sigma}^{\frac{1}{2}}}$.
+Next, we apply the change of variables ${\mathbf{y}}={\bm{\Sigma}^{\frac{1}{2}}}(v\bar{{\bm{\mu}}}+{\bm{w}})$ s.t.
+${\bm{w}}\perp\bar{{\bm{\mu}}}$ and $\bar{{\bm{\mu}}}:={\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}/\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$.
+Note that the Jacobian of this variable change is $\sqrt{\det{\bm{\Sigma}}}$ and
+${\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}}=(v\bar{{\bm{\mu}}}+{\bm{w}})^{T}{\bm{\Sigma}^{\frac{1}{2}}}{\bm{\Sigma}}^{-1}{\bm{\Sigma}^{\frac{1}{2}}}(v\bar{{\bm{\mu}}}+{\bm{w}})=\|v\bar{{\bm{\mu}}}+{\bm{w}}\|^{2}=v^{2}+\|{\bm{w}}\|^{2},$ (63)
+$({\mathbf{y}}-{\bm{\mu}})^{T}{\bm{\Sigma}}^{-1}({\mathbf{y}}-{\bm{\mu}})$ $=((v-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert)\bar{{\bm{\mu}}}+{\bm{w}})^{T}((v-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert)\bar{{\bm{\mu}}}+{\bm{w}})$ (64)
+$=(v-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert)^{2}+\|{\bm{w}}\|^{2},$ (65)
+$Q({\mathbf{y}})=-\frac{1}{2}\left(v^{2}+\lVert{\bm{w}}\rVert^{2}-(v-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert)^{2}-\lVert{\bm{w}}\rVert^{2}\right)=-\frac{1}{2}\left(2v\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}\right).$ (66)
+The region $H=\{Q({\mathbf{y}})\leq 0\}$ can then be expressed as $H=\left\{(v,{\bm{w}})\big|v\geq\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert/2,{\bm{w}}\perp{\bm{\mu}}\right\}$ and $H^{c}=\{(v,{\bm{w}})\big|v<\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert/2,{\bm{w}}\perp{\bm{\mu}}\}$.
+Next, we apply the change of variables to $I_{1}$ and decompose the integral into two parts:
+$I_{1}$ $=\int_{H}\exp\left(-\frac{1}{2}\left(-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}+2v\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert+v^{2}+\lVert{\bm{w}}\rVert^{2}\right)\right)dvd{\bm{w}}$ (67)
+$=\exp(\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2})\int_{v\geq\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert/2}\exp\left(-\frac{1}{2}(v+\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert)^{2}\right)dv\int_{{\mathbb{R}}^{t-1}}\exp\left(-\frac{\lVert{\bm{w}}\rVert^{2}}{2}\right)d{\bm{w}}$ (68)
+$=\exp(\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2})\cdot\sqrt{2\pi}\left(1-\Phi\left(\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\right)\right)\cdot\sqrt{(2\pi)^{t-1}}$ (69)
+$=\sqrt{(2\pi)^{t}}\exp(\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2})\Phi\left(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\right)\qquad\qquad\qquad\qquad\left(\Phi(-x)=1-\Phi(x)\right)$ (70)
+where $\Phi(\cdot)$ is the CDF of the unit-Gaussian. Similarly, we apply the change of variables to $I_{2}$ and decompose the integral into four parts:
+$I_{2}$ $=\int_{H^{c}}\Bigg(\exp\left(-\frac{1}{2}\left(v^{2}+\lVert{\bm{w}}\rVert^{2}\right)\right)$ (71)
+$\qquad\qquad-\frac{1-c}{c}\exp\left(-\frac{1}{2}(-2v\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert+\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}+v^{2}+\lVert{\bm{w}}\rVert^{2})\right)\Bigg)dvd{\bm{w}}$ (72)
+$=\int_{v<\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert/2}\exp\left(-\frac{v^{2}}{2}\right)dv\int_{{\mathbb{R}}^{t-1}}\exp\left(-\frac{w^{2}}{2}\right)d{\bm{w}}$ (73)
+$\quad-\frac{1-c}{c}\int_{v<\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert/2}\exp\left(-\frac{1}{2}\left(v-\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\right)^{2}\right)dv\int_{{\mathbb{R}}^{t-1}}\exp\left(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}{2}\right)d{\bm{w}}$ (74)
+$=\sqrt{2\pi}\Phi\left(\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\right)\cdot\sqrt{(2\pi)^{t-1}}-\frac{1-c}{c}\sqrt{2\pi}\Phi\left(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\right)\cdot\sqrt{(2\pi)^{t-1}}$ (75)
+$=\sqrt{(2\pi)^{t}}\left(1-\frac{1}{c}\Phi\left(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\right)\right).\qquad\qquad\qquad\qquad\left(\Phi(-x)=1-\Phi(x)\right)$ (76)
+Combining the results for $I_{1}$ and $I_{2}$, we get that
+${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|x_{1:t}\right]$ $\geq 1+P_{0}(p)e^{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}\Phi\left(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\right)-\frac{1}{P_{0}(p)}\Phi\left(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\right)$ (77)
+Note that a sharp bound of Eq.˜77 in terms of elementary functions can be obtained using bounds on the error function (Cook, 2018; Abramowitz & Stegun, 1972):
+${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|x_{1:t}\right]$ $\geq 1+P_{0}(p)\sqrt{\frac{2}{\pi}}e^{-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}{8}}\frac{2}{3\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert+\sqrt{9\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}+16}}$ (78)
+$\quad-\frac{1}{P_{0}(p)}\sqrt{\frac{2}{\pi}}e^{-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}{8}}\frac{2}{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert+\sqrt{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}+32\pi^{-1}}}.$ (79)
+A.4 Information-theoretic regret bound for GP-TS
+We begin by showing that the rewards are subgaussian.
+{lemma}[] Fix $x\in{\mathcal{X}}$ and let $Z=f(x)+\epsilon_{t}-{\mathbb{E}}_{t}[f(x)]$.
+If $p^{*}\sim P_{1}$, $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ with $k_{1,p^{*}}(\cdot,\cdot)\leq\sigma_{0}^{2}$, then $Z|H_{t}$ is $\sqrt{\sigma_{0}^{2}+\sigma^{2}}$-subgaussian.
+Proof.
+Start by considering $Z=\sum_{p\in P}\mathds{1}\{p^{*}=p\}(f_{p}(x)-\mu_{t,p}(x))$ where $f_{p}\sim{\mathcal{GP}}_{p}$ such that $f_{p}(x)-\mu_{t,p}(x)$ is $\sigma_{0}$-subgaussian. Then,
+${\mathbb{E}}[\lambda Z|H_{t}]$ $={\mathbb{E}}[{\mathbb{E}}[\exp(\lambda Z)\,|\,p^{*}=p,H_{t}]$ (80)
+$=\sum_{p\in P}{\mathbb{P}}_{t}(p^{*}=p){\mathbb{E}}[\exp(\lambda(f_{p}(x)-\mu_{t,p}(x)))\,|\,p^{*}=p,H_{t}]$ (81)
+$\leq\sum_{p\in P}{\mathbb{P}}_{t}(p^{*}=p)\exp(\sigma_{t,p}^{2}(x)\lambda^{2}/2)$ (82)
+$\leq\sum_{p\in P}{\mathbb{P}}_{t}(p^{*}=p)\exp(\sigma_{0}^{2}\lambda^{2}/2)$ (83)
+$=\exp(\sigma^{2}_{0}\lambda/2)$ (84)
+Hence, $Z$ is $\sigma$-subgaussian. Then, $Z+\epsilon_{t}|H_{t}$ is $\sqrt{\sigma^{2}_{0}+\sigma^{2}}$-subgaussian since $Z|H_{t}$ and $\epsilon_{t}|H_{t}$ are independent. ∎
+Then, we state and prove the information-theoretic regret bound for HP-GP-TS. See 4.3
+Proof.
+Here, we analyze the regret of HP-GP-TS using the information-theoretic framework of Russo & Van Roy (2016).
+The proof presented here follows the proof in section D.2 of Russo & Van Roy (2016) with subgaussian noise.
+The idea of the information-theoretic framework is to express the regret as follows
+${\mathbb{E}}\left[{\sum_{t\in[T]}}f(x^{*})-f(x_{t})\right]$ $={\mathbb{E}}\left[{\sum_{t\in[T]}}{\mathbb{E}}\big[\underbrace{f(x^{*})-f(x_{t})}_{\geq 0}|H_{t}\big]\right]$ $(\text{Tower rule})$ (85)
+$={\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{{\mathbb{E}}\left[f(x^{*})-f(x_{t})|H_{t}\right]^{2}\cdot\frac{I(\cdot\,;\cdot\,|H_{t})}{I(\cdot\,;\cdot\,|H_{t})}}\right]$ (86)
+where $I(\cdot\,;\cdot\,|H_{t})$ is the mutual information between two carefully chosen variables such that ${\mathbb{E}}\left[f(x^{*})-f(x_{t})|H_{t}\right]^{2}\leq CI(\cdot\,;\cdot\,|H_{t})$ for some $C$.
+Let ${\mathbb{E}}_{t}[\cdot]={\mathbb{E}}[\cdot|H_{t}],{\mathbb{P}}_{t}[\cdot]={\mathbb{P}}[\cdot|H_{t}]$. Then,
+${\mathbb{E}}_{t}[f(x^{*})-f(x_{t})]=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x){\mathbb{E}}_{t}[f(x)|x^{*}=x]-\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x){\mathbb{E}}_{t}[f(x)|x_{t}=x]$ (87)
+$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x)\left({\mathbb{E}}_{t}[f(x)|x^{*}=x]-{\mathbb{E}}_{t}[f(x)|x_{t}=x]\right)\quad$ $\left({\mathbb{P}}_{t}(x_{t}=x)={\mathbb{P}}_{t}(x^{*}=x)\right)$ (88)
 […]
+Let $Z=f(x)+\epsilon_{t}-{\mathbb{E}}_{t}[f(x)]$.
+Note that $Z|H_{t}$ is $\sqrt{\sigma_{0}^{2}+\sigma^{2}}$-subgaussian.
+Consequently, $\log{\mathbb{E}}_{t}[\exp(\lambda Z)]\leq\frac{\lambda^{2}(\sigma_{0}^{2}+\sigma^{2})}{2}$ $\forall\lambda\in{\mathbb{R}}$.
+Fix $x^{*},x\in{\mathcal{X}}$. Then,
+$\lambda\Big({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\Big)-\frac{\lambda^{2}(\sigma_{0}^{2}+\sigma^{2})}{2}$ (90)
+$\leq\lambda\Big({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\Big)$ (91)
+$\quad-\log{\mathbb{E}}_{t}[\exp(\lambda(f(x)+\epsilon_{t}-{\mathbb{E}}_{t}[f(x)])]$ $(Z|H_{t}\text{ is }\sqrt{\sigma_{0}^{2}+\sigma^{2}}\text{-subgaussian})$ (92)
+$=\lambda{\mathbb{E}}_{t}[Z|x^{*}=x^{*}]-\log{\mathbb{E}}_{t}[\exp(\lambda Z)]$ (93)
 […]
-Combining the bounds for (1) and (2), we obtain
-${\text{BR}}(T)\leq\frac{\pi^{2}}{6}+\sqrt{CT\beta_{t}{\mathbb{E}}_{p^{*}}[\gamma_{T,p^{*}}]}.$ (60)
-Appendix B Additional experimental details
-In this section, we provide some additional experimental details and results.
-First, we include the full set of confusion matrices for the synthetic lengthscale and hill experiments in Fig.
-10.
+Now, let $\lambda=\frac{1}{\sigma_{0}^{2}+\sigma^{2}}\left({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\right)$, then the following holds for all $x^{*},x\in{\mathcal{X}}$
+${\mathbb{E}}_{t}\left[f(x)|x^{*}=x^{*}\right]-{\mathbb{E}}_{t}[f(x)]\leq\sqrt{2(\sigma_{0}^{2}+\sigma^{2})D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})||{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)}.$ (95)
+Let $I_{t}(\cdot\,;\cdot\,)=I(\cdot\,;\cdot\,|H_{t}=h_{t})$. Then,
+$I_{t}(x^{*};(x_{t},y_{t}))=I_{t}(x^{*};x_{t})+I_{t}(x^{*};y_{t}|x_{t})$ $(\text{Chain rule})$ (96)
+$=I_{t}(x^{*};y_{t}|x_{t})$ $(x^{*}|H_{t}{\perp\!\!\!\!\perp}x_{t}|H_{t})$ (97)
+$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};y_{t}|x_{t}=x)$ (98)
+$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};f(x_{t})+\epsilon_{t}|x_{t}=x)$ (99)
+$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};f(x)+\epsilon_{t})$ $(f,x^{*}|H_{t}{\perp\!\!\!\!\perp}x_{t}|H_{t})$ (100)
+$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)\Bigg(\sum_{x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ (101)
+$\quad\quad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})\,\Big|\Big|\,{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)\Bigg)$ (102)
+$=\sum_{x,x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ $(x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t})$ (103)
+$\quad\quad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})\,\Big|\Big|\,{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)$ (104)
+Putting the above together, we now bound ${\mathbb{E}}_{t}[f(x^{*})-f(x_{t})]^{2}$:
+${\mathbb{E}}_{t}[f(x^{*})-f(x_{t})]^{2}=\left(\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x)\left({\mathbb{E}}_{t}[f(x)|x^{*}=x]-{\mathbb{E}}_{t}[f(x)]\right)\right)^{2}$ (105)
+$\leq|{\mathcal{X}}|\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x)^{2}\left({\mathbb{E}}_{t}[f(x)|x^{*}=x]-{\mathbb{E}}_{t}[f(x)]\right)^{2}$ $(\text{Cauchy-Schwarz})$ (106)
+$\leq|{\mathcal{X}}|\sum_{x,x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\left({\mathbb{E}}_{t}[f(x)|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\right)^{2}$ (107)
+$\leq 2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})\sum_{x,x^{*}\in{\mathcal{X}}}\Big({\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ (108)
+$\qquad\qquad\qquad\qquad\qquad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})||{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)\Big)$ (109)
+$\leq 2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})I_{t}(x^{*};(x_{t},y_{t})).$ (110)
+Returning to the full regret,
+${\mathbb{E}}\bigg[{\sum_{t\in[T]}}f(x^{*})-f(x_{t})\bigg]$ $={\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{{\mathbb{E}}_{t}\left[f(x^{*})-f(x_{t})\right]^{2}\cdot\frac{I_{t}(x^{*}\,;(x_{t},y_{t}))}{I_{t}(x^{*}\,;(x_{t},y_{t}))}}\right]$ (111)
+$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}{\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\right]$ (112)
+Then, the expectation can then be bounded as
+${\mathbb{E}}\bigg[{\sum_{t\in[T]}}\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\bigg]=$ $\leq{\sum_{t\in[T]}}{\mathbb{E}}\left[\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\right]$ (114)
+$\leq{\sum_{t\in[T]}}1\cdot\sqrt{{\mathbb{E}}\left[I_{t}(x^{*}\,;(x_{t},y_{t}))\right]}$ $\left(\text{Jensen's inequality}\right)$ (115)
+$={\sum_{t\in[T]}}\sqrt{{\mathbb{E}}_{H_{t}}\left[I(x^{*}\,;(x_{t},y_{t})|H_{t}=h_{t})\right]}$ (116)
+$={\sum_{t\in[T]}}\sqrt{I(x^{*}\,;(x_{t},y_{t})|H_{t})}$ (117)
+$\leq\sqrt{{\sum_{t\in[T]}}1^{2}{\sum_{t\in[T]}}I(x^{*}\,;(x_{t},y_{t})|H_{t})}$ $(\text{Cauchy-Schwarz})$ (118)
+Finally, the regret can be bounded as
+${\mathbb{E}}\bigg[{\sum_{t\in[T]}}f(x^{*})-f(x_{t})\bigg]$ $=\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{TI(x^{*};H_{T})}$ (119)
+$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{TH(x^{*})}$ (120)
+$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{T\log|{\mathcal{X}}|}$ (121)
+Appendix B Description of kernels
+[…]
+Appendix C Additional experimental details
+In this section, we provide some additional details about the experiments.
+For all the real-world datasets, sensors containing any null measurements have been filtered out.
+The Intel Berkeley dataset consists of measurements from 46 temperature sensors across 19 days.
+The training set consists of the first 12 days of measurements and the remaining 7 days constitute the test set.
+The noise variance is set to $\sigma^{2}=0.7^{2}$.
+The PeMS data consists of measurements from 211 sensors along the I-880 highway from all of 2023.
+The goal is to find the sensors with low speeds to identify congestions.
+We use the 5-min averages provided by PeMS.
+Data between 2023-01-01 and 2023-09-01 is put into the training set whilst the data until 2023-12-31 is put into the test set.
+The noise variance is set to $\sigma^{2}=2.25^{2}$.
+The PNW precipication data consists of daily precipication data from 1949 to 1950 across 167 $50\times 50$ km regions in the Pacific Northwest.
+The goal is to find the region with the highest precipitation for any given day.
+The training data consists of the measurements made prior to 1980 and the test data consists of the measurements between 1980 and 1994.
+The original data is stated to be given in mm/day however the data seems to be off by a factor of 10.
+We rescale the data to a log-scale using $\log(\cdot/10+0.1)$, similar to Krause et al.
+(2008).
+The noise variance is set to $\sigma^{2}=0.41^{2}$.
+In the Intel experiment, we removed one outlier seed.
+All methods had a final cumulative regret around 6000°C, note that the average for the worst performing model across the other seeds was $\approx 150$°C.
+The outlier is shown Fig.˜8.
+We can see that one of the sensors display very high temperatures compared to all other sensors, which is why all methods performed poorly on this seed.
+It should be noted that many of the sensors in the Intel data logged degrees above 100°C after a certain time - likely due to sensor failure rather boiling temperatures in an office environment.
+Also note that these days were excluded from both our training and test data.
+The outlier could be an indication that this particular sensor was starting to fail earlier than others.
+Figure 8: Removed sample from the test data in the Intel experiment. One of the sensors display very high temperatures.
+Appendix D Additional experimental results
+In this section, we provide some additional experimental results.
+Figure 9: Mean number of priors remaining in $P_{t}$ over time for PE-GP-UCB and -TS.
+Figure 10: Average entropy in the hyperposterior $P_{t}$ over time for HP- and MAP GP-TS.
+The dashed reference values correspond to entropies of discrete distributions with prob.
+$q$ on one choice and prob.
+$\frac{1-q}{|P|-1}$ on the other |P|1|P|-1 choices.
+First, we include the mean number of priors in $P_{t}$ for all experiments in Fig.˜9.
+Similarly, we include the average entropy of the hyperposterior for all experiments in Fig.˜10.
+For the lengthscale, subspace, PeMS and PNW precipication experiments, hardly any priors are eliminated.
+In contrast, the hyperposterior entropy concentrates rapidly across all experiments with the subspace and PNW precipication having the most and least concentrated hyperposterior.
+We include the full set of confusion matrices for the lengthscale and subspace experiments in Fig.˜11.
 In the lengthscale experiments, we observe that PE-GP-UCB and -TS oversample the shortest lengthscale.
@@
 However, we see that HP-GP-TS and MAP GP-TS do not suffer from this optimistic bias.
-In the hill experiment, we see that PE-GP-UCB almost never selects Hill 0, i.e.
-the mean function with only smaller hills.
-PE-GP-TS selects the prior almost uniformly whilst HP-GP-TS and MAP-GP-TS select the correct mean function in a majority of the time steps.
-In Fig. 11, we visualize the median cumulative regret on the real-world data experiments.
-In the Intel experiment, we removed one outlier seed.
-All methods had a final cumulative regret around 6000°C, note that the average for the worst performing model across the other seeds was $\approx 150$°C.
-The outlier is shown Fig.
-12.
-We can see that one of the sensors display very high temperatures compared to all other sensors, which is why all methods performed poorly on this seed.
-It should be noted that many of the sensors in the Intel data logged degrees above 100°C after a certain time - likely due to sensor failure rather boiling temperatures in an office environment.
-Also note that these days were excluded from both our training and test data.
-The outlier could be indication that this particular sensor was starting to fail earlier than others.
-(a) Lengthscale experiment
-(b) Hill experiment
-Figure 10: Confusion matrices for the true prior $p^{*}$ and $p_{t}$ across all time steps of the synthetic lengthscale and hill experiments.
-Figure 11: Median cumulative regret on Intel temperature data (top) and PeMS data (bottom). Errorbars correspond to first and last decile.
-Figure 12: Removed sample from the test data in the Intel experiment. One of the sensors display very high temperatures but
+In the subspace experiment, HP- and MAP GP-TS have an accuracy of around 96% where as PE-GP-TS and -UCB have accuracies 30% and 36% respectively.
+The priors are equivalent up to coordinate permutations and are therefore difficult.
+The PE-methods do not oversample any specific prior but commit to much time to exploring along the irrelevant dimensions.
+In Tables˜1 and 2, the total regret for the lengthscale and subspace scaling experiments are shown.
+In Fig.˜12, we visualize the median cumulative regret on the real-world data experiments.
+Similarly, in Fig.˜13, we show further quantiles of the final cumulative regret.
+MAP and HP-GP-TS exhibit similar median regret across the experiments but MAP GP-TS has a larger variance and tail values.
+(a) Kernel experiment
+(b) Lengthscale experiment
+(c) Subspace experiment
+Figure 11: Confusion matrices for the true prior $p^{*}$ and $p_{t}$ across all time steps of the synthetic experiments.
+Table 1: Average total regret and $\pm 1$ standard error for the lengthscale experiment as $|P|$ increases.
+Algorithm Lengthscales, $|P|$
+8 16 32 64 128
+MAP GP-TS 30.2 $\pm$ 1.2 32.4 $\pm$ 2.5 32.5 $\pm$ 2.1 28.7 $\pm$ 1.1 30.8 $\pm$ 1.9
+HP-GP-TS 31.4 $\pm$ 1.0 31.7 $\pm$ 0.9 30.8 $\pm$ 0.8 30.7 $\pm$ 1.0 31.0 $\pm$ 1.4
+PE-GP-TS 61.8 $\pm$ 0.5 61.3 $\pm$ 0.5 62.2 $\pm$ 0.5 62.4 $\pm$ 0.4 64.3 $\pm$ 0.4
+PE-GP-UCB 114.2 $\pm$ 0.6 114.8 $\pm$ 0.6 115.5 $\pm$ 0.6 114.5 $\pm$ 0.6 114.8 $\pm$ 0.6
+Oracle GP-TS 28.1 $\pm$ 0.8 26.4 $\pm$ 0.8 27.3 $\pm$ 0.8 26.5 $\pm$ 0.7 25.7 $\pm$ 0.7
+Oracle GP-UCB 48.3 $\pm$ 1.2 46.9 $\pm$ 1.1 48.4 $\pm$ 1.1 46.5 $\pm$ 1.0 45.6 $\pm$ 1.0
+Table 2: Average total regret and $\pm 1$ standard error for the subspace experiment as $|P|$ increases.
+Algorithm Subspaces, $|P|$
+5 8 12 16
+MAP GP-TS 87.2 $\pm$ 1.0 89.9 $\pm$ 1.1 89.1 $\pm$ 0.9 90.9 $\pm$ 1.2
+HP-GP-TS 88.3 $\pm$ 0.9 88.8 $\pm$ 0.9 89.5 $\pm$ 0.9 90.8 $\pm$ 0.9
+PE-GP-TS 177.1 $\pm$ 1.4 269.5 $\pm$ 1.9 344.7 $\pm$ 2.3 396.9 $\pm$ 2.5
+PE-GP-UCB 389.0 $\pm$ 1.5 526.0 $\pm$ 1.8 622.4 $\pm$ 2.3 688.0 $\pm$ 2.7
+Oracle GP-TS 86.0 $\pm$ 1.0 84.1 $\pm$ 0.9 84.6 $\pm$ 1.0 84.8 $\pm$ 1.0
+Oracle GP-UCB 217.3 $\pm$ 1.0 218.2 $\pm$ 1.0 218.6 $\pm$ 1.0 218.9 $\pm$ 0.9
+Figure 12: Median cumulative regret on Intel temperature data (top), PeMS speed data (middle) and PNW precipication data (bottom). Errorbars correspond to first and last decile.
+(a) Intel
+(b) PeMS
+(c) PNW Precipication
+Figure 13: Quantiles of the final cumulative regret on the real-world data experiments.
+The median is shown with a black line.
+The whiskers correspond to the 5th and 95th percentile and the lower and upper edges of the boxes show the first and third quantile.

### v3

@@
-Efficient Prior Selection in Gaussian Process Bandits with Thompson Sampling
+Adaptive Prior Selection in Gaussian Process Bandits with Thompson Sampling
 Abstract
@@
 Instead, practitioners often rely on maximum likelihood estimation to select the hyperparameters of the prior - which lacks theoretical guarantees.
-In this work, we propose two algorithms for joint prior selection and regret minimization in GP bandits based on GP Thompson sampling (GP-TS): Prior-Elimination GP-TS (PE-GP-TS) and HyperPrior GP-TS (HP-GP-TS).
-We theoretically analyze the algorithms and establish upper bounds for the regret of HP-GP-TS.
-In addition, we demonstrate the effectiveness of our algorithms compared to the alternatives through experiments with synthetic and real-world data.
+In this work, we propose two algorithms for joint prior selection and regret minimization in GP bandits based on GP Thompson sampling (GP-TS): Prior-Elimination GP-TS (PE-GP-TS) that disqualifies priors with poor predictive performance, and HyperPrior GP-TS (HP-GP-TS) that utilizes a bi-level Thompson sampling scheme.
+We theoretically analyze the algorithms and establish upper bounds for their respective regret.
+In addition, we demonstrate the effectiveness of our algorithms compared to the alternatives through extensive experiments with synthetic and real-world data.
 1 Introduction
@@
 However, in a sequential decision making problem MLE is not guaranteed to recover the correct parameters.
-In the literature, Wang & de Freitas (2014); Berkenkamp et al.
+In the literature, Wang and de Freitas (2014); Berkenkamp et al.
 (2019); Ziomek et al.
 (2024) provided algorithms with theoretical guarantees when the kernel lengthscale is unknown.
+Hong et al.
+(2022a) introduced Thompson sampling with mixture priors (MixTS) and provided theoretical guarantees for standard and linear bandits.
+However, we identify apparent technical issues with their proof.
 More recently, Ziomek et al.
@@
 We analyze the regret of PE-GP-TS.
-For the terms we can bound, we obtain a regret bound for PE-GP-TS of order ${\mathcal{O}}(\sqrt{T\log T|P|\hat{\gamma}_{T}})$ where $T$ is the horizon, $|P|$ is the number of priors and $\hat{\gamma}_{T}$ is the worst-case maximum information gain, which matches that of PE-GP-UCB.
+We obtain a regret bound for PE-GP-TS of order ${\mathcal{O}}(\sqrt{T\log T|P|\hat{\gamma}_{T}})$ (which matches that of PE-GP-UCB) plus a term (left unbounded) depending on the uncertainty of the optimal arm under the correct prior.
+Here, $T$ is the horizon, $|P|$ is the number of priors and $\hat{\gamma}_{T}$ is the worst-case maximum information gain.
 The second algorithm, HyperPrior GP-TS (HP-GP-TS), uses bi-level posterior sampling to efficiently explore the priors and arms.
-An UCB-based analysis of HP-GP-TS yields a regret bound of order ${\mathcal{O}}(\sqrt{T\log T\bar{\gamma}_{T}})$ (where $\bar{\gamma}_{T}$ is the average maximum information gain) plus a term that corresponds to the cost of learning the optimal prior.
-An information-theoretic analysis yields a regret bound of order ${\mathcal{O}}(\sqrt{T|{\mathcal{X}}|\log|{\mathcal{X}}|})$ where $|{\mathcal{X}}|$ is the number of arms.
+Our analysis of HP-GP-TS yields a regret bound of order ${\mathcal{O}}(\sqrt{T\log T\bar{\gamma}_{T}})$ (where $\bar{\gamma}_{T}$ is the average maximum information gain) plus a term (left unbounded) that corresponds to the cost of learning the optimal prior.
 We evaluate our methods on three sets of synthetic experiments and three experiments with real-world data.
@@
   • We propose two novel algorithms for GP-bandits with unknown prior: PE- and HP-GP-TS.
-  • We theoretically analyze the regret of HP-GP-TS using a UCB framework and an information-theoretic framework which provides a regret bound of order ${\mathcal{O}}(\sqrt{T|{\mathcal{X}}|\log|{\mathcal{X}}|})$.
-Additionally, we analyze the regret of PE-GP-TS.
-  • We experimentally evaluate our algorithms on both synthetic and real-world data, demonstrating that they achieve superior performance and that the regret of HP-GP-TS does not increase with $|P|$.
+  • We theoretically analyze the regret of PE-GP-TS and HP-GP-TS.
+  • We analyze the proof of the regret bound for MixTS (Hong et al., 2022a) in the linear setting, and identify technical issues.
+  • We experimentally evaluate our algorithms on both synthetic and real-world data, demonstrating that they achieve competitive performance and that the regret of HP-GP-TS does not increase with $|P|$.
 2 Background and problem statement
-Figure 1: Agent deciding between priors based on observed data.
-The unknown function $f$ is sampled from a periodic prior and the observed data is in the top left.
-The agent discards the priors on the right since they do not match the new observation (star).
 Problem statement
@@
 An effective agent must distinguish which prior the means are sampled from to ensure it explores efficiently.
-We illustrate this in Fig.˜1.
 Now, let us formally state the problem.
-Let ${\mathcal{X}}\subseteq[0,r]^{d}\subset{\mathbb{R}}^{d}$ denote the finite set of arms and $P$ a finite set of priors with associated prior mean and kernel functions $\mu_{1,p}:{\mathcal{X}}\mapsto{\mathbb{R}}$ and $k_{1,p}:{\mathcal{X}}\times{\mathcal{X}}\mapsto{\mathbb{R}}$, $\forall p\in P$.
+Let ${\mathcal{X}}\subseteq[0,r]^{d}\subset{\mathbb{R}}^{d}$ denote the finite set of arms and $P$ a finite set of priors with associated prior mean and kernel functions $\mu_{1,p}:{\mathcal{X}}\mapsto{\mathbb{R}}$ and $k_{1,p}:{\mathcal{X}}\times{\mathcal{X}}\mapsto[-1,1]$, $\forall p\in P$.
 Let $p^{*}\in P$ denote the true prior and assume the expected reward function $f:{\mathcal{X}}\mapsto{\mathbb{R}}\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ is a sample from a Gaussian process with prior $p^{*}$.
@@
 Previous work
-Plenty of previous work have proposed fully Bayesian approaches that integrate the acquisition function over the hyperposterior (Osborne et al., 2009; Benassi et al., 2011; Snoek et al., 2012; Hernández-Lobato et al., 2014; Wang & Jegelka, 2017; De Ath et al., 2021; Hvarfner et al., 2023).
+Plenty of previous work have proposed fully Bayesian approaches that integrate the acquisition function over the hyperposterior (Osborne et al., 2009; Benassi et al., 2011; Snoek et al., 2012; Hernández-Lobato et al., 2014; Wang and Jegelka, 2017; De Ath et al., 2021).
 In contrast, HP-GP-TS optimizes a single hyperposterior sample instead of expected values over the hyperposterior.
-Wang & de Freitas (2014) first derived regret bounds for GP bandits with unknown lengthscale for the Expected Improvement algorithm (Močkus, 1975).
+Hvarfner et al.
+(2023) proposed Self-Correcting Bayesian Optimization (SCoreBO) whose objective function balances reducing the uncertainty of $(x^{*},f^{*})$ and reducing the uncertainty of the true prior $p^{*}$.
+Notably, SCoreBO explicitly tries to identify the prior rather than integrating out the uncertainty of the prior.
+Wang and de Freitas (2014) first derived regret bounds for GP bandits with unknown lengthscale for the Expected Improvement algorithm (Močkus, 1975).
 However, the proposed algorithm requires a lower bound on the lengthscale and the regret bound depends on the worst-case MIG.
@@
 (2024) built on this idea and introduced length-scale balancing GP-UCB which can adaptively explore smaller lengthscales but can return to longer ones, unlike A-GP-UCB.
-The aforementioned works are based on UCB, EI, PI or regret balancing.
-However, another line of work has studied Thompson sampling in standard and linear bandits with unknown prior distribution (Kveton et al., 2021; Basu et al., 2021; Hong et al., 2022; Li et al., 2024).
-In their setting (meta or hierarchical bandits), the agent plays multiple bandit instances, either simultaneously or sequentially.
-The unknown means are sampled from the same (unknown) prior and by gathering knowledge across instances, the agent can solve later instances more efficiently once it has identified the prior.
-We emphasize that these methods have been studied only for standard stochastic and linear bandits, not for GP bandits.
-Figure 2: a) Elimination procedure of PE-GP-TS.
+The closest work to ours is Hong et al.
+(2022a).
+Hong et al.
+(2022a) studied Thompson sampling with mixture priors (MixTS) for standard bandits (with bounded rewards) and linear bandits.
+HP-GP-TS can be seen as an instantiation of MixTS for Gaussian process bandits.
+Hong et al.
+(2022a) provide the outline of a regret bound for a standard bandit setting and a detailed proof for the linear setting.
+However, we highlight some issues that we believe invalidate their analysis for the linear case in Appendix˜B.
+Kveton et al.
+(2021); Basu et al.
+(2021); Hong et al.
+(2022b); Li et al.
+(2024) have studied Thompson sampling based algorithms with unknown prior distributions where the agent plays multiple bandit instances.
+We emphasize that these methods have not been studied for GP bandits.
+Figure 1: Elimination procedure of PE-GP-TS.
 The solid lines correspond to posterior means and the shaded regions are confidence intervals.
@@
 The dashed lines are samples from the posteriors.
-b) Overview of HP-GP-TS.
-The orange star corresponds to $y_{t}$.
 3 Algorithms
-As discussed by Russo & Van Roy (2014), TS can offer advantages over UCB algorithms for problems where constructing tight confidence bounds is difficult.
-In addition, Thompson sampling is often observed to perform better than UCB in practice (Chapelle & Li, 2011; Wen et al., 2015; Kandasamy et al., 2018; Åkerblom et al., 2023b, a).
+As discussed by Russo and Van Roy (2014), TS can offer advantages over UCB algorithms for problems where constructing tight confidence bounds is difficult.
+In addition, Thompson sampling is often observed to perform better than UCB in practice (Chapelle and Li, 2011; Wen et al., 2015; Kandasamy et al., 2018; Åkerblom et al., 2023b, a).
 Motivated by this, we present two algorithms for efficient prior selection based on TS.
-3.1 Prior-Elimination with Thompson sampling
 Algorithm 1 Prior Elimination GP-TS (PE-GP-TS)
@@
 3:  Sample $\tilde{f}_{t,p}\sim{\mathcal{GP}}(\mu_{t,p},k_{t,p})$ $\forall p\in P_{t}$
-4:  Set $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t-1}}\tilde{f}_{t,p}(x)$
+4:  Set $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t}}\tilde{f}_{t,p}(x)$
 5:  $S_{t,p_{t}}=S_{t-1,p_{t}}\cup\{t\}$ and $S_{t,p}=S_{t-1,p}$ for $p\in P\setminus\{p_{t}\}$
@@
 12:   $P_{t+1}=P_{t}$
+3.1 Prior-Elimination with Thompson sampling
 Our first algorithm is an extension of PE-GP-UCB (Ziomek et al., 2025) to be employed with Thompson sampling - instead of UCB.
-The key difference is that instead of maximizing the upper confidence bound $U_{t}(x,p)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$ over ${\mathcal{X}}\times P_{t-1}$, we instead sample $\tilde{f}_{t,p}$ from the posterior ${\mathcal{GP}}(\mu_{t,p},k_{t,p})$ for all priors $p\in P_{t-1}$ where $P_{t-1}$ is the set of active priors.
-Then, we select the arm and prior $x_{t},p_{t}$ such that $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t-1}}\tilde{f}_{t,p}(x)$.
+The key difference is that instead of maximizing the upper confidence bound $U_{t}(x,p)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$ over ${\mathcal{X}}\times P_{t}$, we instead sample $\tilde{f}_{t,p}$ from the posterior ${\mathcal{GP}}(\mu_{t,p},k_{t,p})$ for all priors $p\in P_{t}$ where $P_{t}$ is the set of active priors.
+Then, we select the arm and prior $x_{t},p_{t}$ such that $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t}}\tilde{f}_{t,p}(x)$.
 Whilst PE-GP-UCB has two layers of optimism, the upper confidence bound and joint maximization of $x$ and $p$, PE-GP-TS has only a single layer of optimism - which should alleviate potential overexploration issues.
-The elimination procedure of PE-GP-TS is illustrated in Fig.˜2.
-Samples $\tilde{f}_{t,p}$ are drawn from the active prior $p\in P_{t-1}$.
+The elimination procedure of PE-GP-TS is illustrated in Figure˜1.
+Samples $\tilde{f}_{t,p}$ are drawn from the active prior $p\in P_{t}$.
 Then, the unknown function $f$ is queried at the selected arm $x_{t}$.
@@
 3.2 HyperPrior Thompson sampling
+In our first algorithm, we removed one layer of optimism.
+In our second algorithm, we adopt a fully Bayesian algorithm by using a hyperposterior sampling scheme where both the prior and the mean function are sampled from their respective posteriors.
+By shedding the optimism over the selected prior $p_{t}$, HP-GP-TS should be able avoid costly exploration by selecting likely priors instead of optimistic ones.
+The algorithm is visualized in Figure˜2 and presented in detail in Algorithm˜2.
+In the first step, the current prior $p_{t}$ is sampled from the hyperposterior $P_{t}$.
+Then, a single sample $\tilde{f}_{t}$ is taken from the selected posterior ${\mathcal{GP}}(\mu_{t,p_{t}},k_{t,p_{t}})$ and is used to select the current arm: $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\tilde{f}_{t}(x)$.
+After observing $y_{t}$, the hyperposterior is updated by computing the likelihood of $y_{t}$ under the different priors.
+Note that since the set of priors $P$ is finite, computing the posterior is tractable albeit computationally costly with a complexity of ${\mathcal{O}}(t^{3}|P|)$.
+The algorithm can be extended to continuous priors $P$ using MCMC sampling.
+In comparison to SCoreBO and other fully Bayesian algorithms that compute expected values over the hyperposterior through sampling, HP-GP-TS requires only one sample from the posterior and hyperposterior – potentially reducing the computational cost significantly.
+The likelihood $\mathbb{P}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)={\mathcal{N}}(y_{t};\mu_{t,p}(x_{t}),\sigma_{t,p}^{2}(x_{t})+\sigma^{2})$ is simply the Gaussian likelihood of the posterior at $x_{t}$ with added Gaussian noise with variance $\sigma^{2}$.
+Figure 2: Overview of HP-GP-TS. The orange star represents $y_{t}$.
 Algorithm 2 HyperPrior GP-TS (HP-GP-TS)
@@
 6:  Set ${P}_{t+1}(p)\propto\mathbb{P}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)\cdot P_{t}(p)$ $\triangleright$ Update hyperposterior
-In our first algorithm, we removed one layer of optimism.
-In our second algorithm, we adopt a fully Bayesian algorithm by using a hyperposterior sampling scheme where both the prior and the mean function are sampled from their respective posteriors.
-By shedding the optimism over the selected prior $p_{t}$, HP-GP-TS should be able avoid costly exploration by selecting likely priors instead of optimistic ones.
-The algorithm is visualized in Fig.˜2 and presented with more details in Algorithm˜2.
-In the first step, the current prior $p_{t}$ is sampled from the hyperposterior $P_{t-1}$.
-Then, a single sample $\tilde{f}_{t}$ is taken from the selected posterior ${\mathcal{GP}}(\mu_{t,p_{t}},k_{t,p_{t}})$ and is used to select the current arm: $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\tilde{f}_{t}(x)$.
-After observing $y_{t}$, the hyperposterior is updated by computing the likelihood of $y_{t}$ under the different priors.
-Note that since the set of priors $P$ is finite, computing the posterior is tractable albeit computationally costly with a complexity of ${\mathcal{O}}(t^{3}|P|)$.
-The likelihood $\mathbb{P}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)={\mathcal{N}}(y_{t};\mu_{t,p}(x_{t}),\sigma_{t,p}^{2}(x_{t})+\sigma^{2})$ is simply the Gaussian likelihood of the posterior at $x_{t}$ with added Gaussian noise with variance $\sigma^{2}$.
 4 Regret analysis
@@
 Ziomek et al.
-(2025) structured the proof of the regret bound of PE-GP-UCB into 4 larger steps; First, showing that $p^{*}$ is never eliminated with high probability.
+(2025) structured the proof of the regret bound for PE-GP-UCB into 4 larger steps; First, showing that $p^{*}$ is never eliminated with high probability.
 Second, establishing a bound on the simple regret.
@@
 Section˜4.1 is based on Lemma 5.1 of Srinivas et al.
-(2012) but adapted to TS by specifying that it holds for any sequence of $x_{1},\ldots,x_{T}$, as discussed by Russo & Van Roy (2014).
-Additionally, we add Eq.˜5 which can be shown through the same steps and an additional union bound over $P$.
+(2012) but adapted to TS by specifying that it holds for any sequence of $x_{1},\ldots,x_{T}$, as discussed by Russo and Van Roy (2014).
+Additionally, we add Equation˜5 which can be shown through the same steps and an additional union bound over $P$.
 Next, we state our bound for the simple regret of PE-GP-TS.
@@
 If $p^{*}\in P$ and $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, then PE-GP-TS with confidence parameters $\beta_{t}=2\log(2|{\mathcal{X}}||P|\pi^{2}t^{2}/3\delta)$ and $\xi_{t}=2\sigma^{2}\log(|P|\pi^{2}t^{2}/3\delta)$, satisfies the following regret bound with probability at least $1-\delta$:
-$R(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}+2\sqrt{CT\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}$ (7)
+$R(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}\\ +2\sqrt{T\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}$ (7)
 The bound of the first three terms is of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma}_{T}})$ w.r.t.
@@
 To our knowledge, the best lower bound for standard GP bandits in the Bayesian setting, where $f$ is sampled from a GP, is $\Omega(\sqrt{T})$ for $d=1$ (Scarlett, 2018).
-This would suggest that our bound is tight up to a factor ${\mathcal{O}}(\sqrt{\beta_{T}\hat{\gamma}_{T}})$.
+This would suggest that our bound is tight up to a factor ${\mathcal{O}}(\sqrt{\beta_{T}\hat{\gamma}_{T}})$ when considering only the first three terms.
+However, note that the sublinearity of ${\sum_{t\in[T]}}\sigma_{t,p^{*}}^{2}(x^{*})$ is not demonstrated.
 4.2 Analysis of HP-GP-TS
-We analyze the regret of HP-GP-TS in two ways: using the UCB-based framework of Russo & Van Roy (2014) and the information-theoretic framework of Russo & Van Roy (2016).
+We analyze the regret of HP-GP-TS using the UCB-based framework of Russo and Van Roy (2014).
 First, note that HP-GP-TS inherits the probability matching property of GP-TS that $x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t}$ where $\overset{d}{=}$ denotes equal in distribution.
 In addition, $p_{t}|H_{t}\overset{d}{=}p^{*}|H_{t}$ since $p_{t}$ is sampled from the posterior distribution of $p^{*}$.
-In the UCB-based framework, we decompose the regret into three terms:
-${\mathbb{E}}[f(x^{*})-f(x_{t})]={\mathbb{E}}\Big[\underbrace{f(x^{*})-U_{t,p^{*}}(x^{*})}_{(1)}+\underbrace{U_{t,p^{*}}(x^{*})-U_{t,p^{*}}(x_{t})}_{(2)}+\underbrace{U_{t,p^{*}}(x_{t})-f(x_{t})}_{(3)}\Big]$ (8)
+We decompose the regret into three terms:
+${\mathbb{E}}[f(x^{*})-f(x_{t})]={\mathbb{E}}\Big[\underbrace{f(x^{*})-U_{t,p^{*}}(x^{*})}_{(1)}\\ +\underbrace{U_{t,p^{*}}(x^{*})-U_{t,p^{*}}(x_{t})}_{(2)}+\underbrace{U_{t,p^{*}}(x_{t})-f(x_{t})}_{(3)}\Big]$ (8)
 where $U_{t,p}(x)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$.
 Summing over $t\in[T]$, the first term can be bounded by a constant whilst the third term is bounded by $\sqrt{CT\beta_{T}\bar{\gamma}_{T}}$.
-Together, these two terms match the Bayesian regret bound for GP-TS with known prior.
-For the second term, one can utilize that $p^{*},x_{t}|H_{t}\overset{d}{=}p_{t},x^{*}|H_{t}$ to re-express it as ${\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$.
+Together, these two terms match the Bayesian regret bound for GP-TS with a known prior.
+For standard GP-TS when the true prior $p^{*}$ is known, the second term has expectation zero.
+However for HP-GP-TS, one can utilize that $p^{*},x_{t}|H_{t}\overset{d}{=}p_{t},x^{*}|H_{t}$ to re-express the second term as ${\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$.
 Hence, the second term can be seen as the cost of learning the true prior.
@@
 {lemma}[] If $|P|=2$ and the two priors share the same kernel function ($k_{p}=k$ $\forall p\in P$), then for any fixed sequence of arms $x_{1:t}=\{x_{i}\}_{i=1}^{t}$ the posterior probability of the true prior $p^{*}$ satisfies
-${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|p^{*}=p,x_{1:t}\right]\geq 1+P_{0}(p)e^{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}\Phi\Big(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\Big)-\frac{1}{P_{0}(p)}\Phi\Big(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\Big)$ (9)
+${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|p^{*}=p,x_{1:t}\right]\geq 1+P_{0}(p)e^{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}\\ \times\Phi\Big(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\Big)-\frac{1}{P_{0}(p)}\Phi\Big(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\Big)$ (9)
 where ${\bm{\mu}}\in{\mathbb{R}}^{t}$ such that $({\bm{\mu}})_{i}=\mu_{p^{*}}(x_{i})-\mu_{\tilde{p}}(x_{i})$ for $\tilde{p}\neq p^{*}$ and $({\bm{\Sigma}})_{i,j}=k(x_{i},x_{j})$.
-From Eq.˜8, we note that $P_{t}(p^{*})$ increases the most when $\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$ is maximized, which corresponds to selecting arms with low correlation and large difference in prior mean (relative to the variance).
+In Equation˜8, $\bm{\mu}$} is the difference in mean between the two priors for the given sequence of arms and ${\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}$ is the whitened difference vector.
+As the norm of the whitened difference vector increases, $\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$, the sum of the two rightmost terms in Equation˜9 goes to zero from below.
+The norm $\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$, and consequently the posterior probability of the true prior $P_{t+1}(p^{*})$, is maximized when arms with large difference in prior mean (relative to their variance) and with low correlation are selected.
 Hence, $P_{t}(p^{*})$ increases quickly at first when HP-GP-TS explores but levels out when the algorithm starts exploiting.
@@
 If $p^{*}\sim P_{1}$, $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ and $\beta_{t}=2\log(|{\mathcal{X}}|t^{2}/\sqrt{2\pi)}$, then the Bayesian regret of HP-GP-TS is bounded by
-${\text{BR}}(T)\leq\pi^{2}/6+{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})\right]+\sqrt{CT\beta_{T}\bar{\gamma}_{T}}.$ (10)
+${\text{BR}}(T)\leq\pi^{2}/6+{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})\right]\\ +\sqrt{CT\beta_{T}\bar{\gamma}_{T}}.$ (10)
 Unlike PE-GP-TS and PE-GP-UCB, the regret bound of the third term for HP-GP-TS depends on the average MIG $\sqrt{\bar{\gamma}_{T}}$ rather than the worst case $\sqrt{|P|\hat{\gamma}_{T}}$ which can impact the theoretical regret significantly if the complexity of the priors differ and the prior is weighted towards simple priors.
@@
 If the hyperprior is deterministic then the regret bound for HP-GP-TS matches that of GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ (Takeno et al., 2024) and $\bar{\gamma}_{T}$ would be equal to the worst case $\hat{\gamma}_{T}$.
-Again, using the lower bound of Scarlett (2018), our upper bound would be tight up to a factor of ${\mathcal{O}}(\sqrt{\beta_{T}\bar{\gamma}_{T}})$.
-The information-theoretic framework of Russo & Van Roy (2016) can be applied generally if the probability matching property is satisfied and the rewards are subgaussian (Vashishtha & Maillard, 2025).
-Theorem 4.3.
-If $p^{*}\sim P_{1}$, $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, the Bayesian regret of HP-GP-TS is bounded by
-${\text{BR}}(T)\leq\sqrt{2|{\mathcal{X}}|\log(|{\mathcal{X}}|)(\sigma^{2}_{0}+\sigma^{2})T}.$ (11)
-The proof of Theorem˜4.3 follows the proof in section D.2 of Russo & Van Roy (2016) with subgaussian noise.
-For completeness, we provide a proof using the GP-bandit notation in Algorithm˜2.
-The information-theoretic regret bound is ${\mathcal{O}}(\sqrt{T})$ which improves upon the ${\mathcal{O}}(\sqrt{T\beta_{T}\bar{\gamma}_{T}})$ obtained previously and would match the lower bound of Scarlett (2018) in terms of $T$.
-However, we also obtain a ${\mathcal{O}}(\sqrt{|{\mathcal{X}}|\log|{\mathcal{X}}|})$ dependency.
+Again, using the lower bound of Scarlett (2018), our upper bound would be tight up to a factor of ${\mathcal{O}}(\sqrt{\beta_{T}\bar{\gamma}_{T}})$ when considering only the first and third terms.
+We have not shown that ${\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$ is sublinear.
+As noted before, this sum corresponds to the cost of learning the true prior.
+Conceptually, we expect the terms in the sum to decrease over time as either $p_{t}$ or $x_{t}$ converge to $p^{*}$ or $x^{*}$ respectively.
+Empirically, we observe this in our synthetic experiments, see Figure˜13 in Appendix˜F.
+4.3 Comparison to MixTS
+Hong et al.
+(2022a) study MixTS, a Thompson sampling algorithm that assumes the prior is a mixture distribution, for standard and linear bandits.
+Using our GP-notation, their analysis starts from the decomposition
+${\mathbb{E}}[f(x^{*})-f(x_{t})]={\mathbb{E}}[\underbrace{f(x^{*})-\mu_{t,p^{*}}(x^{*})}_{(1)}\\ +\underbrace{\mu_{t,p_{t}}(x_{t})-f(x_{t})}_{(2)}].$ (11)
+The first term is bounded using standard UCB techniques.
+For the second term, Hong et al.
+(2022a) introduce confidence sets for the priors $C_{t}$ based on the excess reward each prior observes and prove that $p^{*}\in C_{t}$ with high probability.
+Their key idea is to introduce confidence sets for the priors $C_{t}$ based on the excess reward each prior observes and prove that $p^{*}\in C_{t}$ holds with high probability, similar to Ziomek et al.
+(2025).
+In our GP setting, their excess reward can be defined as
+$G_{t}(p)=\sum_{s=1}^{t-1}\mathds{1}\{p_{s}=p\}(\mu_{s,p}(a_{s})-\sqrt{\beta_{t}}\sigma_{s,p}(a_{s})-y_{s})$ (12)
+and the prior confidence set $C_{t}=\{p\in P:G_{t}(p)\leq 2\sigma\sqrt{N_{t}(p)\log T}\}$ where $N_{t}(p)$ is the number of times prior $p$ has been selected before time $t$.
+Unlike the definition of Ziomek et al.
+(2025), priors can return to the confidence set after previously being removed from it.
+When bounding the second term in their decomposition for a standard bandit with bounded rewards, Hong et al.
+(2022a) condition on $p_{t}\in C_{t}$.
+If $p_{t}\in C_{t}$, the second term can be bounded by all the excess $\sum_{p\in P}G_{T+1}(p)$.
+If $p_{t}\notin C_{t}$ and the rewards are in $[0,1]$, then the second term is bounded by the probability of $p^{*}\in C_{t}$.
+For the linear setting with unbounded rewards, the proof requires conditioning on the linear parameter vector $\theta^{*}$ to lie close to its prior mean.
+Under this additional event $E_{0}$, the distribution of the true parameter vector $\theta^{*}$, the true prior and the optimal arm can shift.
+But, conditioned on the history $H_{t}$, MixTS is unaffected by $E_{0}$ at time step $t$.
+However, Hong et al.
+(2022a) use that $x^{*},p^{*}|H_{t},E_{0}\overset{d}{=}x_{t},p_{t}|H_{t},E_{0}$ without proof, potentially invalidating Theorem 1 of Hong et al.
+(2022a).
+We describe this and other technical issues in detail in Appendix˜B.
 Figure 3: Cumulative regret for synthetic experiments with varying kernel (left), lengthscale (center) and mean function (right).
-The average final regret for PE-GP-UCB is 116.5 and 389.0 in the lengthscale and subspace experiments.
+The final regret for PE-GP-UCB is 114 and 389 in the lengthscale and subspace experiments, and 181 for SCoreBO in the lengthscale experiment.
 Errorbars correspond to $\pm 1$ standard error.
+5 Experiments
+In this section, we describe our experiments based on synthetic and real-world data.
+Synthetic experiments
+We consider three synthetic setups with different choices of priors in $P$.
+For the first setup, the priors have one of the following kernels: i) RBF kernel, ii) the rational quadratic kernel with $\alpha=0.5$, iii) Matérn kernel with $\nu=5/2$, iv) Matérn kernel with $\nu=3/2$, v) periodic kernel with period $\rho=5$, vi) linear kernel with $v=0.05^{2}$.
+For the second setup, 8 priors use the RBF kernel with different lengthscales equidistantly spaced between $1/2$ and $4$.
+For the third setup, the total dimensions $d=16$ but each of the 5 priors $p_{i}$ assumes $f(x)$ depends on $d_{s}=4$ subdimensions.
+The 4 subdimensions are designed such that the priors are equally difficult to distinguish.
+All priors use the RBF kernel with lengthscale $\ell=8$.
+For all three setups, the true prior $p^{*}$ is sampled uniformly from $P$, the noise variance $\sigma^{2}=0.25^{2}$, and the horizon $T=500$.
+For the first two setups, 500 arms are equidistantly spaced in $[0,20]$ and for the third 500 arms are sampled uniformly on $0,20]^{16}$.
+All models are evaluated on 500 seeds on each setup.
+As baselines, we use PE-GP-UCB, SCoreBO, fully Bayesian Expected Improvement (EEI) and Maximum A Posteriori (MAP) GP-TS.
+MAP GP-TS is identical to HP-GP-TS except for greedily selecting $p_{t}$ from the posterior: $p_{t}=\operatorname*{arg\,max}_{p}P_{t-1}(p)$ [^1] Note that since the hyperprior is uniform, MAP is equivalent to discrete maximum likelihood estimation..
+In addition, we compare against the oracle variants of PE-GP-TS and PE-GP-UCB that are only given the true prior: $P_{1}=\{p^{*}\}$.
+The cumulative regret for the three synthetic experiments is shown in Figure˜3 and the final regret is shown in Table˜1 in Appendix˜F.
+Across all three experiments, we observe that HP-GP-TS and EEI has lower regret than the other methods and performs close to the oracle GP-TS.
+For the kernel and subspace experiments, PE-GP-TS has lower regret than the oracle GP-UCB.
+Hence, even if PE-GP-UCB was optimized to perform as well as the oracle, it would still not achieve the regret of our proposed methods.
+MAP GP-TS has slightly higher regret than HP-GP-TS for the lengthscale and subspace experiments but has significantly higher regret and variance for the kernel experiment.
+The greedy selection of MAP (MLE) leads to under-exploration for MAP GP-TS in certain instances.
+SCoreBO has the highest regret in the kernel and lengthscale experiment but has more comparable performance in the subspace experiment.
+In Figure˜10 in Appendix˜F, we report the regret for the two most competitive methods, HP-GP-TS and EEI, with an extended horizon $T=1500$ where we observe that HP-GP-TS yields noticeably lower regret.
+The number of priors remaining $|P_{t}|$ and the hyperposterior entropy for the kernel experiment is shown in Figure˜4.
+The PE-methods eliminate at most one prior on average.
+In contrast, the final hyperposterior entropy across all algorithms is equivalent to 70-99% of the probability mass being assigned to one prior showing that the hyperposterior adapts more effectively.
+Across the experiments, SCoreBO has the lowest hyperposterior entropy followed by HP-GP-TS and EEI has the highest (except for the kernel experiment), see Figures˜11 and 12 in Appendix˜F.
+Thus, HP-GP-TS has similar regret to EEI but lower hyperposterior entropy.
+In Figure˜5, we visualize how often the methods select the true prior $p^{*}$ (or kernel) in the kernel experiment as confusion matrices.
+PE-GP-UCB selects the Matérn-3/2 kernel more than 96% of the rounds.
+The Matérn-3/2 kernel induces a distribution over functions that are less smooth compared to the other kernels and produces much higher confidence intervals outside the observed data leading to excessive optimistic exploration.
+PE-GP-TS also shows a bias towards the Matérn-3/2 kernel but does not select it as frequently as PE-GP-UCB - demonstrating that one layer of optimism has been removed.
+The overall “accuracy" of the selected priors, i.e.
+${\sum_{t\in[T]}}\mathds{1}\{p_{t}=p^{*}\}/T$, for the elimination-based methods is around $17\%$ in the kernel experiment compared to $62.9\%$ and $63.2\%$ for MAP and HP-GP-TS respectively.
+For HP-GP-TS, we observe that it can easily identify the periodic and linear kernels.
+However, the RBF, Matérn and RQ kernels are often confused with each other.
+These kernels do not have as easily distinguishable characteristics and are likely to produce similar posteriors even with a small amount of data.
+See Figure˜14 in Appendix˜F for confusion matrices in the lengthscale and subspace experiments.
 Figure 4: Mean number of priors remaining in $P_{t}$ over time for PE-GP-UCB and -TS (left).
@@
 $\frac{1-q}{|P|-1}$ on the other |P|1|P|-1 choices.
-5 Experiments
-Synthetic experiments
-We consider three synthetic setups with different choices of priors in $P$.
-For the first setup, the priors have one of the following kernels: i) RBF kernel, ii) the rational quadratic kernel with $\alpha=0.5$, iii) Matérn kernel with $\nu=5/2$, iv) Matérn kernel with $\nu=3/2$, v) periodic kernel with period $\rho=5$, vi) linear kernel with $v=0.05^{2}$.
-All kernels use a lengthscale of $1.0$ and are scaled s.t.
-$k(x,\tilde{x})\leq 1$.
-In addition, the mean function for all priors is zero everywhere.
-For the second setup, the priors use the RBF kernel with lengthscales $4$, 22, 11 or $1/2$.
-For the third setup, the total dimensions $d=16$ but each prior $p_{i}$ assumes $f(x)$ depends on $d_{s}=4$ subdimensions: $[i,i+1,i+2,i+3]$ for $i\in[5]$.
-Dimensions larger than $5$ are wrapped around 1, i.e.
-$((j-1)\mod 5)+1$, such that the priors are equally difficult to distinguish and optimize.
-All priors use the RBF kernel with lengthscale $\ell=8$.
-For all three setups, the true prior $p^{*}$ is sampled uniformly from $P$, the noise variance $\sigma^{2}=0.25^{2}$, and the horizon $T=500$.
-For the first two setups, 500 arms are equidistantly spaced in $[0,20]$ and for the third 500 arms are sampled uniformly on $0,20]^{16}$.
-The prior elimination methods use $\delta=0.05$.
-All models are evaluated on 500 seeds on each setup.
-As baselines, we use PE-GP-UCB and Maximum A Posteriori (MAP) GP-TS where MAP GP-TS is identical to HP-GP-TS except for greedily selecting $p_{t}$ from the posterior: $p_{t}=\operatorname*{arg\,max}_{p}P_{t-1}(p)$ [^1] Note that since the hyperprior is uniform, MAP is equivalent to discrete maximum likelihood estimation..
-Regardless of the selected prior, the weighting of the hyperposterior $P_{t}$ is updated for all priors.
-Hence, greedily selecting the prior could reduce unnecessary exploration.
-In addition, we investigate the oracle variants of PE-GP-TS and PE-GP-UCB with $\delta=0.05$ that are only given the true prior: $P_{1}=\{p^{*}\}$.
-The cumulative regret for the three synthetic experiments is shown in Fig.˜3.
-Across all three experiments, we observe that HP-GP-TS has lower regret than the other methods and performs close to the oracle GP-TS.
-For the kernel and subspace experiments, PE-GP-TS has lower regret than the oracle GP-UCB.
-Hence, even if PE-GP-UCB was optimized to perform as well as the oracle, it would still not achieve the regret of our proposed methods.
-MAP GP-TS has slightly higher regret than HP-GP-TS for the lengthscale and subspace experiments but has significantly higher regret and variance for the kernel experiment.
-The greedy selection of MAP (MLE) leads to under-exploration for MAP GP-TS in certain instances.
-The number of priors remaining $|P_{t}|$ and the hyperposterior entropy for the kernel experiment is shown in Fig.˜4.
-The PE-methods eliminate at most one prior on average.
-In contrast, the hyperposterior entropy of HP-GP-TS is equivalent to 80-90% of the probability mass being assigned to one prior.
-HP- and MAP-GP-TS thus effectively discards priors at a much faster rate.
-The same pattern holds for the lengthscale and subspace experiments, see Figs.˜9 and 10 in Appendix˜D.
 Figure 5: Confusion matrices for the true prior $p^{*}$ and the selected priors $p_{t}$ for the kernel experiment. Row-wise normalized to 100%.
-In Fig.˜5, we visualize how often the methods select the true prior $p^{*}$ (or kernel) in the kernel experiment as confusion matrices.
-PE-GP-UCB selects the Matérn-3/2 kernel more than 96% of the rounds.
-The Matérn-3/2 kernel induces a distribution over functions that are less smooth compared to the other kernels and produces much higher confidence intervals outside the observed data leading to excessive optimistic exploration.
-PE-GP-TS also shows a bias towards the Matérn-3/2 kernel but does not select it as frequently as PE-GP-UCB - demonstrating that one layer of optimism has been removed.
-The overall “accuracy" of the selected priors, i.e.
-${\sum_{t\in[T]}}\mathds{1}\{p_{t}=p^{*}\}/T$, for the elimination-based methods is around $17\%$ in the kernel experiment compared to $62.5\%$ and $63.2\%$ for MAP and HP-GP-TS respectively.
-For HP-GP-TS, we observe that it can easily identify the periodic and linear kernels.
-However, the RBF, Matérn and RQ kernels are often confused with each other.
-These kernels do not have as easily distinguishable characteristics and are likely to produce similar posteriors even with a small amount of data.
-See Fig.˜11 in Appendix˜D for confusion matrices in the lengthscale and subspace experiments.
-Figure 6: Total regret for the lengthscale and subspace experiments as the number of priors $|P|$ increases.
+Figure 6: Total regret for the lengthscale and subspace experiments as $|P|$ increases.
+Figure 7: Cumulative regret on the real-world data experiments. Errorbars correspond to $\pm 1$ standard error. The average final regret for SCoreBO and PE-GP-UCB is 861 and 506 on PNW.
 Scaling $|P|$
@@
 Each prior can share at most 3 out of 4 dimensions with other priors which ensures the priors remain meaningfully different.
-The total regret as the number of priors increases is shown in Fig.˜6.
-For the lengthscale experiment, increasing the number of priors above 8 does not affect the regret for any algorithm.
-This is likely due to the redundancy in the priors.
-However in the subspace experiment, the regret of the prior elimination algorithms scales approximately as $\sqrt{|P|}$ whilst MAP- and HP-GP-TS are consistently close to the constant regret of the oracle.
-Figure 7: Cumulative regret on Intel temperature data (top) and PeMS data (bottom). Errorbars correspond to $\pm 1$ standard error. The average final regret for PE-GP-UCB is 511.9 for PNW.
+The total regret as the number of priors increases is shown in Figure˜6.
+For the lengthscale experiment, increasing the number of priors above 8 does not affect the regret for any algorithm, likely due to the increased redundancy in the priors.
+The one exception is SCoreBO, whose regret increases as $|P|$ is increased to 32 but levels off beyond that.
+In the subspace experiment, the regret of the prior elimination algorithms scales approximately as $\sqrt{|P|}$ whilst MAP- and HP-GP-TS are consistently close to the constant regret of the oracle.
+The regret of EEI drops initially but is otherwise constant and SCoreBO also has constant regret.
 Real-world data
-We perform three experiments with real-world data from the Intel Berkeley dataset (Madden et al., 2004), California Performance Measurement System (PeMS) (Chen et al., 2001; California Department of Transportation, 2024) and Pacific Northwest (PNW) daily precipitation dataset (Widmann & Bretherton, 1999, 2000).
+We perform three experiments with real-world data from the Intel Berkeley dataset (Madden and others, 2004), California Performance Measurement System (PeMS) (Chen et al., 2001; California Department of Transportation, 2024) and Pacific Northwest (PNW) daily precipitation dataset (Widmann and Bretherton, 1999, 2000).
 Each dataset contains measurements from a set of sensors over time.
@@
 The buckets in the Intel data corresponds to the 12 days in the training dataset.
-For the PeMS data, each hour between 06:00 and 13:00 defines one prior, giving 7 priors.
-For the daily precipication data, each month in the year constitutes a prior, yielding 12 priors.
+For the PeMS data, each hour between 06:00 and 13:00 defines one bucket, giving 7 priors.
+For the daily precipitation data, each month in the year constitutes a bucket, yielding 12 priors.
 When running the experiments, we select a measurement of all sensors from the test data uniformly at random.
-The selected measurements correspond to the unknown function $f(x)$ where $x$ is the sensor index and the goal is then to identify sensors measuring large temperatures, small speeds or high precipication respectively for the three datasets.
-When the algorithms select an arm to evaluate, we add Gaussian noise to $y_{t}$ with variance $\sigma^{2}$ around 5% of the signal variance, similar to Srinivas et al.
+The selected measurements correspond to the unknown function $f(x)$ where $x$ is the sensor index and the goal is then to identify sensors measuring large temperatures, small speeds or high precipitation respectively for the three datasets.
+When the algorithms select an arm to evaluate, we add Gaussian noise with variance $\sigma^{2}$ around 5% of the signal variance, similar to Srinivas et al.
 (2012); Bogunovic et al.
 (2016).
-See Appendix˜C for more details about the experimental setup.
-The cumulative regret for the experiments with real-world data is presented in Fig.˜7.
-For the Intel and PNW data, HP- and MAP GP-TS obtain the lowest and second lowest cumulative regret respectively.
-HP- and MAP GP-TS have lower regret than PE-GP-TS initially but PE-GP-TS catches up and has the lowest total regret.
-To understand this better, we visualize quantiles of the total regret in Fig.˜13.
-MAP- and HP-GP-TS have the lowest median regret for all three experiments and hence perform best in a majority of instances.
-However, the 90th and 95th quantiles are considerably larger for the PeMS data which impacts the average regret significantly.
-Hence, for the PeMS data, the prior elimination methods seem to yield more stable results.
-In Fig.˜4, the number of priors remaining in $|P_{t}|$ and the hyperposterior entropy is shown for the Intel experiment.
-Similar to the synthetic experiment, on average, the prior elimination methods eliminate less than 1 prior whereas the hyperposteriors of HP- and MAP GP-TS concentrate the equivalent of 80-90% of the probability mass to one prior.
-The results for PeMS and PNW experiment are shown in Figs.˜9 and 10 in Appendix˜D.
-Here, effectively no priors are eliminated.
-For the PNW experiment, the hyperposteriors do not concentrate as much compared to the other experiments.
-This could indicate that knowing the exact prior is not as important for the PNW data.
+See Appendix˜E for details about the experimental setup.
+The cumulative regret for the experiments with real-world data is presented in Figure˜7.
+Across these experiments, HP-GP-TS has either the lowest regret or is within 1 standard error of the algorithm with the lowest regret.
+SCoreBO has significantly higher regret than all other methods.
+Notably for the PeMS data, PE-GP-TS has the lowest average regret whereas MAP GP-TS and EEI perform worse compared to the other experiments.
+The number of priors remaining in $|P_{t}|$ and the hyperposterior entropy for the real-world data experiments is shown in Figures˜11 and 12 in Appendix˜F.
+Similar to the synthetic experiments, on average, the prior elimination methods eliminate less than 1 prior at best and no priors (across all 500 seeds) at worst.
+In contrast, the hyperposterior of HP-GP-TS concentrates to the equivalent of 60-80% of the probability mass to one prior.
+The relative standing in terms of reduced hyperposterior uncertainty between SCoreBO, HP-GP-TS and EEI remains consistent across all the experiments.
+SCoreBO reduces the hyperposterior uncertainty the most at the cost of significantly higher regret whereas HP-GP-TS provides a better balance between low regret and low hyperposterior uncertainty.
 6 Conclusion
-In this paper, we have proposed two algorithms for joint prior selection and regret minimization in GP bandits based on GP-TS.
+In this paper, we have proposed two algorithms for adaptive prior selection and regret minimization in GP bandits based on GP-TS.
 We have analyzed the algorithms theoretically and have experimentally evaluated both algorithms on synthetic and real-world data.
-We find that they both select the true prior more often and obtain lower regret than previous work due to lowering the amount of optimistic exploration.
+We find that they both select the true prior more often and obtain lower or comparable regret than previous work due to lowering the amount of optimistic exploration.
 Acknowledgments
@@
 Follows by the same steps as Lemma 5.1 of Srinivas except we condition on the complete history $H_{t}$ instead of only ${\mathbf{y}}_{1:t-1}$.
-Additionally, for Eq.˜5 we must take an additional union bound over $p\in P$.
+Additionally, for Equation˜5 we must take an additional union bound over $p\in P$.
 Fix $t,x,p\in[T]\times{\mathcal{X}}\times P$.
@@
 Using that $\mathbb{P}(Z>c)\leq 1/2e^{-c^{2}/2}$ for $Z\sim{\mathcal{N}}(0,1)$, we get that
-$\mathbb{P}\left(\left|\frac{\tilde{f}_{t,p}(x)-\mu_{t,p^{*}}(x)}{\sigma_{t,p^{*}}(x)}\right|>\sqrt{\beta_{t}}\right)$ $\leq\exp(-\beta_{t}/2)$ (12)
+$\mathbb{P}\left(\left|\frac{\tilde{f}_{t,p}(x)-\mu_{t,p}(x)}{\sigma_{t,p}(x)}\right|>\sqrt{\beta_{t}}\right)$ $\leq\exp(-\beta_{t}/2)$ (13)
 $=\frac{3\delta}{|{\mathcal{X}}||P|\pi^{2}t^{2}}$ (14)
-Note that $\sum_{t\geq 1}\frac{1}{t^{2}}=\frac{\pi^{2}}{6}$By taking the union bound over $\mathcal{X}$}, $P$ and $t\geq 1$, Eq.˜5 holds w.p.
+Note that $\sum_{t\geq 1}\frac{1}{t^{2}}=\frac{\pi^{2}}{6}$By taking the union bound over $\mathcal{X}$}, $P$ and $t\geq 1$, Equation˜5 holds w.p.
 at least $1-\delta/2$.
-By the same reasoning and skipping the union bound over $P$, Eq.˜4 holds w.p.
+By the same reasoning and skipping the union bound over $P$, Equation˜4 holds w.p.
 at least $1-\delta/2$.
@@
 […]
-Combining, Eqs.˜18 and 20 we obtain
+Combining, Equations˜19 and 21 we obtain
 $f(x^{*})-f(x_{t})\leq 2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-\eta_{t}+\epsilon_{t}.$ (22)
@@
 $\mathcal{C}=\left\{t\in[T]:\left|\sum_{i\in S_{t,p_{t}}}\eta_{i}\right|>\sqrt{\xi_{t}S_{t,p_{t}}}+\sum_{i\in S_{t,p_{t}}}\sqrt{\beta_{i}}\sigma_{i,p_{t}}(x_{i})\right\}.$ (26)
-Using Eqs.˜14 and 21, we can bound the cumulative regret as follows:
+Using Equations˜15 and 22, we can bound the cumulative regret as follows:
 ${\text{BR}}(T)$ $=\sum_{t\in\mathcal{C}}{\text{BR}}_{t}+\sum_{t\notin\mathcal{C}}{\text{BR}}_{t}$ (27)
@@
 ${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sum_{t\notin\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})+2\sum_{t\notin\mathcal{C}}\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t}).$ (33)
-Finally, applying Eq.˜15, we obtain the result
-${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}.$ (33)
+Finally, applying Equation˜16, we obtain the result
+${\text{BR}}(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{T\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}.$ (34)
 A.2 UCB-analysis of HP-GP-TS
@@
 $\leq{\sum_{t\in[T]}}\sum_{x\in{\mathcal{X}}}{\mathbb{E}}_{p^{*},H_{t}}\left[{\mathbb{E}}_{t}\left[\left[f(x)-\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)\right]_{+}\big|\,p^{*},H_{t}\right]\right]$ $(\text{Tower rule})$ (42)
-Recall that for $Z\sim{\mathcal{N}}(\mu,\sigma)$ with $\mu\leq 0$, ${\mathbb{E}}[[Z]_{+}]=\frac{\sigma}{\sqrt{2\pi}}\exp\left(\frac{-\mu^{2}}{2\sigma^{2}}\right)$.
+Recall that for $Z\sim{\mathcal{N}}(\mu,\sigma)$ with $\mu\leq 0$, ${\mathbb{E}}[[Z]_{+}]\leq\frac{\sigma}{\sqrt{2\pi}}\exp\left(\frac{-\mu^{2}}{2\sigma^{2}}\right)$.
 In our case, note that $f(x)|p^{*},H_{t}\sim{\mathcal{N}}(\mu_{t,p^{*}}(x),\sigma^{2}_{t,p^{*}}(x))$ and $\mu_{t,p^{*}}(x)-\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x)$ is deterministic given $p^{*},H_{t}$.
@@
 $P_{t+1}(p)$ $=\frac{N_{p}({\mathbf{y}})P_{0}(p)}{N_{p}({\mathbf{y}})P_{0}(p)+N_{\tilde{p}}({\mathbf{y}})P_{0}(\tilde{p})}=\frac{\frac{N_{p}({\mathbf{y}})}{N_{\tilde{p}}({\mathbf{y}})}P_{0}(p)}{\frac{N_{p}({\mathbf{y}})}{N_{\tilde{p}}({\mathbf{y}})}P_{0}(p)+(1-P_{0}(p))}.$ (60)
-Note that $N_{p}({\mathbf{y}})/N_{\tilde{p}}({\mathbf{y}})=\exp\left(-\frac{1}{2}\left({\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}}-({\mathbf{y}}^{T}-{\bm{\mu}}){\bm{\Sigma}}^{-1}({\mathbf{y}}-{\bm{\mu}})\right)\right)=:\exp(Q({\mathbf{y}}))$ and that the quotient in the integrand can be written as $q(x)=\frac{e^{x}c}{e^{x}c+(1-c)}$ for $c:=P_{0}(p)\in(0,1)$.
+Note that $N_{p}({\mathbf{y}})/N_{\tilde{p}}({\mathbf{y}})=\exp\left(-\frac{1}{2}\left({\mathbf{y}}^{T}{\bm{\Sigma}}^{-1}{\mathbf{y}}-({\mathbf{y}}-{\bm{\mu}})^{T}{\bm{\Sigma}}^{-1}({\mathbf{y}}-{\bm{\mu}})\right)\right)=:\exp(Q({\mathbf{y}}))$ and that the quotient in the integrand can be written as $q(x)=\frac{e^{x}c}{e^{x}c+(1-c)}$ for $c:=P_{0}(p)\in(0,1)$.
 Since we wish to lower bound the expectation of $P_{t+1}(p)$, we will utilize the lower bounds $q(x)\geq ce^{x}$ $\forall x\leq 0$ and $q(x)\geq 1-\frac{1-c}{c}e^{-x}$ $\forall x\in{\mathbb{R}}$.
@@
 ${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|x_{1:t}\right]$ $\geq 1+P_{0}(p)e^{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}\Phi\left(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\right)-\frac{1}{P_{0}(p)}\Phi\left(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\right)$ (78)
-Note that a sharp bound of Eq.˜77 in terms of elementary functions can be obtained using bounds on the error function (Cook, 2018; Abramowitz & Stegun, 1972):
+Note that a sharp bound of Equation˜78 in terms of elementary functions can be obtained using bounds on the error function (Cook, 2018; Abramowitz and Stegun, 1972):
 ${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|x_{1:t}\right]$ $\geq 1+P_{0}(p)\sqrt{\frac{2}{\pi}}e^{-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}{8}}\frac{2}{3\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert+\sqrt{9\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}+16}}$ (79)
 $\quad-\frac{1}{P_{0}(p)}\sqrt{\frac{2}{\pi}}e^{-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}{8}}\frac{2}{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert+\sqrt{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}+32\pi^{-1}}}.$ (80)
-A.4 Information-theoretic regret bound for GP-TS
-We begin by showing that the rewards are subgaussian.
-{lemma}[] Fix $x\in{\mathcal{X}}$ and let $Z=f(x)+\epsilon_{t}-{\mathbb{E}}_{t}[f(x)]$.
-If $p^{*}\sim P_{1}$, $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ with $k_{1,p^{*}}(\cdot,\cdot)\leq\sigma_{0}^{2}$, then $Z|H_{t}$ is $\sqrt{\sigma_{0}^{2}+\sigma^{2}}$-subgaussian.
-Proof.
-Start by considering $Z=\sum_{p\in P}\mathds{1}\{p^{*}=p\}(f_{p}(x)-\mu_{t,p}(x))$ where $f_{p}\sim{\mathcal{GP}}_{p}$ such that $f_{p}(x)-\mu_{t,p}(x)$ is $\sigma_{0}$-subgaussian. Then,
-${\mathbb{E}}[\lambda Z|H_{t}]$ $={\mathbb{E}}[{\mathbb{E}}[\exp(\lambda Z)\,|\,p^{*}=p,H_{t}]$ (80)
-$=\sum_{p\in P}{\mathbb{P}}_{t}(p^{*}=p){\mathbb{E}}[\exp(\lambda(f_{p}(x)-\mu_{t,p}(x)))\,|\,p^{*}=p,H_{t}]$ (81)
-$\leq\sum_{p\in P}{\mathbb{P}}_{t}(p^{*}=p)\exp(\sigma_{t,p}^{2}(x)\lambda^{2}/2)$ (82)
-$\leq\sum_{p\in P}{\mathbb{P}}_{t}(p^{*}=p)\exp(\sigma_{0}^{2}\lambda^{2}/2)$ (83)
-$=\exp(\sigma^{2}_{0}\lambda/2)$ (84)
-Hence, $Z$ is $\sigma$-subgaussian. Then, $Z+\epsilon_{t}|H_{t}$ is $\sqrt{\sigma^{2}_{0}+\sigma^{2}}$-subgaussian since $Z|H_{t}$ and $\epsilon_{t}|H_{t}$ are independent. ∎
-Then, we state and prove the information-theoretic regret bound for HP-GP-TS. See 4.3
-Proof.
-Here, we analyze the regret of HP-GP-TS using the information-theoretic framework of Russo & Van Roy (2016).
-The proof presented here follows the proof in section D.2 of Russo & Van Roy (2016) with subgaussian noise.
-The idea of the information-theoretic framework is to express the regret as follows
-${\mathbb{E}}\left[{\sum_{t\in[T]}}f(x^{*})-f(x_{t})\right]$ $={\mathbb{E}}\left[{\sum_{t\in[T]}}{\mathbb{E}}\big[\underbrace{f(x^{*})-f(x_{t})}_{\geq 0}|H_{t}\big]\right]$ $(\text{Tower rule})$ (85)
-$={\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{{\mathbb{E}}\left[f(x^{*})-f(x_{t})|H_{t}\right]^{2}\cdot\frac{I(\cdot\,;\cdot\,|H_{t})}{I(\cdot\,;\cdot\,|H_{t})}}\right]$ (86)
-where $I(\cdot\,;\cdot\,|H_{t})$ is the mutual information between two carefully chosen variables such that ${\mathbb{E}}\left[f(x^{*})-f(x_{t})|H_{t}\right]^{2}\leq CI(\cdot\,;\cdot\,|H_{t})$ for some $C$.
-Let ${\mathbb{E}}_{t}[\cdot]={\mathbb{E}}[\cdot|H_{t}],{\mathbb{P}}_{t}[\cdot]={\mathbb{P}}[\cdot|H_{t}]$. Then,
-${\mathbb{E}}_{t}[f(x^{*})-f(x_{t})]=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x){\mathbb{E}}_{t}[f(x)|x^{*}=x]-\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x){\mathbb{E}}_{t}[f(x)|x_{t}=x]$ (87)
-$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x)\left({\mathbb{E}}_{t}[f(x)|x^{*}=x]-{\mathbb{E}}_{t}[f(x)|x_{t}=x]\right)\quad$ $\left({\mathbb{P}}_{t}(x_{t}=x)={\mathbb{P}}_{t}(x^{*}=x)\right)$ (88)
+Appendix B Technical issues with MixTS regret bound in the linear setting
+Theorem 1 in Hong et al.
+(2022a) provides a regret bound for MixTS in the linear setting.
+The linear setting assumes that the true parameter $\theta^{*}\sim{\mathcal{N}}(\theta_{0,S_{*}},\Sigma_{0,S_{*}})$ where the latent state $S_{*}$ is sampled from a discrete mixture prior $P_{0}$.
+The proof of Theorem 1 in Hong et al.
+(2022a) contains a few non-obvious steps that seem difficult to motivate.
+We use the notation of Hong et al.
+(2022a).
+  1. 1.
+    Eq.
+(9) in Hong et al.
+(2022a) uses the TS property that the true prior and optimal arm is equal in distribution to the selected prior and selected arm given the history: $A_{t,*}^{\top}\bar{\theta}_{t,S_{*}}|H_{t}\overset{d}{=}A_{t}^{\top}\bar{\theta}_{t,S_{t}}|H_{t}$ (equivalent to $\mu_{t,p^{*}}(x^{*})|H_{t}\overset{d}{=}\mu_{t,p_{t}}(x_{t})|H_{t}$ in our notation).
+However, Eq.
+(9) additionally conditions on the event $E_{0}=\{\|\theta_{*}-\theta_{0,S_{*}}\|_{\Sigma_{0,S_{*}}^{-1}}\leq\sqrt{2d\log(dn)}\}$ where $\theta_{*}$ lies close to its prior mean.
+The TS property does not hold under this event since it modifies the distribution of the linear parameter $\theta^{*}$ but not the sampled parameters $\theta_{t}$, thus changing the distribution of the optimal arm $A_{t,*}$ but not the selected arm $A_{t}$.
+Consider the example in Figure˜8, if $E_{0}$ holds then $\theta_{*}$ lies in the blue regions and thus $a_{2}$ is optimal w.p.
+0.
+If $E_{0}^{c}$ holds, then $a_{2}$ is optimal with a non-zero probability.
+However, MixTS is oblivious to $E_{0}$ given the history and thus $A_{t,*}|H_{t},E_{0}\not\overset{d}{=}A_{t}|H_{t}$.
+This counterexample illustrates the overall idea but we have not validated that the scale of the arms and the blue regions are feasible.
+  2. 2.
+    Five lines above Eq.
+(9) in Hong et al.
+(2022a), it is stated that the regret is upper-bounded by a constant $M$ whenever $E_{0}$ occurs.
+However, from Eq (9) to the first term in step 3 of their analysis (page 15), the bound of $M$ is applied implicitly to $A_{t}^{\top}\bar{\theta}_{t,S_{t}}-A_{t}^{\top}\theta_{*}$ without motivation.
+For the setting with bounded rewards, then $\bar{\theta}_{t,S_{t}}$ is also bounded but for Gaussian rewards $\bar{\theta}_{t,S_{t}}$ can be unbounded.
+  3. 3.
+    The second term in Eq.
+(9) contains the indicator function $\mathbf{1}\{E_{0}\}$: $\mathbb{E}[(A_{t}^{\top}\bar{\theta}_{t,S_{t}}-A_{t}^{\top}\theta_{*})\mathbf{1}\{E_{0}\}]$.
+In step 3 (page 15), this indicator function is dropped without motivation: $\mathbb{E}[\langle A_{t}^{\top}\bar{\theta}_{t,S_{t}}-A_{t}^{\top}\theta_{*}\rangle_{M}]$ where $\langle\cdot\rangle_{M}=\min(\cdot,M)$ for the bound $M$.
+If the expression inside is non-negative w.p.
+1, then this step would be valid but this is not the case.
+  4. 4.
+    From our understanding, the final equation on page 15 adds and subtracts the confidence bound and adds a zero-mean Gaussian inside a minimum.
+However, adding a zero-mean Gaussian inside a minimum reduces the expectation but the analysis seems to assume that it would increase the expectation.
+I.e.
+it is seemingly assumed that $\mathbb{E}[\min(M,X)]\leq\mathbb{E}[\min(M,X+\epsilon_{t})$ for a constant M and random variable $X$.
+However, the reverse inequality is true.
+Theorem 1 of Hong et al. (2022a) provides a novel proof technique that could potentially be applied to HP-GP-TS. But the technical issues presented above present a few challenges.
+Figure 8: Potential counterexample to Eq (9) in Hong et al. (2022a). The blue regions represent when the event $E_{0}$ holds and the black dots represent the two arms $a_{1}$ and $a_{2}$.
+Appendix C Extending the analysis of MixTS to HP-GP-TS
+In this section, we explore how the analysis of Hong et al.
+(2022a) can be applied to the Gaussian process setting considered in this paper.
+The core idea of the analysis of Hong et al.
+(2022a) is to introduce confidence sets for the priors (or latent variables in their terminology).
+The confidence intervals closely resemble the prior elimination criteria introduced by Ziomek et al.
+(2025).
+In our notation, the equivalent confidence set $C_{t}$ of Hong et al.
+(2022a) is defined as
+$C_{t}=\left\{p\in P:\sum_{s=1}^{t-1}\mathds{1}\{p_{s}=p\}(\mu_{s,p}(x_{s})-\sqrt{\beta_{T}}\sigma_{s,p_{s}}(x_{s})-y_{s})\leq 2\sigma\sqrt{N_{t}(p)\log T}\right\}.$ (81)
+We define $G_{t}(p)=\sum_{s=1}^{t-1}\mathds{1}\{p_{s}=p\}(\mu_{s,p}(x_{s})-\sqrt{\beta_{T}}\sigma_{s,p_{s}}(x_{s})-y_{s})$.
+The prior elimination criteria for PE-GP-UCB and PE-GP-TS can be expressed as a confidence set in the following way:
+$\tilde{C}_{t}=\Bigg\{p\in P:\forall s<t\quad\Bigg|\sum_{\begin{subarray}{c}i\leq s\\ p_{i}=p\end{subarray}}y_{i}-\mu_{i,p}(x_{i})\Bigg|\leq\sqrt{2\sigma^{2}N_{t}(p)\log(t^{2}|P|\pi^{2}/3)}+\sum_{\begin{subarray}{c}i\leq s\\ p_{i}=p\end{subarray}}\sqrt{\beta_{i}}\sigma_{i,p}(x_{i})\Bigg\}.$ (82)
+The main difference between $C_{t}$ and $\tilde{C}_{t}$ is that $\tilde{C}_{t}$ is monotonically decreasing in cardinality since eliminated priors cannot be selected by PE-GP-UCB and -TS.
+Additionally, $\tilde{C}_{t}$ uses a symmetric criteria and the confidence parameters in $C_{t}$ use the final time step $T$ rather than the current time step $t$.
+To apply the analysis of Hong et al.
+(2022a) directly to a GP-setting would require resolving the issues discussed in Appendix˜B.
+To circumvent some of these issues, we attempt to analyze the regret of HP-GP-TS by directly applying a UCB-decomposition and manipulate the decomposition such an expression similar to $G_{t}(p_{t})$ appears.
+Recall, the regular UCB-decomposition to the Bayesian cumulative regret:
+${\text{BR}}(T)$ $={\sum_{t\in[T]}}{\mathbb{E}}[f(x^{*})-f(x_{t})]$ (83)
+$={\sum_{t\in[T]}}{\mathbb{E}}[f(x^{*})-U_{t,p^{*}}(x^{*})+U_{t,p^{*}}(x^{*})-f(x_{t})]$ (84)
+$=\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}[f(x^{*})-U_{t,p^{*}}(x^{*})]}_{(1)}+\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p_{t}}(x_{t})-f(x_{t})]}_{(2)}\quad(p_{t},x_{t}|H_{t}\overset{d}{=}p^{*},x^{*}|H_{t})$ (85)
+By the same steps as in the proof of Theorem˜4.2, $(1)\leq\pi^{2}/6$. Hence, we focus on the second term and condition on whether the selected prior lies in the confidence set:
+$(2)={\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t}))\mathds{1}\{p_{t}\in C_{t}\}]+{\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t}))\mathds{1}\{p_{t}\notin C_{t}\}].$ (86)
+To bound the right-term above, we would like $(U_{t,p_{t}}(x_{t})-f(x_{t}))$ to be bounded such that it is sufficient to analyze the probability $\mathds{1}\{p_{t}\notin C_{t}\}$.
+However, we have not identified a method of doing so.
+We continue to study the left term.
+Since the noise $\epsilon_{t}$ is independent of $\mathds{1}\{p_{t}\in C_{t}\}$ and has zero mean, we can add it inside the expectation and add-and-subtract the confidence radius:
+${\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t}))\mathds{1}\{p_{t}\in C_{t}\}]$ (87)
+$={\sum_{t\in[T]}}{\mathbb{E}}[(U_{t,p_{t}}(x_{t})-f(x_{t})-\epsilon_{t})\mathds{1}\{p_{t}\in C_{t}\}]$ (88)
+$={\sum_{t\in[T]}}{\mathbb{E}}[(\mu_{t,p_{t}}(x_{t})-\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-f(x_{t})-\epsilon_{t})\mathds{1}\{p_{t}\in C_{t}\}]$ (89)
+$\quad+{\sum_{t\in[T]}}2{\mathbb{E}}[\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\mathds{1}\{p_{t}\in C_{t}\}]$ (90)
+$\leq\underbrace{{\sum_{t\in[T]}}{\mathbb{E}}[(\mu_{t,p_{t}}(x_{t})-\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})-f(x_{t})-\epsilon_{t})\mathds{1}\{p_{t}\in C_{t}\}]}_{(3)}$ (91)
+$\quad+{\sum_{t\in[T]}}2{\mathbb{E}}[\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})]$ $(\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})\geq 0)$ (92)
+Note that the sum of $\sqrt{\beta_{t}}\sigma_{t,p_{t}}(x_{t})$ can be bounded in terms of the worst-case MIG $\hat{\gamma}_{T}$} by Equation˜16.
+To bound (3), following the argument of Hong et al.
+(2022a), we note that for each prior $p\in P$ there must be some last time $t^{\prime}(p)\leq T$ that $p$ was in the confidence set and chosen by the algorithm: $p\in C_{t^{\prime}}$ and $p=p_{t^{\prime}}$.
+Since $p\in C_{t^{\prime}}$, then $G_{t^{\prime}}(p)\leq 2\sigma\sqrt{N_{t^{\prime}}(p)\log T}$.
+Therefore, for each prior $p$, we can split the sum into $G_{t^{\prime}}(p)$ and the final term at time $t^{\prime}$:
+$(3)$ $\leq{\mathbb{E}}\bigg[\sum_{p\in P}(G_{t^{\prime}}(p)+(\mu_{t^{\prime},p}(x_{t^{\prime}})-\sqrt{\beta_{t^{\prime}}}\sigma_{t^{\prime},p}(x_{t^{\prime}})-f(x_{t^{\prime}})-\epsilon_{t^{\prime}}))\mathds{1}\{p\in C_{t^{\prime}}\}\bigg]$ (93)
+$\leq{\mathbb{E}}\bigg[\sum_{p\in P}2\sigma\sqrt{N_{t^{\prime}}(p)\log T}\bigg]$ (94)
+$\quad+{\mathbb{E}}\bigg[\sum_{p\in P}(\mu_{t^{\prime},p}(x_{t^{\prime}})-\sqrt{\beta_{t^{\prime}}}\sigma_{t^{\prime},p}(x_{t^{\prime}})-f(x_{t^{\prime}})-\epsilon_{t^{\prime}}))\mathds{1}\{p\in C_{t^{\prime}}\}\bigg].$ (95)
+The first sum above can be bounded using that $\sum_{p\in P}N_{T}(p)=T$ and Jensen’s inequality.
+The major challenge is to show that the final term is bounded.
+This requires creating an event s.t.
+the terms inside are bounded with high probability.
+However, as the issues in Appendix˜B highlight, if the event is in introduced at an earlier step then it must not invalidate the subsequent manipulations.
+Appendix D Description of kernels
 […]
-Let $Z=f(x)+\epsilon_{t}-{\mathbb{E}}_{t}[f(x)]$.
-Note that $Z|H_{t}$ is $\sqrt{\sigma_{0}^{2}+\sigma^{2}}$-subgaussian.
-Consequently, $\log{\mathbb{E}}_{t}[\exp(\lambda Z)]\leq\frac{\lambda^{2}(\sigma_{0}^{2}+\sigma^{2})}{2}$ $\forall\lambda\in{\mathbb{R}}$.
-Fix $x^{*},x\in{\mathcal{X}}$. Then,
-$\lambda\Big({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\Big)-\frac{\lambda^{2}(\sigma_{0}^{2}+\sigma^{2})}{2}$ (90)
-$\leq\lambda\Big({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\Big)$ (91)
-$\quad-\log{\mathbb{E}}_{t}[\exp(\lambda(f(x)+\epsilon_{t}-{\mathbb{E}}_{t}[f(x)])]$ $(Z|H_{t}\text{ is }\sqrt{\sigma_{0}^{2}+\sigma^{2}}\text{-subgaussian})$ (92)
-$=\lambda{\mathbb{E}}_{t}[Z|x^{*}=x^{*}]-\log{\mathbb{E}}_{t}[\exp(\lambda Z)]$ (93)
-[…]
-Now, let $\lambda=\frac{1}{\sigma_{0}^{2}+\sigma^{2}}\left({\mathbb{E}}_{t}[f(x)+\epsilon_{t}|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\right)$, then the following holds for all $x^{*},x\in{\mathcal{X}}$
-${\mathbb{E}}_{t}\left[f(x)|x^{*}=x^{*}\right]-{\mathbb{E}}_{t}[f(x)]\leq\sqrt{2(\sigma_{0}^{2}+\sigma^{2})D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})||{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)}.$ (95)
-Let $I_{t}(\cdot\,;\cdot\,)=I(\cdot\,;\cdot\,|H_{t}=h_{t})$. Then,
-$I_{t}(x^{*};(x_{t},y_{t}))=I_{t}(x^{*};x_{t})+I_{t}(x^{*};y_{t}|x_{t})$ $(\text{Chain rule})$ (96)
-$=I_{t}(x^{*};y_{t}|x_{t})$ $(x^{*}|H_{t}{\perp\!\!\!\!\perp}x_{t}|H_{t})$ (97)
-$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};y_{t}|x_{t}=x)$ (98)
-$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};f(x_{t})+\epsilon_{t}|x_{t}=x)$ (99)
-$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)I_{t}(x^{*};f(x)+\epsilon_{t})$ $(f,x^{*}|H_{t}{\perp\!\!\!\!\perp}x_{t}|H_{t})$ (100)
-$=\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x_{t}=x)\Bigg(\sum_{x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ (101)
-$\quad\quad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})\,\Big|\Big|\,{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)\Bigg)$ (102)
-$=\sum_{x,x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ $(x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t})$ (103)
-$\quad\quad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})\,\Big|\Big|\,{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)$ (104)
-Putting the above together, we now bound ${\mathbb{E}}_{t}[f(x^{*})-f(x_{t})]^{2}$:
-${\mathbb{E}}_{t}[f(x^{*})-f(x_{t})]^{2}=\left(\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x)\left({\mathbb{E}}_{t}[f(x)|x^{*}=x]-{\mathbb{E}}_{t}[f(x)]\right)\right)^{2}$ (105)
-$\leq|{\mathcal{X}}|\sum_{x\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x)^{2}\left({\mathbb{E}}_{t}[f(x)|x^{*}=x]-{\mathbb{E}}_{t}[f(x)]\right)^{2}$ $(\text{Cauchy-Schwarz})$ (106)
-$\leq|{\mathcal{X}}|\sum_{x,x^{*}\in{\mathcal{X}}}{\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\left({\mathbb{E}}_{t}[f(x)|x^{*}=x^{*}]-{\mathbb{E}}_{t}[f(x)]\right)^{2}$ (107)
-$\leq 2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})\sum_{x,x^{*}\in{\mathcal{X}}}\Big({\mathbb{P}}_{t}(x^{*}=x){\mathbb{P}}_{t}(x^{*}=x^{*})\cdot$ (108)
-$\qquad\qquad\qquad\qquad\qquad D\left({\mathbb{P}}_{t}(f(x)+\epsilon_{t}|x^{*}=x^{*})||{\mathbb{P}}_{t}(f(x)+\epsilon_{t})\right)\Big)$ (109)
-$\leq 2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})I_{t}(x^{*};(x_{t},y_{t})).$ (110)
-Returning to the full regret,
-${\mathbb{E}}\bigg[{\sum_{t\in[T]}}f(x^{*})-f(x_{t})\bigg]$ $={\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{{\mathbb{E}}_{t}\left[f(x^{*})-f(x_{t})\right]^{2}\cdot\frac{I_{t}(x^{*}\,;(x_{t},y_{t}))}{I_{t}(x^{*}\,;(x_{t},y_{t}))}}\right]$ (111)
-$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}{\mathbb{E}}\left[{\sum_{t\in[T]}}\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\right]$ (112)
-Then, the expectation can then be bounded as
-${\mathbb{E}}\bigg[{\sum_{t\in[T]}}\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\bigg]=$ $\leq{\sum_{t\in[T]}}{\mathbb{E}}\left[\sqrt{I_{t}(x^{*}\,;(x_{t},y_{t}))}\right]$ (114)
-$\leq{\sum_{t\in[T]}}1\cdot\sqrt{{\mathbb{E}}\left[I_{t}(x^{*}\,;(x_{t},y_{t}))\right]}$ $\left(\text{Jensen's inequality}\right)$ (115)
-$={\sum_{t\in[T]}}\sqrt{{\mathbb{E}}_{H_{t}}\left[I(x^{*}\,;(x_{t},y_{t})|H_{t}=h_{t})\right]}$ (116)
-$={\sum_{t\in[T]}}\sqrt{I(x^{*}\,;(x_{t},y_{t})|H_{t})}$ (117)
-$\leq\sqrt{{\sum_{t\in[T]}}1^{2}{\sum_{t\in[T]}}I(x^{*}\,;(x_{t},y_{t})|H_{t})}$ $(\text{Cauchy-Schwarz})$ (118)
-Finally, the regret can be bounded as
-${\mathbb{E}}\bigg[{\sum_{t\in[T]}}f(x^{*})-f(x_{t})\bigg]$ $=\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{TI(x^{*};H_{T})}$ (119)
-$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{TH(x^{*})}$ (120)
-$\leq\sqrt{2|{\mathcal{X}}|(\sigma^{2}_{0}+\sigma^{2})}\sqrt{T\log|{\mathcal{X}}|}$ (121)
-Appendix B Description of kernels
-[…]
-Appendix C Additional experimental details
+Appendix E Additional experimental details
 In this section, we provide some additional details about the experiments.
-For all the real-world datasets, sensors containing any null measurements have been filtered out.
+For the kernel experiment, all kernels use a lengthscale of $1.0$ and are scaled s.t.
+$k(x,\tilde{x})\leq 1$.
+In addition, the mean function for all priors is zero everywhere.
+For the subspace experiment, the total dimensions $d=16$ but each prior $p_{i}$ assumes $f(x)$ depends on $d_{s}=4$ subdimensions: $[i,i+1,i+2,i+3]$ for $i\in[5]$.
+Dimensions larger than $5$ are wrapped around 1, i.e.
+$((j-1)\mod 5)+1$, such that the priors are equally difficult to distinguish and optimize.
+The prior elimination methods use $\delta=0.05$ across all experiments, including the oracle methods.
+During every iteration $t$, SCoreBO samples $M$ priors from the hyperposterior $P_{t}$ and samples $N$ optimizers $x^{*},f^{*}$ for each prior sampled through posterior sampling.
+We use $M=16$ and $N=12$ for SCoreBO in all experiments.
+We use the implementation of the SCoreBO acquisition function in BoTorch (Community) (Balandat et al., 2020).
+To make the implementation fast with GPUs, we set linear_operator.settings.stable_qr_cpu_threshold to 8 in order to avoid QR-factorization being performed on CPU (Gardner et al., 2018; Pleiss et al., 2022, 2025).
+To avoid out of memory issues, we replace the default torch.matmul in DefaultPredictionStrategy._exact_predictive_covar_inv_quad_form_root (from gpytorch.models.exact_prediction_strategies) with an equivalent torch.einsum (Gardner et al., 2018).
+Since the priors in our experiments are discrete, we compute the hyperposterior exactly and sample from it directly.
+For all the real-world datasets, sensors containing any null measurements are filtered out.
 The Intel Berkeley dataset consists of measurements from 46 temperature sensors across 19 days.
@@
 The noise variance is set to $\sigma^{2}=2.25^{2}$.
-The PNW precipication data consists of daily precipication data from 1949 to 1950 across 167 $50\times 50$ km regions in the Pacific Northwest.
+The PNW precipitation data consists of daily precipitation data from 1949 to 1994 across 167 $50\times 50$ km regions in the Pacific Northwest.
 The goal is to find the region with the highest precipitation for any given day.
@@
 In the Intel experiment, we removed one outlier seed.
-All methods had a final cumulative regret around 6000°C, note that the average for the worst performing model across the other seeds was $\approx 150$°C.
-The outlier is shown Fig.˜8.
+All methods had a final cumulative regret around 6000°C on this instance, note that the average for the worst performing model across the other seeds was $\approx 150$°C.
+The outlier is shown Figure˜9.
 We can see that one of the sensors display very high temperatures compared to all other sensors, which is why all methods performed poorly on this seed.
@@
 Figure 9: Removed sample from the test data in the Intel experiment. One of the sensors display very high temperatures.
-Appendix D Additional experimental results
+Appendix F Additional experimental results
 In this section, we provide some additional experimental results.
+Table 1: Average total regret and $\pm$ 1 standard error for the synthetic and real-world data experiments. The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
+Algorithm Synthetic Real-world data
+Kernel Lengthscale Subspace Intel PeMS PNW Precip.
+MAP GP-TS 84.3 $\pm$ 8.4 30.2 $\pm$ 1.2 87.2 $\pm$ 1.0 73.8 $\pm$ 7.7 1635.0 $\pm$ 129.3 178.4 $\pm$ 6.9
+HP-GP-TS 39.2 $\pm$ 1.4 31.4 $\pm$ 1.0 88.3 $\pm$ 0.9 54.1 $\pm$ 3.0 1327.8 $\pm$ 107.9 167.7 $\pm$ 5.2
+PE-GP-TS 62.0 $\pm$ 0.6 61.8 $\pm$ 0.5 177.1 $\pm$ 1.4 106.5 $\pm$ 2.1 1214.2 $\pm$ 81.5 200.9 $\pm$ 4.0
+PE-GP-UCB 121.6 $\pm$ 1.2 114.2 $\pm$ 0.6 389.0 $\pm$ 1.5 173.0 $\pm$ 2.7 2159.2 $\pm$ 48.4 506.2 $\pm$ 2.6
+Oracle GP-TS 35.0 $\pm$ 1.1 28.1 $\pm$ 0.8 86.0 $\pm$ 1.0
+Oracle GP-UCB 68.5 $\pm$ 1.9 48.3 $\pm$ 1.2 217.3 $\pm$ 1.0
+SCoreBO 180.4 $\pm$ 7.7 180.8 $\pm$ 5.8 106.6 $\pm$ 0.9 256.8 $\pm$ 9.3 3460.2 $\pm$ 163.7 861.3 $\pm$ 21.0
+EEI 39.0 $\pm$ 2.6 30.1 $\pm$ 2.1 88.3 $\pm$ 4.2 51.6 $\pm$ 4.8 1664.1 $\pm$ 137.7 196.5 $\pm$ 12.3
+First, we provide the average total regret for the synthetic and real-world data experiments in Table˜1.
+We observe that HP-GP-TS either has the lowest regret or is within 1 standard error of the algorithm with the lowest regret across all the experiments.
+In Figure˜3, it can be noted had low regret early in the synthetic experiment but HP-GP-TS either catches up or almost catches up later in the experiments.
+We compare both algorithms with an extended time horizon of $T=1500$, the results are shown in Figures˜10 and 2.
+With the extended time horizon, HP-GP-TS achieves the lowest regret across all synthetic experiments.
+Although, EEI is still within 1 standard error on the lengthscale experiment.
+Figure 10: Cumulative regret for synthetic experiments extended time horizon $T=1500$ with varying kernel (left), lengthscale (center) and mean function (right).
+Errorbars correspond to $\pm 1$ standard error.
+Table 2: Average total regret and $\pm$ 1 standard error for the synthetic experiments with longer horizon $T=1500$.
+The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
+Algorithm Kernel Lengthscale Subspace
+HP-GP-TS 49.1 $\pm$ 1.6 39.7 $\pm$ 1.2 103.4 $\pm$ 1.3
+EEI 62.6 $\pm$ 6.2 45.7 $\pm$ 5.0 138.9 $\pm$ 9.2
 Figure 11: Mean number of priors remaining in $P_{t}$ over time for PE-GP-UCB and -TS.
@@
 $\frac{1-q}{|P|-1}$ on the other |P|1|P|-1 choices.
-First, we include the mean number of priors in $P_{t}$ for all experiments in Fig.˜9.
-Similarly, we include the average entropy of the hyperposterior for all experiments in Fig.˜10.
-For the lengthscale, subspace, PeMS and PNW precipication experiments, hardly any priors are eliminated.
-In contrast, the hyperposterior entropy concentrates rapidly across all experiments with the subspace and PNW precipication having the most and least concentrated hyperposterior.
-We include the full set of confusion matrices for the lengthscale and subspace experiments in Fig.˜11.
+Next, we include the mean number of priors in $P_{t}$ for all experiments in Figure˜11.
+Similarly, we include the average entropy of the hyperposterior for all experiments in Figure˜12.
+For the lengthscale, subspace, PeMS and PNW precipitation experiments, hardly any priors are eliminated.
+In contrast, the hyperposterior entropy concentrates rapidly across all experiments with the subspace and PNW precipitation having the most and least concentrated hyperposteriors.
+The cost of learning the prior, ${\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$, for HP-GP-TS is shown in Figure˜13.
+Whilst this term has not been bounded theoretically in Theorem˜4.2, we observe that empirically its growth diminishes much quicker than the cumulative regret.
+Since ${\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$ is one term in the upper bound for the regret of HP-GP-TS, learning the prior does not seem to increase the growth rate of the regret empirically.
+Figure 13: The cost of learning the prior, ${\sum_{t\in[T]}}U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})$, for HP-GP-TS in the synthetic experiments.
+The solid blue line corresponds to the mean and the dashed lines corresponds to the first and last decile at each time step.
+We include the full set of confusion matrices for the lengthscale and subspace experiments in Figure˜14.
 In the lengthscale experiments, we observe that PE-GP-UCB and -TS oversample the shortest lengthscale.
@@
 In the subspace experiment, HP- and MAP GP-TS have an accuracy of around 96% where as PE-GP-TS and -UCB have accuracies 30% and 36% respectively.
-The priors are equivalent up to coordinate permutations and are therefore difficult.
-The PE-methods do not oversample any specific prior but commit to much time to exploring along the irrelevant dimensions.
-In Tables˜1 and 2, the total regret for the lengthscale and subspace scaling experiments are shown.
-In Fig.˜12, we visualize the median cumulative regret on the real-world data experiments.
-Similarly, in Fig.˜13, we show further quantiles of the final cumulative regret.
-MAP and HP-GP-TS exhibit similar median regret across the experiments but MAP GP-TS has a larger variance and tail values.
+Even though PE-GP-UCB has a higher accuracy than PE-GP-TS, it still has significantly higher regret.
+Additionally, the priors are equivalent up to coordinate permutations and therefore generate functions that are equally difficult to optimize.
+Unlike the kernel and lengthscale experiments, the PE-methods do not oversample any specific prior but commit too much time to exploring along the irrelevant dimensions.
+In Tables˜3 and 4, the total regret for the lengthscale and subspace scaling experiments are shown.
 (a) Kernel experiment
@@
 Figure 14: Confusion matrices for the true prior $p^{*}$ and $p_{t}$ across all time steps of the synthetic experiments.
-Table 1: Average total regret and $\pm 1$ standard error for the lengthscale experiment as $|P|$ increases.
+Table 3: Average total regret and $\pm 1$ standard error for the lengthscale experiment as $|P|$ increases. The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
 Algorithm Lengthscales, $|P|$
@@
 Oracle GP-UCB 48.3 $\pm$ 1.2 46.9 $\pm$ 1.1 48.4 $\pm$ 1.1 46.5 $\pm$ 1.0 45.6 $\pm$ 1.0
-Table 2: Average total regret and $\pm 1$ standard error for the subspace experiment as $|P|$ increases.
+SCoreBO 180.8 $\pm$ 5.8 240.3 $\pm$ 6.3 277.2 $\pm$ 6.8 281.6 $\pm$ 6.7 283.9 $\pm$ 6.8
+EEI 30.1 $\pm$ 2.1 30.9 $\pm$ 2.2 29.4 $\pm$ 2.2 30.8 $\pm$ 2.6 32.1 $\pm$ 2.8
+Table 4: Average total regret and $\pm 1$ standard error for the subspace experiment as $|P|$ increases. The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
 Algorithm Subspaces, $|P|$
@@
 Oracle GP-UCB 217.3 $\pm$ 1.0 218.2 $\pm$ 1.0 218.6 $\pm$ 1.0 218.9 $\pm$ 0.9
-Figure 12: Median cumulative regret on Intel temperature data (top), PeMS speed data (middle) and PNW precipication data (bottom). Errorbars correspond to first and last decile.
-(a) Intel
-(b) PeMS
-(c) PNW Precipication
-Figure 13: Quantiles of the final cumulative regret on the real-world data experiments.
-The median is shown with a black line.
-The whiskers correspond to the 5th and 95th percentile and the lower and upper edges of the boxes show the first and third quantile.
+SCoreBO 106.6 $\pm$ 0.9 108.2 $\pm$ 0.8 108.9 $\pm$ 0.7 109.5 $\pm$ 0.7
+EEI 88.3 $\pm$ 4.2 81.3 $\pm$ 3.8 82.5 $\pm$ 3.6 81.3 $\pm$ 3.8